### v1

# Computable Bases

Abstract.
In computable analysis typically topological spaces with countable bases are considered. The Theorem of Kreitz-Weihrauch implies that the subbase representation of a second-countable ${\mathsf{{T}}}_{0}$ space is admissible with respect to the topology that the subbase generates. We consider generalizations of this setting to bases that are representable, but not necessarily countable. We introduce the notions of a computable presubbase and a computable prebase. We prove a generalization of the Theorem of Kreitz-Weihrauch for the presubbase representation that shows that any such representation is admissible with respect to the topology generated by compact intersections of the presubbase elements. For computable prebases we obtain representations that are admissible with respect to the topology that they generate. These concepts provide a natural way to investigate many topological spaces that have been studied in computable analysis. The benefit of this approach is that topologies can be described by their usual subbases and standard constructions for such subbases can be applied. Finally we discuss a Galois connection between presubbases and representations of ${\mathsf{{T}}}_{0}$ spaces that indicates that presubbases and representations offer particular views on the same mathematical structure from different perspectives.
1991 Mathematics Subject Classification:
[Theory of computation]: Logic; [Mathematics of computing]: Continuous mathematics.
Version of October 10, 2025.
1. Introduction
In the representation based approach to computable analysis, as initiated by Kreitz and Weihrauch [KW85, Wei87, Wei00], we work with representations $\delta:\subseteq{\mathbb{N}}^{\mathbb{N}}\to X$, which are surjective partial maps. In this situation $(X,\delta)$, or briefly $X$, is called a represented space. We say that a function $F:\subseteq{\mathbb{N}}^{\mathbb{N}}\to{\mathbb{N}}^{\mathbb{N}}$ realizes a multivalued map $f:\subseteq X\rightrightarrows Y$ on represented spaces $(X,\delta_{X})$ and $(Y,\delta_{Y})$, in symbols, $F\vdash f$, if $\delta_{Y}F(p)\in f\delta_{X}(p)$ for all $p\in{\mathrm{dom}}(f\delta_{X})$. Then $f$ is called computable or continuous, if it has a realizer of the corresponding type.
On the other hand, every represented space $(X,\delta_{X})$ comes naturally equipped with a topology ${\mathcal{O}}(X)$, namely the final topology
${\mathcal{O}}(X):=\{U\subseteq X:\delta_{X}^{-1}(U)\mbox{ open in }{\mathrm{dom}}(\delta_{X})\}$
induced by the Baire space topology on $\mathbb{N}$$^{\mathbb{N}}$ via the map $\delta_{X}$ on $X$. If not mentioned otherwise, we will assume that ${\mathcal{O}}(X)$ is the final topology of the represented space $X$. The topological spaces whose topology occurs in this way have a special name, they are called qcb spaces (quotients of countably based spaces). These spaces share some special properties [Sch02b], for instance, they are always sequential, i.e., a subset is closed if and only if it is sequentially closed, which means that it is closed under limits of converging sequences. We also automatically have a representation $\delta_{{\mathcal{O}}(X)}$ of this topology ${\mathcal{O}}(X)$ that is induced by a function space representation of the space ${\mathcal{C}}(X,{\mathbb{S}})$ of the continuous functions $f:X\to{\mathbb{S}}$ with the Sierpiński space ${\mathbb{S}}=\{0,1\}$. This is because $U\subseteq X$ is open if and only if its characteristic function $\chi_{U}:X\to{\mathbb{S}}$ is continuous. Matthias Schröder has developed a theory of computable topology based on these concepts [Sch02b, Sch02a, Sch21] (see also Pauly [Pau16] for a concise presentation of some aspects of computable topology).
An essential question is when continuity in terms of the representations $\delta_{X}$ and $\delta_{Y}$ (as defined above) coincides for a singlevalued function $f:X\to Y$ with the ordinary notion of continuity with respect to the topologies ${\mathcal{O}}(X)$ and ${\mathcal{O}}(Y)$. Every function $f$ with a continuous realizer is continuous in the ordinary topological sense, but not necessarily the other way around [Sch02b]. The two concepts of continuity are equivalent if $\delta_{X}$ and $\delta_{Y}$ are so-called admissible representations with respect to these topologies. The notion of admissibility was originally defined by Kreitz and Weihrauch just for second-countable ${\mathsf{{T}}}_{0}$ spaces [KW85] and then later extended to arbitrary topological spaces by Schröder [Sch02b]. We will adopt the latter definition.
Definition 1 (Admissibility).
A representation $\delta_{X}:\subseteq{\mathbb{N}}^{\mathbb{N}}\to X$ is called admissible with respect to a topology $\tau$ on $X$ if the following hold:
  1. (1)
    $\delta_{X}$ is continuous with respect to the topology $\tau$ and
  2. (2)
    any other representation $\delta:\subseteq{\mathbb{N}}^{\mathbb{N}}\to X$ that is continuous with respect to $\tau$ is also continuous with respect to the representation $\delta_{X}$.
A represented space is called admissible if the representation of the space is admissible with respect to the final topology ${\mathcal{O}}(X)$ induced by it.
Admissibility of $\delta_{X}$ means that any continuous representation $\delta$ can be continuously translated into $\delta_{X}$. Hence the names $p\in{\mathbb{N}}^{\mathbb{N}}$ of $\delta_{X}$ provide the minimal information on the points $x\in X$ that is required for the representation to be continuous in the topological sense. Schröder proved [Sch02b, Theorem 7, Lemma 8] that whether a representation $\delta_{X}$ is admissible with respect to some topology $\tau$ only depends on the sequentialization of that topology [^1] The sequentialization ${\mathrm{seq}}(\tau)$ of a topology $\tau$ is the smallest topology that contains $\tau$ and is sequential..
Theorem 2 (Schröder 2002).
Let $(X,\delta_{X})$ be a represented space with final topology ${\mathcal{O}}(X)$ and let $\tau$ be some topology on $X$. Then the following are equivalent:
  1. (1)
    $\delta_{X}$ is admissible with respect to $\tau$.
  2. (2)
    $\delta_{X}$ is admissible with respect to ${\mathrm{seq}}(\tau)$.
If one these conditions holds, then ${\mathrm{seq}}(\tau)={\mathcal{O}}(X)$.
In particular, this means that if $\delta$ is admissible with respect to some topology, then also with respect to ${\mathcal{O}}(X)$. Schröder also showed [Sch02b, Theorem 13] that a topology for which an admissible representation exists is necessarily ${\mathsf{{T}}}_{0}$. The topological spaces that admit an admissible representation are exactly those whose sequentializations are so-called qcb0 spaces (the ${\mathsf{{T}}}_{0}$ spaces among the qcb spaces). Finally, Schröder also proved that the representation $\delta_{{\mathcal{O}}(X)}$ is always admissible with respect to the Scott topology on ${\mathcal{O}}(X)$ [Sch02a] and that admissibility of a represented space can be characterized with the help of the neighborhood map ${\mathcal{U}}:X\to{\mathcal{O}}{\mathcal{O}}(X),x\mapsto\{U\in{\mathcal{O}}(X):x\in U\}$ of the space as follows.
Proposition 3 (Admissibility).
A represented space $X$ is admissible if and only if its neighborhood map ${\mathcal{U}}:X\to{\mathcal{O}}{\mathcal{O}}(X)$ is a continuous embedding with respect to the involved representations.
Being an embedding with respect to the representations means that $\mathcal{U}$} is injective and $\mathcal{U}$} as well as its partial inverse have continuous realizers. Likewise, we define computable embeddings with computable realizers. The map $\mathcal{U}$} is always computable, hence the essential condition is the one that concerns its inverse. In light of this result one can consider admissibility also as an effectivization of the ${\mathsf{{T}}}_{0}$ property as the map $\mathcal{U}$} is injective if and only if the topology ${\mathcal{O}}(X)$ is ${\mathsf{{T}}}_{0}$ (${\mathsf{{T}}}_{0}$ spaces are also known as a Kolmogorov spaces). Consequently, we call $X$ a computable Kolmogorov space if $\mathcal{U}$} is a computable embedding. [^2] We prefer this terminology over the notion of computable admissibility that is also used for this concept, as one and the same topological space can be a computable Kolmogorov space in different inequivalent ways, whereas topological admissibility characterizes an equivalence class. The notion of admissibility is often linked to this uniqueness property.
Kreitz and Weihrauch [KW85] have shown that every second-countable ${\mathsf{{T}}}_{0}$ space has an admissible representation. Namely, if $B=(B_{n})_{n\in{\mathbb{N}}}$ is a subbase of some topology $\tau$ on $X$, then the subbase representation $\delta^{B}:\subseteq{\mathbb{N}}^{\mathbb{N}}\to X$ with
$\delta^{B}(p)=x:\iff{\mathrm{range}}(p)-1=\{n\in{\mathbb{N}}:x\in B_{n}\}$
for all $p\in{\mathbb{N}}^{\mathbb{N}}$ and $x\in X$ is admissible with respect to $\tau$. Here we do not use ${\mathrm{range}}(p)$ but ${\mathrm{range}}(p)-1:=\{n\in{\mathbb{N}}:n+1\in{\mathrm{range}}(p)\}$ in order to allow for the empty set to be enumerated. We formulate the aforementioned result as a theorem.
Theorem 4 (Kreitz-Weihrauch 1985).
If $X$ is a ${\mathsf{{T}}}_{0}$ space with a countable subbase $B$, then the subbase representation $\delta^{B}$ is admissible with respect to the topology of $X$.
This theorem does not only guarantee that there are sufficiently many nice representations, but it also provides a tool that can be used to show that a concrete given representation $\delta_{X}$ is admissible. One just needs to prove that the representation is topologically equivalent to some subbase representation $\delta^{B}$ with respect to a known subbase $B$ of the topology of interest, i.e., that ${\mathrm{id}}:X\to X$ is continuous in both directions with respect to $\delta_{X}$ and $\delta^{B}$, respectively.
One purpose of this article is provide a similar tool for general (not necessarily second-countable) ${\mathsf{{T}}}_{0}$ spaces. A second goal is to develop a theory of computable bases for represented spaces very much along the lines of the second-countable case [BR25]. Finally, we also demonstrate that these concepts have interesting applications.
2. The Presubbase Theorem
We start with the following definition that extends the concept of a subbase representation beyond the countable case.
Definition 5 (Presubbase).
Let $X$ be a set. We call a family $(B_{y})_{y\in Y}$ of subsets of $X$ a presubbase for $X$, if $Y$ is a represented space and its transpose
$B^{\mathsf{{T}}}:X\to{\mathcal{O}}(Y),x\mapsto\{y\in Y:x\in B_{y}\}$
is well-defined and injective.
Injectivity of $B^{\mathsf{{T}}}$ implies that $(B_{y})_{y\in Y}$ is a subbase of some ${\mathsf{{T}}}_{0}$ topology on $X$. We note that every countable subbase $B:{\mathbb{N}}\to{\mathcal{O}}(X)$ of a ${\mathsf{{T}}}_{0}$ topology is a particular instance of a presubbase, as $B^{\mathsf{{T}}}:X\to{\mathcal{O}}({\mathbb{N}})$ is always well-defined. Hence, the following definition generalizes the concept of a subbase representation by Kreitz and Weihrauch as it is specified above.
Definition 6 (Presubbase representation).
Let $(B_{y})_{y\in Y}$ be a presubbase of a set $X$. We define the presubbase representation $\delta^{B}:\subseteq{\mathbb{N}}^{\mathbb{N}}\to X$ by
$\delta^{B}(p)=x:\iff\delta_{{\mathcal{O}}(Y)}(p)=\{y\in Y:x\in B_{y}\}$
for all $p\in{\mathbb{N}}^{\mathbb{N}}$ and $x\in X$.
We note that the map $\delta^{B}$ is well-defined as $B^{\mathsf{{T}}}$ is well-defined and injective. For short, we can also write $\delta^{B}=(B^{\mathsf{{T}}})^{-1}\circ\delta_{{\mathcal{O}}(Y)}$. The reason that we speak about a presubbase representation in this general situation and not about a subbase representation is that $\delta^{B}$ is not necessarily admissible with respect to the topology generated by $(B_{y})_{y\in Y}$. However, it is admissible with respect to a closely related topology generated by compact intersections of the sets $B_{y}$, which is our first main result. By a compact set $K\subseteq X$ we mean a set with the property that every open cover has a finite subcover (no Hausdorff condition involved).
Theorem 7 (Presubbase theorem).
Let $(B_{y})_{y\in Y}$ be a presubbase of a set $X$. Then $(X,\delta^{B})$ is a computable Kolmogorov space and $\delta^{B}$ is admissible with respect to the topology $\tau$ on $X$ that is generated by the base sets $X$ and $\bigcap_{y\in K}B_{y}$ for every compact $K\subseteq Y$.
We note that this result generalizes Theorem 4 by Kreitz-Weihrauch as for countable subbases $B:{\mathbb{N}}\to{\mathcal{O}}(X)$ the compact subsets $K\subseteq{\mathbb{N}}$ are exactly the finite subsets and hence the topology generated by $X$ and $\bigcap_{n\in K}B_{n}$ for compact $K\subseteq{\mathbb{N}}$ is exactly the same topology as the topology generated by the subbase $B$ itself.
[…]
Proposition 8 (Scott convergence).
Let $X$ be a sequential topological space. Let $(U_{n})$ be a sequence in ${\mathcal{O}}(X)$ and $U\in{\mathcal{O}}(X)$. Then the following are equivalent:
  1. (1)
    $U_{n}\to U$ with respect to the Scott topology.
  2. (2)
    $U_{n}\to U$ with respect to the topology, which is generated by the base of sets ${\mathcal{F}}_{K}$ over all compact $K\subseteq X$.
  3. (3)
    $U_{n}\to U$ with respect to the topology, which is generated by the subbase of sets ${\mathcal{F}}_{\{x,x_{n}:n\in{\mathbb{N}}\}}$ over all sequences $(x_{n})$ and $x$ in $X$ such that $x_{n}\to x$.
  4. (4)
    $U\subseteq\bigcup_{k\in{\mathbb{N}}}\left(\bigcap_{n\geq k}U_{n}\right)^{\circ}$.
Here $A^{\circ}$ denotes the interior of the set $A$. As a consequence of this result and Theorem 2 we obtain the following characterization of the topology of the space ${\mathcal{O}}(X)$.
Corollary 9 (Schröder 2002).
Let $X$ be a represented space. Then ${\mathcal{O}}(X)$ is endowed with the Scott topology that is the sequentialization of the compact-open topology and the standard representation of ${\mathcal{O}}(X)$ is admissible with respect to both topologies.
Now we are actually prepared to prove Theorem 7.
Proof of Theorem 7.
Since $(B_{y})_{y\in Y}$ is a presubbase of $X$, the map $B^{\mathsf{{T}}}:X\to{\mathcal{O}}(Y)$ is well-defined and injective. We consider the represented space $(X,\delta^{B})$ and we prove that it is a computable Kolmogorov space. We represent ${\mathcal{O}}(X)$ and ${\mathcal{O}}(Y)$ in the usual way, which means that they are endowed with the respective Scott topologies by Corollary 9. The definition $\delta^{B}=(B^{\mathsf{{T}}})^{-1}\circ\delta_{{\mathcal{O}}(Y)}$ ensures that $B^{\mathsf{{T}}}:X\to{\mathcal{O}}(Y)$ becomes a computable embedding with respect to $\delta^{B}$ as a representation of $X$. This implies that $B:Y\to{\mathcal{O}}(X)$ is computable too. We need to prove that also ${\mathcal{U}}:X\to{\mathcal{O}}{\mathcal{O}}(X)$ is a computable embedding. Given ${\mathcal{U}}_{x}=\{U\in{\mathcal{O}}(X):x\in U\}\in{\mathcal{O}}{\mathcal{O}}(X)$ for some $x\in X$, we can compute
$B^{\mathsf{{T}}}_{x}=\{y\in Y:x\in B_{y}\}=B^{-1}({\mathcal{U}}_{x})\in{\mathcal{O}}(Y)$
since $B$ is computable. And since $B^{\mathsf{{T}}}$ is a computable embedding, we can compute $x\in X$ from this set. Hence, also ${\mathcal{U}}:X\to{\mathcal{O}}{\mathcal{O}}(X)$ is a computable embedding and $(X,\delta^{B})$ is a computable Kolmogorov space.
In particular, $\delta^{B}$ is admissible with respect to its final topology ${\mathcal{O}}(X)$ by Proposition 3. We need to prove that it is also admissible with respect to the topology $\tau$. We already know by Corollary 9 that ${\mathcal{O}}(Y)$ is admissibly represented with respect to the compact-open topology generated by the sets ${\mathcal{F}}_{K}=\{U\in{\mathcal{O}}(Y):K\subseteq U\}$ over all compact $K\subseteq Y$. Now we note that
$\bigcap_{y\in K}B_{y}=\{x\in X:K\subseteq B_{x}^{\mathsf{{T}}}\}=(B^{\mathsf{{T}}})^{-1}({\mathcal{F}}_{K})$
and $(B^{\mathsf{{T}}})^{-1}({\mathcal{O}}(Y))=X$. Hence, $\tau$ is the initial topology of $B^{\mathsf{{T}}}$ with respect to the compact-open topology on ${\mathcal{O}}(Y)$, and hence $\delta^{B}=(B^{\mathsf{{T}}})^{-1}\circ\delta_{{\mathcal{O}}(Y)}$ is admissible with respect to $\tau$ by [Sch02b, Section 4.2] (see also [Sch02a, Proposition 4.1.4]). ∎
For simplicity we denote convergent sequences $(x_{n})_{n\in{\mathbb{N}}}$ that converge to $x_{\infty}$ as $(x_{n})_{n\in{\mathbb{N}}_{\infty}}$ with $\mathbb{N}$$_{\infty}$$\mathbb{N}$$\cup\{\infty\}$. By Proposition 8 we could alternatively also use the subbase sets ${\mathcal{F}}_{K}$ with $K=\{y_{n}:n\in{\mathbb{N}}_{\infty}\}$ over all convergent sequences $(y_{n})_{n\in{\mathbb{N}}_{\infty}}$ in $Y$ in the previous proof. Hence, we obtain the following.
Remark 10.
The statement of Theorem 7 remains true if we replace the intersections $\bigcap_{y\in K}B_{y}$ by intersections $\bigcap_{n\in{\mathbb{N}}_{\infty}}B_{y_{n}}$ over all convergent sequences $(y_{n})_{n\in{\mathbb{N}}_{\infty}}$ in $Y$.
General so called $Y$-indexed bases of type $B:Y\to{\mathcal{O}}(X)$ have also been independently considered in [dBSS16, Definition 6.1]. The remark above relates our results regarding presubbases to the so-called sequentially YY–indexed generating systems considered in [dBSS16, Definition 7.1].
When we have a presubbase $B$ of a represented space $X$, then we have to deal in general with at least four different topologies, all generated by $B$ in different ways. We illustrate these in Figure 1. By $\bigcap_{\mathcal{K}}B$ we denote the subbase described in Theorem 7 and by $\bigcap_{\infty}B$ we denote the subbase generated over all convergent sequences, as described in Remark 10.
topology generated
by $B$ as a subbase
$\subseteq\,\,$ topology generated by $\bigcap_{\infty}B$ as a subbase $\subseteq\,\,$ topology generated by $\bigcap_{\mathcal{K}}B$ as a subbase $\subseteq\,\,$ final topology of the representation $\delta^{B}$
Figure 1. Topologies associated to a presubbase $B$ of a represented space.
The fourth (right-hand side) topology in Figure 1 is always the sequentialization of the second and of the third one. These are the statements of Remark 10 and Theorem 7, respectively. In the case of countable presubbases $B:{\mathbb{N}}\to{\mathcal{O}}(X)$ all four topologies coincide. We introduce a name for presubbases for which the first and the second topologies coincide.
Definition 11 (Convergent intersection property).
We say that a presubbase $B:Y\to{\mathcal{O}}(X)$ of a represented space $X$ satisfies the convergent intersection property if for every convergent sequence $(y_{n})_{n\in{\mathbb{N}}_{\infty}}$ in $Y$ there exists a finite set $F\subseteq Y$ with $\bigcap_{n\in{\mathbb{N}}_{\infty}}B_{y_{n}}=\bigcap_{y\in F}B_{y}$.
For presubbases $B$ with the convergent intersection property all four topologies in Figure 1 have the fourth one as sequentialization.
Corollary 12 (Convergent intersection property).
Let $(B_{y})_{y\in Y}$ be a presubbase of a set $X$ that satisfies the convergent intersection property. Then $\delta^{B}$ is admissible with respect to the topology $\tau$ on $X$ that is generated by $B$.
3. Computable Presubbases, Prebases and Bases
In this section we want to develop a theory of computable bases $B:Y\to{\mathcal{O}}(X)$ along the lines of the theory of computable bases for second-countable spaces [BR25]. Bases with more complicated index sets have already been studied in other contexts, see, e.g., [dBSS16]. Firstly, the proof of Theorem 7 already indicates how we can define computable presubbases.
Definition 13 (Computable presubbase).
Let $X$ and $Y$ be represented spaces. Then $B:Y\to{\mathcal{O}}(X)$ is called a computable presubbase of $X$ if the transpose
$B^{\mathsf{{T}}}:X\to{\mathcal{O}}(Y),x\mapsto\{y\in Y:x\in B_{y}\}$
is a well-defined computable embedding.
We note that $B^{\mathsf{{T}}}$ is well-defined and computable if and only if $B$ is so. Obviously, a presubbase $B$ of $X$ is a computable presubbase of $X$ if and only if the representation of $X$ is computably equivalent to $\delta^{B}$. This observation together with Theorem 7 yield the following corollary.
Corollary 14 (Presubbases).
If $X$ is a represented space with a computable presubbase $B:Y\to{\mathcal{O}}(X)$, then $X$ is a computable Kolmogorov space and ${\mathcal{O}}(X)={\mathrm{seq}}(\tau)$ for the topology $\tau$ that is generated by the base sets $X$ and $\bigcap_{y\in K}B_{y}$ for every compact $K\subseteq Y$.
In analogy to the countable case we can now define the concept of a computable base. In the second-countable case we demand that every finite intersection of subbase sets should be computably representable as a countable union of subbase sets. If we replace “finite” by “compact” and “countable” by “overt”, then we obtain the general concept of a computable base. By ${\mathcal{K}}_{-}(Y)$ we denote the set of saturated compact subsets $K\subseteq Y$ represented via continuous maps $\forall_{K}:{\mathcal{O}}(X)\to{\mathbb{S}}$ and by ${\mathcal{A}}_{+}(Y)$ we denote the space of closed subsets $A\subseteq Y$ represented via continuous maps $\exists_{A}:{\mathcal{O}}(X)\to{\mathbb{S}}$. Here $\forall_{K}$ is just the characteristic function of ${\mathcal{F}}_{K}$ and $\exists_{A}$ is the characteristic function of ${\mathcal{T}}_{A}:=\{U\in{\mathcal{O}}(X):A\cap U\not$$\emptyset\}$ (see [Pau16, Sch21]). Using this terminology we obtain the following straightforward definition.
Definition 15 (Computable prebase).
Let $X$ a represented space with a computable presubbase $B:Y\to{\mathcal{O}}(X)$. Then $B$ is called a computable prebase of $X$, if there is a computable $R:{\mathcal{K}}_{-}(Y)\rightrightarrows{\mathcal{A}}_{+}(Y)$ such that
$\bigcap_{y\in K}B_{y}=\bigcup_{y\in A}B_{y}$
for every $K\in{\mathcal{K}}_{-}(Y)$ and $A\in R(K)$ and $X=\bigcup_{y\in Y}B_{y}$. We call $B$ a computable base of $X$ if $B$ is actually a base of $X$
Our definition of a computable base is a computable analogue of the definition of a YY–base provided in [dBSS16, Definition 6.1]. However, we note that the more important concept for us is that of a computable prebase, which is related to the sequential bases in [dBSS16, Definition 7.1].
From a purely topological perspective, computable prebases have in particular the property that compact intersections are open in the topology generated by the prebase. Hence, by Corollary 14 computable prebases characterize the topology of their spaces up to sequentialization.
Corollary 16 (Computable prebases).
If $X$ is a represented space that has a computable prebase that generates a topology $\tau$, then ${\mathcal{O}}(X)={\mathrm{seq}}(\tau)$.
In the following we will need a couple of computable operations for the spaces ${\mathcal{O}}(X)$, ${\mathcal{A}}_{+}(X)$ and ${\mathcal{K}}_{-}(X)$, respectively. Most of these operations have been considered before (see [Pau16]).
Proposition 17 (Computable operations on hyperspaces).
Let $X$ be a represented space and $Y$ a computable Kolmogorov space. Then the following hold:
  1. (1)
    ${\mathcal{U}}:X\to{\mathcal{O}}{\mathcal{O}}(X)$ is computable (neighborhood map)
  2. (2)
    ${\mathrm{inj}}:X\to{\mathcal{A}}_{+}(X),x\mapsto\overline{\{x\}}$ is computable (closed injection)
  3. (3)
    ${\mathrm{inj}}:X\to{\mathcal{K}}_{-}(X),x\mapsto{\mathrm{sat}}\{x\}$ is computable (compact injection)
  4. (4)
    ${\mathcal{K}}_{-}:{\mathcal{C}}(X,Y)\to{\mathcal{C}}({\mathcal{K}}_{-}(X),{\mathcal{K}}_{-}(Y)),f\mapsto(K\mapsto{\mathrm{sat}}f(K))$ (compact images)
  5. (5)
    ${\mathcal{A}}_{+}:{\mathcal{C}}(X,Y)\to{\mathcal{C}}({\mathcal{A}}_{+}(X),{\mathcal{A}}_{+}(Y)),f\mapsto(A\mapsto\overline{f(A)})$ (closed images)
  6. (6)
    $\sec:X\times{\mathcal{O}}(X\times Y)\to{\mathcal{O}}(Y),(x,U)\mapsto\{y\in Y:(x,y)\in U\}$ is computable (section)
  7. (7)
    $\times:{\mathcal{O}}(X)\times{\mathcal{O}}(Y)\to{\mathcal{O}}(X\times Y),(V,U)\mapsto V\times U$ is computable (product)
  8. (8)
    $\times:{\mathcal{A}}_{+}(X)\times{\mathcal{A}}_{+}(Y)\to{\mathcal{A}}_{+}(X\times Y),(A,B)\mapsto A\times B$ is computable (product)
  9. (9)
    $\bigcup:{\mathcal{A}}_{+}{\mathcal{O}}(X)\to{\mathcal{O}}(X),{\mathcal{A}}\mapsto\bigcup{\mathcal{A}}$ is computable (overt union of open)
  10. (10)
    $\bigcap:{\mathcal{K}}_{-}{\mathcal{O}}(X)\to{\mathcal{O}}(X),{\mathcal{K}}\mapsto\bigcap{\mathcal{K}}$ is computable (compact intersection of open)
  11. (11)
    $\bigcup:{\mathcal{K}}_{-}{\mathcal{K}}_{-}(X)\to{\mathcal{K}}_{-}(X),{\mathcal{K}}\mapsto{\mathrm{sat}}\bigcup{\mathcal{K}}$ is computable (compact union of compact)
  12. (12)
    ${\mathcal{F}}:{\mathcal{K}}_{-}(X)\to{\mathcal{O}}{\mathcal{O}}(X),K\mapsto{\mathcal{F}}_{K}$ is a computable embedding (filter)
  13. (13)
    ${\mathcal{T}}:{\mathcal{A}}_{+}(X)\to{\mathcal{O}}{\mathcal{O}}(X),A\mapsto{\mathcal{T}}_{A}$ is a computable embedding (trace)
  14. (14)
    $\Box:{\mathcal{O}}(X)\to{\mathcal{O}}{\mathcal{K}}_{-}(X),U\mapsto\{K\in{\mathcal{K}}_{-}(X):K\subseteq U\}$ is a computable embedding (box)
  15. (15)
    ${\mathcal{C}}{\mathcal{O}}:{\mathcal{C}}(X,Y)\to{\mathcal{O}}({\mathcal{K}}_{-}(X)\times{\mathcal{O}}(Y)),f\mapsto\{(K,U)\in{\mathcal{K}}_{-}(X)\times{\mathcal{O}}(Y):f(K)\subseteq U\}$
    is a computable embedding (compact-open)
For arbitrary sets ${\mathcal{A}}\subseteq{\mathcal{O}}(X)$ we have $\bigcup{\mathcal{A}}=\bigcup\overline{{\mathcal{A}}}$ and for arbitrary compact sets ${\mathcal{K}}\subseteq{\mathcal{O}}(X)$ we have $\bigcap{\mathcal{K}}=\bigcap{\mathrm{sat}}({\mathcal{K}})$.
Proof.
The results (1)–(10) can all be found in [Pau16], see Propositions 4.2, 5.5, 7.4, Section 9 and Corollaries 10.2 and 10.4 therein. We briefly discuss (11)–(15), which are all easy to see.
(12) We have $U\in{\mathcal{F}}_{K}\iff K\subseteq U$ and hence $\chi_{{\mathcal{F}}_{K}}(U)=\forall_{K}(U)$.
(13) We have $U\cap{\mathcal{T}}_{A}\iff A\cap U\not$$\emptyset$ and hence $\chi_{{\mathcal{T}}_{A}}(U)=\exists_{A}(U)$.
(14) We have $K\in\Box(U)\iff K\subseteq U\iff U\in{\mathcal{F}}_{K}$. Hence computability of $\mathcal{F}$} in (9) implies computability of $\Box$. We also have $x\in U\iff{\mathrm{sat}}\{x\}\subseteq U\iff{\mathrm{sat}}\{x\}\in\Box(U)$. Hence $\Box$ is a computable embedding as ${\mathrm{inj}}:X\to{\mathcal{K}}_{-}(X)$ is computable by (3).
(11) We have ${\mathrm{sat}}\bigcup{\mathcal{K}}\subseteq{\mathcal{U}}\iff\bigcup{\mathcal{K}}\subseteq U\iff(\forall K\in{\mathcal{K}})\;K\in\Box U\iff\forall_{\mathcal{K}}(\Box U)=1$. Hence, (11) follows from (14).
(15) Since $f(K)\subseteq U\iff{\mathrm{sat}}f(K)\subseteq U$, computability of ${\mathcal{C}}{\mathcal{O}}$ follows from (1) and (4). If $Y$ is a computable Kolmogorov space, then ${\mathcal{U}}:Y\to{\mathcal{O}}{\mathcal{O}}(Y)$ is a computable embedding. Since $U\in{\mathcal{U}}_{f(x)}\iff f(x)\in U\iff{\mathrm{sat}}f\{x\}\subseteq U$, it follows that ${\mathcal{C}}{\mathcal{O}}$ is a computable embedding in this case too.
For arbitrary ${\mathcal{A}}\subseteq{\mathcal{O}}(X)$ and compact ${\mathcal{K}}\subseteq{\mathcal{O}}(X)$ we have
  • $x\in\bigcup\overline{{\mathcal{A}}}\iff\overline{{\mathcal{A}}}\cap{\mathcal{U}}_{x}\not=\emptyset\iff{\mathcal{A}}\cap{\mathcal{U}}_{x}\not=\emptyset\iff x\in\bigcup{\mathcal{A}}$ and
  • $x\in\bigcap{\mathrm{sat}}{\mathcal{K}}\iff{\mathrm{sat}}{\mathcal{K}}\subseteq{\mathcal{U}}_{x}\iff{\mathcal{K}}\subseteq{\mathcal{U}}_{x}\iff x\in\bigcup{\mathcal{K}}$,
which proves the additional remark. ∎
Often, we can establish that a computable presubbase is even a computable prebase because it is already closed under compact intersections.
Lemma 18 (Closure under compact intersections).
Let $X$ be a represented space with a computable presubbase $B:Y\to{\mathcal{O}}(X)$. If there is a computable problem $R:{\mathcal{K}}_{-}(Y)\rightrightarrows Y$} such that $\bigcap_{y\in K}B_{y}=B_{z}$ for every $z\in R(K)$, then $B$ is a computable prebase.
Proof.
This follows as ${\mathrm{inj}}:Y\to{\mathcal{A}}_{+}(Y),z\mapsto\overline{\{z\}}$ is computable and $\bigcup_{y\in\overline{\{z\}}}B_{y}=B_{z}$ by Proposition 17. ∎
Every computable presubbase can be converted into a computable prebase just by taking the closure under compact intersections.
Proposition 19 (Computable prebases).
Let $X$ be a represented space with a computable presubbase $B:Y\to{\mathcal{O}}(X)$. Then
$\cap_{\mathcal{K}}B:{\mathcal{K}}_{-}(Y)\to{\mathcal{O}}(X),K\mapsto\bigcap_{y\in K}B_{y},$
is a computable prebase with the understanding that the empty intersection is $X$.
Proof.
We use Proposition 17. Firstly, we note that $\cap_{\mathcal{K}}B(\emptyset)=X$. Next, we need to prove that $\cap_{\mathcal{K}}B$ is a computable presubbase, i.e., that
$(\cap_{\mathcal{K}}B)^{\mathsf{{T}}}:X\to{\mathcal{O}}{\mathcal{K}}_{-}(Y),x\mapsto\{K\in{\mathcal{K}}_{-}(Y):x\in\cap_{y\in K}B_{y}\}$
is well-defined and a computable embedding. Since $B^{\mathsf{{T}}}$ is a computable embedding, it follows that $B$ is computable and hence so is $\cap_{\mathcal{K}}B$. Hence $(\cap_{\mathcal{K}}B)^{\mathsf{{T}}}$ also is well-defined and computable. Now, given $(\cap_{\mathcal{K}}B)^{\mathsf{{T}}}_{x}$, we can compute $B^{\mathsf{{T}}}_{x}=\{y\in Y:x\in B_{y}\}\in{\mathcal{O}}(Y)$, as ${\mathrm{inj}}:Y\to{\mathcal{K}}_{-}(Y),y\mapsto{\mathrm{sat}}\{y\}$ is computable and the saturation is irrelevant as $\cap_{y\in K}B_{y}=\cap_{y\in{\mathrm{sat}}(K)}B_{y}$. That is $B^{\mathsf{{T}}}_{x}={\mathrm{inj}}^{-1}((\bigcap_{\mathcal{K}}B)^{\mathsf{{T}}}_{x})$. As $B^{\mathsf{{T}}}$ is a computable embedding, this implies that $(\cap_{\mathcal{K}}B)^{\mathsf{{T}}}$ is a computable embedding too.
Finally, we need to prove that $\cap_{\mathcal{K}}B$ is a computable prebase, i.e., that all compact intersections of it can be obtained computably as overt unions. To this end, we prove that $\cap_{\mathcal{K}}\cap_{\mathcal{K}}B:{\mathcal{K}}_{-}{\mathcal{K}}_{-}(Y)\to{\mathcal{O}}(X)$ is computable. Since $\bigcup:{\mathcal{K}}_{-}{\mathcal{K}}_{-}(Y)\to{\mathcal{K}}_{-}(Y),{\mathcal{K}}\mapsto{\mathrm{sat}}\bigcup{\mathcal{K}}$ is computable and
$\bigcap_{K\in{\mathcal{K}}}\bigcap_{y\in K}B_{y}=\bigcap_{y\in\bigcup{\mathcal{K}}}B_{y}=\bigcap_{y\in{\mathrm{sat}}(\bigcup{\mathcal{K}})}B_{y},$
the claim follows with Lemma 18. ∎
In light of the countable case [BR25] there is also a stronger notion of a computable base described by unions, which we call a Lacombe base. We call a map $f:X\to Y$ computably surjective if it is computable, surjective and has a multivalued computable inverse $F:Y\rightrightarrows X$.
Definition 20 (Lacombe base).
Let $X$ be a computable Kolmogorov space and let $Y$ be a represented space. Then a map $B:Y\to{\mathcal{O}}(X)$ is called a computable Lacombe base of $X$ if
$\bigcup:{\mathcal{A}}_{+}(Y)\to{\mathcal{O}}(X),A\mapsto\bigcup_{y\in Y}B_{y}$
is computably surjective.
Similar definitions of topological bases have been considered in a constructive setting before [BL12, Definition 6.3]. It is not too difficult to see and essentially a consequence of Proposition 17 that every computable Lacombe base is a computable base.
Proposition 21 (Computable Lacombe bases).
Let $X$ be a computable Kolmogorov space. Then every computable Lacombe base of $X$ is also a computable base.
Proof.
We use Proposition 17. If $\bigcup B:{\mathcal{A}}_{+}(Y)\to{\mathcal{O}}(X)$ is computable, then so is the map $B:Y\to{\mathcal{O}}(X)$. This is because ${\mathrm{inj}}:Y\to{\mathcal{A}}_{+}(Y),y\mapsto\overline{\{y\}}$ is computable and $\bigcup_{z\in\overline{\{y\}}}B_{z}=B_{y}$. We also have that $\bigcap:{\mathcal{K}}_{-}{\mathcal{O}}(X)\to{\mathcal{O}}(X)$ is computable and $\bigcup:{\mathcal{A}}_{+}(Y)\to{\mathcal{O}}(X)$ has a computable right inverse $S:{\mathcal{O}}(X)\rightrightarrows{\mathcal{A}}_{+}(Y)$. Hence, given a set $K\in{\mathcal{K}}_{-}(Y)$ we can compute ${\mathcal{K}}:={\mathrm{sat}}B(K)={\mathrm{sat}}\{B_{y}:y\in K\}\in{\mathcal{K}}_{-}{\mathcal{O}}(X)$ and hence $\bigcap{\mathcal{K}}=\bigcap_{y\in K}B_{y}\in{\mathcal{O}}(X)$. Using $S$ we obtain $A\in{\mathcal{A}}_{+}(Y)$ with $\bigcap_{y\in K}B_{y}=\bigcup_{y\in A}B_{y}$. Altogether this describes a computation of a problem $R:{\mathcal{K}}_{-}(Y)\rightrightarrows{\mathcal{A}}_{+}(Y)$ as it is required for a computable prebase. Because $\bigcup:{\mathcal{A}}_{+}(Y)\to{\mathcal{O}}(X)$ is surjective, we also have $\bigcup_{y\in Y}B_{y}=X$ and $B$ is actually a base of $X$. It remains to prove that $B:Y\to{\mathcal{O}}(X)$ is a computable presubbase, i.e., that the transpose $B^{\mathsf{{T}}}:X\to{\mathcal{O}}(Y)$ is injective and has a computable right inverse. If the set ${\mathcal{U}}=\{y\in Y:x\in B_{y}\}\in{\mathcal{O}}(Y)$ is given, then we can also compute ${\mathcal{U}}_{x}=\{U\in{\mathcal{O}}(X):x\in U\}\in{\mathcal{O}}{\mathcal{O}}(X)$. This is because given $U\in{\mathcal{O}}(X)$ we can compute $A\in{\mathcal{A}}_{+}(Y)$ with $U=\bigcup_{y\in A}B_{y}$ and hence $x\in U\iff(\exists y\in A)\;x\in B_{y}\iff(\exists y\in A)\;y\in{\mathcal{U}}$. But given ${\mathcal{U}}_{x}\in{\mathcal{O}}{\mathcal{O}}(X)$ we can compute $x\in X$, as $X$ is a computable Kolmogorov space. This completes the proof that $B:Y\to{\mathcal{O}}(X)$ is a computable base of $X$. ∎
In fact, every computable Kolmogorov space has a computable Lacombe base, namely simply the identity.
Proposition 22 (Identity as a base).
Let $X$ be a represented space. Then the identity ${\mathrm{id}}:{\mathcal{O}}(X)\to{\mathcal{O}}(X)$ is a computable Lacombe base/base/prebase/presubbase of $X$ if and only if $X$ is a computable Kolmogorov space.
Proof.
It follows from ${\mathrm{id}}^{\mathsf{{T}}}={\mathcal{U}}:X\to{\mathcal{O}}{\mathcal{O}}(X)$ that $\mathrm{id}$} is a computable prebase if and only if $X$ is a computable Kolmogorov space. Let $X$ now be a computable Kolmogorov space. We claim that $\mathrm{id}$} is even a computable Lacombe base. This follows from Proposition 17 since ${\mathrm{inj}}:{\mathcal{O}}(X)\to{\mathcal{A}}_{+}{\mathcal{O}}(X),U\mapsto\overline{\{U\}}$ is computable and $\bigcup\overline{\{U\}}=U$. ∎
Altogether, we obtain the following characterization of computable Kolmogorov spaces in terms of their bases.
Theorem 23 (Computable Kolmogorov spaces and bases).
Let $X$ be a represented space. Then the following are pairwise equivalent:
  1. (1)
    $X$ is a computable Kolmogorov space.
  2. (2)
    $X$ has a computable presubbase.
  3. (3)
    $X$ has a computable prebase.
  4. (4)
    $X$ has a computable base.
  5. (5)
    $X$ has a computable Lacombe base.
  6. (6)
    ${\mathrm{id}}:{\mathcal{O}}(X)\to{\mathcal{O}}(X)$ is a computable Lacombe base of $X$.
Proof.
[…]
For computably second-countable spaces every computable base is even a computable Lacombe base [BR25]. That is for $Y={\mathbb{N}}$ the inverse implication in Proposition 21 holds true as well. We do not know whether this holds in general for all index spaces $Y$.
Problem 24.
Is every computable base $B:Y\to{\mathcal{O}}(X)$ of a computable Kolmogorov space $X$ a computable Lacombe base? If not, characterize the index spaces $Y$ for which this holds!
By a represented ${\mathsf{{T}}}_{0}$ space we mean a represented space $X$ whose final topology is ${\mathsf{{T}}}_{0}$. We note that another consequence of the above results is that every represented ${\mathsf{{T}}}_{0}$ space $X$ can be converted into a computable Kolmogorov space. This is because ${\mathrm{id}}:{\mathcal{O}}(X)\to{\mathcal{O}}(X)$ is always a presubbase. In fact, this conversion preserves the topology and the equivalence class of its representation. For two representations $\delta_{1}$ of a set $X_{1}$ and $\delta_{2}$ of a set $X_{2}$ we write $\delta_{1}\leq\delta_{2}$ if $X_{1}\subseteq X_{2}$ and the representation $\delta_{1}$ can be computably reduced to $\delta_{2}$, i.e., if the identity ${\mathrm{id}}:X_{1}\to X_{2}$ is computable.
Theorem 25 (Represented spaces as computable Kolmogorov spaces).
Let $X$ be a represented space with a ${\mathsf{{T}}}_{0}$ topology ${\mathcal{O}}(X)$. Then we can endow $X$ with another representation that turns $X$ into a computable Kolmogorov space without changing the topology ${\mathcal{O}}(X)$ (and without changing the computable equivalence class of the representation of ${\mathcal{O}}(X)$).
Proof.
Let $(X,\delta)$ be a represented space with ${\mathsf{{T}}}_{0}$ topology ${\mathcal{O}}(X)$ and canonical representation $\delta^{\circ}$ of ${\mathcal{O}}(X)$ and $\delta^{\circ\circ}$ of ${\mathcal{O}}{\mathcal{O}}(X)$. Then ${\mathrm{id}}:{\mathcal{O}}(X)\to{\mathcal{O}}(X)$ is a presubbase of this space, as ${\mathrm{id}}^{\mathsf{{T}}}={\mathcal{U}}:X\to{\mathcal{O}}{\mathcal{O}}(X)$ is well-defined and injective. Hence, the presubbase representation $\delta^{\bullet}={\mathcal{U}}^{-1}\circ\delta^{\circ\circ}$ turns $X$ into a computable Kolmogorov space $X^{\bullet}$ by the Presubbase Theorem 7. Let $\delta^{\bullet\circ}$ denote the representation of the open sets ${\mathcal{O}}(X^{\bullet})$ induced by $\delta^{\bullet}$. We need to prove that ${\mathcal{O}}(X)={\mathcal{O}}(X^{\bullet})$ and that $\delta^{\circ}\equiv\delta^{\bullet\circ}$. Since $\mathcal{U}$} is computable, it is clear that $\delta\leq\delta^{\bullet}$ and hence $\delta^{\bullet\circ}\leq\delta^{\circ}$ and ${\mathcal{O}}(X^{\bullet})\subseteq{\mathcal{O}}(X)$. We need to prove the inverse reduction. If $x=\delta^{\bullet}(p)$ and $U=\delta^{\circ}(q)$ are given by $p,q\in{\mathbb{N}}^{\mathbb{N}}$, then we have $x\in U\iff U\in{\mathcal{U}}_{x}\iff\delta^{\circ}(q)\in\delta^{\circ\circ}(p)$ which can be confirmed with the help of $p,q\in{\mathbb{N}}^{\mathbb{N}}$. This shows $\delta^{\circ}\leq\delta^{\bullet\circ}$ and ${\mathcal{O}}(X)\subseteq{\mathcal{O}}(X^{\bullet})$. ∎
In fact, $\delta\mapsto\delta^{\bullet}$ is a closure operator in the lattice of representations of ${\mathsf{{T}}}_{0}$ spaces that was introduced and studied by Schröder [Sch02a]. We will get back to discussing this closure operator in the Epilogue.
4. Closure properties
In this section we provide a number of examples of computable presubbases and prebases and we demonstrate that these concepts can be used to derive interesting results. All these results are essentially well-known (mostly due to the work of Schröder [Sch02a]). The purpose here is not to claim originality, but to show how these results can be easily derived using the concepts of bases. We start with providing a number of examples that show that computable presubbases of hyper and function spaces occur very naturally.
Proposition 26 (Computable presubbases).
Let $X$ be a represented space and $Y$ a computable Kolmogorov space. The following are computable presubbases:
  1. (1)
    ${\mathcal{U}}:X\to{\mathcal{O}}{\mathcal{O}}(X),x\mapsto\{U\in{\mathcal{O}}(X):x\in U\}$.
  2. (2)
    ${\mathcal{F}}:{\mathcal{K}}_{-}(X)\to{\mathcal{O}}{\mathcal{O}}(X),K\mapsto\{U\in{\mathcal{O}}(X):K\subseteq U\}$.
  3. (3)
    $\Box:{\mathcal{O}}(X)\to{\mathcal{O}}{\mathcal{K}}_{-}(X),U\mapsto\{K\in{\mathcal{K}}_{-}(X):K\subseteq U\}$.
  4. (4)
    $\Diamond:{\mathcal{O}}(X)\to{\mathcal{O}}{\mathcal{A}}_{+}(X),U\mapsto\{A\in{\mathcal{A}}_{+}(X):A\cap U\not$$\emptyset\}$.
  5. (5)
    $\triangleright:{\mathcal{K}}_{-}(X)\times{\mathcal{O}}(Y)\to{\mathcal{O}}{\mathcal{C}}(X,Y),(K,U)\mapsto\{f\in{\mathcal{C}}(X,Y):f(K)\subseteq U\}$.
The maps $\mathcal{F}$} and $\Box$ are even computable prebases.
Proof.
We have ${\mathcal{U}}^{\mathsf{{T}}}={\mathrm{id}}_{{\mathcal{O}}(X)}$, $\Box^{\mathsf{{T}}}={\mathcal{F}}$, $\Diamond^{\mathsf{{T}}}={\mathcal{T}}$ and $\triangleright^{\mathsf{{T}}}={\mathcal{C}}{\mathcal{O}}$ where the maps $\mathcal{F}$}, $\mathcal{T}$}, ${\mathcal{C}}{\mathcal{O}}$ and $\Box$ are computable embeddings according to Proposition 17. Hence, $\mathcal{U}$}, $\mathcal{F}$}, $\Box$, $\Diamond$ and $\triangleright$ are all computable presubbases.
We still need to prove that $\mathcal{F}$} and $\Box$ are even computable prebases. Given ${\mathcal{K}}\in{\mathcal{K}}_{-}{\mathcal{K}}_{-}(X)$ we can compute ${\mathrm{sat}}\bigcup_{K\in{\mathcal{K}}}K\in{\mathcal{K}}_{-}(X)$ according to Proposition 17 and because $\bigcap_{K\in{\mathcal{K}}}{\mathcal{F}}_{K}={\mathcal{F}}_{{\mathrm{sat}}\bigcup_{K\in{\mathcal{K}}}K}$ it follows by Lemma 18 that $\mathcal{F}$} is a computable prebase. Likewise, given ${\mathcal{K}}\in{\mathcal{K}}_{-}{\mathcal{O}}(X)$ we can compute $\bigcap_{U\in{\mathcal{K}}}U\in{\mathcal{O}}(X)$ and because $\bigcap_{U\in{\mathcal{K}}}\Box(U)=\Box(\bigcap_{U\in{\mathcal{K}}}U)$ it follows by Lemma 18 that $\Box$ is a computable prebase. ∎
In particular, we can conclude that all the involved spaces ${\mathcal{O}}(X)$, ${\mathcal{K}}_{-}(X)$, ${\mathcal{A}}_{+}(X)$ and ${\mathcal{C}}(X,Y)$ are computable Kolmogorov spaces under the given conditions (see Corollary 31). The maps specified in Proposition 26 are all known as subbases of some well-known topologies on the respective spaces. We summarize the terminology in the table in Figure 2.
space subbase name of topology
${\mathcal{C}}(X,Y)$ $\triangleright$ compact-open topology
${\mathcal{A}}_{+}(X)$ $\Diamond$ lower Fell topology
${\mathcal{K}}_{-}(X)$ $\Box$ upper Vietoris topology
${\mathcal{O}}(X)$ $\mathcal{F}$} compact-open topology
${\mathcal{O}}(X)$ $\mathcal{U}$} point-open topology
Figure 2. Function and hyperspace topologies and their subbases.
Proposition 26 can also be used as an example to illustrate the different scopes of presubbases, prebases, and bases, respectively.
Example 27.
Let $X$ be some represented space.
  1. (1)
    The neighborhood map $\mathcal{U}$} is a computable presubbase of ${\mathcal{O}}(X)$ that generates the point-open topology on ${\mathcal{O}}(X)$.
  2. (2)
    The canonical computable prebase associated to $\mathcal{U}$} (see Proposition 19) is $\bigcap_{\mathcal{K}}{\mathcal{U}}={\mathcal{F}}$, i.e., the filter map $\mathcal{F}$} that generates the compact-open topology on ${\mathcal{O}}(X)$, whose sequentialization is hence the Scott topology on ${\mathcal{O}}(X)$.
  3. (3)
    The map $\mathcal{F}$} is a computable base of ${\mathcal{O}}(X)$ with respect to the Scott topology if and only if the space $X$ is consonant.
  4. (4)
    The map ${\mathrm{id}}_{{\mathcal{O}}{\mathcal{O}}(X)}$ is always a computable (Lacombe) base of ${\mathcal{O}}(X)$ with respect to the Scott topology.
These examples show that in the general case the concept of a computable presubbase and of a computable prebase might actually be more fruitful and interesting than the concept of a computable base. We have lots of natural presubbases and prebases that reveal useful information about our spaces. The presubbase $\mathcal{U}$} does in general not satisfy the convergent intersection property, while the other two presubbases mentioned in Example 27 do.
With the following result we provide a few natural constructions of prebases, where we exploit that the index spaces are overt. We recall that a represented space $X$ is called overt, if $X$ is a computable point in ${\mathcal{A}}_{+}(X)$. Overt spaces are exactly those spaces that allow computable projections on open sets.
Proposition 28 (Overt space).
A represented spaces $X$ is overt if and only if the projection
${\mathrm{pr}}_{Y}:{\mathcal{O}}(Y\times X)\to{\mathcal{O}}(Y),U\mapsto\{y\in Y:(\exists x\in X)\;(y,x)\in U\}$
is computable for every represented space $Y$
Proof.
We use the fact that sections $U_{y}:=\{x\in X:(y,x)\in U\}$ and products are computable by Proposition 17.
[…]
Now we can prove the following closure properties for bases $B:R\to{\mathcal{O}}(X)$ with overt index spaces $R$. We note that ${\mathcal{O}}(X)$ is always overt, hence every computable Kolmogorov space $X$ has a computable base ${\mathrm{id}}_{{\mathcal{O}}(X)}:{\mathcal{O}}(X)\to{\mathcal{O}}(X)$ with an overt index space. By $X\sqcup Y$ we denote the coproduct of two represented spaces $X$ and $Y$, which is the set $X\sqcup Y=(\{0\}\times X)\cup(\{1\}\times Y)$ with its canonical representation and by $X\sqcap Y$ we denote the meet $X\sqcap Y=X\cap Y$, which is represented with names that are pairs, where one component represents the point as a point in $X$ and the second component the same point as a point in $Y$. We can identify the space $X\sqcap Y$ with the subspace $\{(x,y)\in X\times Y:x=y\}$ of $X\times Y$. Hence, the results on $X\sqcap Y$ could also be derived from the results on products and subspaces.
Proposition 29 (Constructions of overt prebases).
Let $X$ and $Y$ be represented spaces and let $Z\subseteq X$. Let $S$ and $R$ be overt represented spaces. If $B_{X}:R\to{\mathcal{O}}(X)$ and $B_{Y}:S\to{\mathcal{O}}(Y)$ are computable presubbases (prebases), then so are:
  1. (1)
    $B_{X\times Y}:R\times S\to{\mathcal{O}}(X\times Y),(r,s)\mapsto B_{X}(r)\times B_{Y}(s)$.
  2. (2)
    $B_{Y^{\mathbb{N}}}:S^{*}\to{\mathcal{O}}(Y^{\mathbb{N}}),(s_{1},...,s_{n})\mapsto B_{Y}(s_{1})\times...\times B_{Y}(s_{n})\times Y^{\mathbb{N}}$.
  3. (3)
    $B_{Z}:R\to{\mathcal{O}}(Z),r\mapsto B_{X}(r)\cap Z$.
  4. (4)
    $B_{X\sqcup Y}:R\sqcup S\to{\mathcal{O}}(X\sqcup Y),t\mapsto\left\{\begin{array}[]{ll}B_{X}(t)&\mbox{if $t\in R$}\\ B_{Y}(t)&\mbox{if $t\in S$}\end{array}\right.$.
  5. (5)
    $B_{X\sqcap Y}:R\times S\to{\mathcal{O}}(X\sqcap Y),(r,s)\mapsto B_{X}(r)\cap B_{Y}(s)$.
If $B_{X}$ and $B_{Y}$ are even computable bases, then so is $B_{X\sqcup Y}$. Overtness of $R$ is not needed for (3).
Proof.
Let $B_{X}$ and $B_{Y}$ be computable presubbases. We obtain the transposes
  1. (1)
    $B_{X\times Y}^{\mathsf{{T}}}:X\times Y\to{\mathcal{O}}(R\times S),(x,y)\mapsto B_{X}^{\mathsf{{T}}}(x)\times B_{Y}^{\mathsf{{T}}}(y)$.
  2. (2)
    $B_{Y^{\mathbb{N}}}^{\mathsf{{T}}}:Y^{\mathbb{N}}\to{\mathcal{O}}(S^{*}),(y_{n})_{n\in{\mathbb{N}}}\mapsto\bigsqcup_{n\in{\mathbb{N}}}\left(B_{Y}^{\mathsf{{T}}}(y_{0})\times...\times B_{Y}^{\mathsf{{T}}}(y_{n-1})\right)$.
  3. (3)
    $B_{Z}^{\mathsf{{T}}}:Z\to{\mathcal{O}}(R),z\mapsto B_{X}^{\mathsf{{T}}}(z)$.
  4. (4)
    $B_{X\sqcup Y}^{\mathsf{{T}}}:X\sqcup Y\to{\mathcal{O}}(R\sqcup S),z\mapsto\left\{\begin{array}[]{ll}B_{X}^{\mathsf{{T}}}(z)&\mbox{if $z\in X$}\\ B_{Y}^{\mathsf{{T}}}(z)&\mbox{if $z\in Y$}\end{array}\right.$.
  5. (5)
    $B_{X\sqcap Y}^{\mathsf{{T}}}:X\sqcap Y\to{\mathcal{O}}(R\times S),z\mapsto B_{X}^{\mathsf{{T}}}(z)\times B_{Y}^{\mathsf{{T}}}(z)$.
It is easy to see that all these maps are computable and we are going to show that they are computable embeddings. Hence the maps listed in the proposition are all computable presubbases.
(1) and (2) We prove this in detail for the countable case of $B_{Y^{\mathbb{N}}}$ in (2) and we leave the finite case to the reader. Given $U:=\bigsqcup_{n\in{\mathbb{N}}}\left(B_{Y}^{\mathsf{{T}}}(y_{0})\times...\times B_{Y}^{\mathsf{{T}}}(y_{n-1})\right)\in{\mathcal{O}}(S^{*})$ for some $(y_{n})_{n\in{\mathbb{N}}}\in Y^{\mathbb{N}}$ and $n\in{\mathbb{N}}$, we can compute the sections
$V:=\sec_{n+1}(U)=B_{Y}^{\mathsf{{T}}}(y_{0})\times...\times B_{Y}^{\mathsf{{T}}}(y_{n})\in{\mathcal{O}}(S^{n+1})$
and since $S$ is overt, and hence $S^{n}$ too, we can then compute the projection $W:=B_{Y}^{\mathsf{{T}}}(y_{n})\in{\mathcal{O}}(S)$ by Proposition 28 (which also holds uniformly in $n$). From $W$ we can reconstruct $y_{n}$, as $B_{Y}^{\mathsf{{T}}}$ is a computable embedding.
(3) It immediately follows that $B_{Z}^{\mathsf{{T}}}$ is a computable embedding because $B_{X}^{\mathsf{{T}}}$ is one. Overtness of $R$ is not required here.
(4) Given $U\in{\mathcal{O}}(R\sqcup S)$, where $R\sqcup S=(\{0\}\times R)\cup(\{1\}\times S)$, we can compute the sections $U_{0}\in{\mathcal{O}}(R)$ and $U_{1}\in{\mathcal{O}}(S)$. If $U$ is in the image of $B_{X\sqcup Y}^{\mathsf{{T}}}$ exactly one of $U_{0}$ and $U_{1}$ is non-empty and since $R$ and $S$ are overt, we can find out which one it is. If, for instance, $U_{0}\not$$\emptyset$, then $U_{0}=B_{X}^{\mathsf{{T}}}(x)$ and since $B_{X}^{\mathsf{{T}}}$ is a computable embedding, we can reconstruct $x\in X$ from $U_{0}$. Likewise, we can reconstruct $y\in Y$ in the case of $U_{1}\not$$\emptyset$.
(5) Given $U=B_{X}^{\mathsf{{T}}}(z)\times B_{Y}^{\mathsf{{T}}}(z)\in{\mathcal{O}}(R\times S)$ we can compute the projections $B_{X}^{\mathsf{{T}}}(z)$ and $B_{Y}^{\mathsf{{T}}}(z)$ by Proposition 29 and hence we can reconstruct $z\in X$ as well as $z\in Y$, as $B_{X}^{\mathsf{{T}}}$ and $B_{Y}^{\mathsf{{T}}}$ are computable embeddings. This yields $z\in X\sqcap Y$.
Let now $B_{X}$ and $B_{Y}$ even be computable prebases. We claim that in this case all the maps listed in the proposition are computable prebases too. To this end, we need to prove that compact intersections of the given map can be computably obtained as overt unions.
(1) and (2) We prove the claim for $B_{Y^{\mathbb{N}}}$. Given $K\in{\mathcal{K}}_{-}(S^{*})$ we can compute the projection $K_{0}:={\mathrm{pr}}_{0}(K)\in{\mathcal{K}}_{-}({\mathbb{N}})$ onto the natural number component and we can compute some upper bound $m\in{\mathbb{N}}$ on $\max(K_{0})$. Hence, we can compute also the projections $K_{i}:={\mathrm{sat}}({\mathrm{pr}}_{i}(K))\in{\mathcal{K}}_{-}(S)$ on the components i=1,,mi=1,...,m of $K$ by Proposition 17 (which holds uniformly in $i$). For each i=1,,mi=1,...,m we can compute some set $A_{i}\in{\mathcal{A}}_{+}(S)$ with $\bigcap_{s\in K_{i}}B_{Y}(s)=\bigcup_{s\in A_{i}}B_{Y}(s)$. By Proposition 17 (which holds analogously for finite products) we can hence compute $A:=A_{1}\times...\times A_{m}\in{\mathcal{A}}_{+}(S^{*})$ and we obtain for all $n\in{\mathbb{N}}$
$\bigcap_{(s_{1},...,s_{n})\in K}B_{Y^{\mathbb{N}}}(s_{1},...,s_{n})$ $=$ $\bigcap_{(s_{1},...,s_{n})\in K}(B_{Y}(s_{1})\times...\times B_{Y}(s_{n})\times Y^{\mathbb{N}})$
$=$ $\left(\bigcap_{s_{1}\in K_{1}}B_{Y}(s_{1})\right)\times...\times\left(\bigcap_{s_{m}\in K_{m}}B_{Y}(s_{m})\right)\times Y^{\mathbb{N}}$
$=$ $\left(\bigcup_{s_{1}\in A_{1}}B_{Y}(s_{1})\right)\times...\times\left(\bigcup_{s_{m}\in A_{m}}B_{Y}(s_{m})\right)\times Y^{\mathbb{N}}$
$=$ $\bigcup_{(s_{1},...,s_{m})\in A}B_{Y^{\mathbb{N}}}(s_{1},...,s_{m}).$
We note that the second and fourth equalities hold as the conditions on the $s_{i}$ are independent and for the second equality we also use that the saturated sets $K_{i}={\mathrm{sat}}({\mathrm{pr}}_{i}(K))$ can be replaced by the unsaturated sets ${\mathrm{pr}}_{i}(K)$ in the intersections by Proposition 17.
(3) This is obvious as the intersection with $Z$ can be associated to a compact intersection as well as to an overt union.
(4) It is easy to see that given $K\in{\mathcal{K}}_{-}(R\sqcup S)$ we can compute the sections $K_{0}\in{\mathcal{K}}_{-}(R)$ and $K_{1}\in{\mathcal{K}}_{-}(S)$. This is because $K_{0}\subseteq U\iff K\subseteq U\sqcup Y$ and $K_{1}\subseteq V\iff K\subseteq X\sqcup V$ for $U\in{\mathcal{O}}(R)$ and $V\in{\mathcal{O}}(S)$. With these sections we obtain
$\bigcap_{t\in K}B_{X\sqcup Y}(t)$ $=$ $\left(\bigcap_{r\in K_{0}}B_{X}(t)\right)\sqcup\left(\bigcap_{s\in K_{1}}B_{Y}(s)\right)$
$=$ $\left(\bigcup_{r\in A_{0}}B_{X}(t)\right)\sqcup\left(\bigcup_{s\in A_{1}}B_{Y}(s)\right)=\bigcup_{t\in A_{0}\sqcup A_{1}}B_{X\sqcup Y}(t)$
where $A_{0}\in{\mathcal{A}}_{+}(R)$ and $A_{1}\in{\mathcal{A}}_{+}(S)$ are sets that we can compute from $K_{0}$ and $K_{1}$ by assumption. Hence we can also compute $A:=A_{0}\sqcup A_{1}\in{\mathcal{A}}_{+}(R\sqcup S)$ because $A\cap U\not$$\emptyset$ for $U\in{\mathcal{O}}(R\sqcup S)$ if and only if $A_{0}\cap U_{0}\not$$\emptyset$ or $A_{1}\cap U_{1}\not$$\emptyset$.
(5) If $K\in{\mathcal{K}}_{-}(R\times S)$, then we can compute the projections $K_{1}:={\mathrm{sat}}{\mathrm{pr}}_{1}(K)\in{\mathcal{K}}_{-}(R)$ and $K_{2}:={\mathrm{sat}}{\mathrm{pr}}_{2}(K)\in{\mathcal{K}}_{-}(S)$. We obtain
$\bigcap_{(r,s)\in K}B_{X\sqcap Y}(r,s)$ $=$ $\bigcap_{r\in K_{1}}B_{X}(r)\cap\bigcap_{s\in K_{2}}B_{Y}(s)$
$=$ $\left(\bigcup_{r\in A_{1}}B_{X}(r)\right)\cap\left(\bigcup_{s\in A_{2}}B_{Y}(s)\right)=\bigcup_{(r,s)\in A_{1}\times A_{2}}B_{X}(r)\cap B_{Y}(s)$
and by assumption we can compute corresponding sets $A_{1}\in{\mathcal{A}}_{+}(R)$ and $A_{2}\in{\mathcal{A}}_{+}(S)$.
Let now $B_{X}$ and $B_{Y}$ even be computable bases. That $B_{X\sqcup Y}$ is a base follows from ${\mathcal{O}}(X\sqcup Y)={\mathcal{O}}(X)\sqcup{\mathcal{O}}(Y)$. ∎
Using Proposition 29 and the computable base ${\mathrm{id}}_{{\mathcal{O}}(X)}:{\mathcal{O}}(X)\to{\mathcal{O}}(X)$ we even obtain specific computable prebases for spaces associated to computable Kolmogorov spaces.
Corollary 30 (Computable prebases).
Let $X$ and $Y$ be computable Kolmogorov spaces and let $Z\subseteq X$. Then the following are computable prebases:
  1. (1)
    $B_{X\times Y}:{\mathcal{O}}(X)\times{\mathcal{O}}(Y)\to{\mathcal{O}}(X\times Y),(U,V)\mapsto U\times V$.
  2. (2)
    $B_{Y^{\mathbb{N}}}:{\mathcal{O}}(X)^{*}\to{\mathcal{O}}(Y^{\mathbb{N}}),(U_{1},...,U_{n})\mapsto U_{1}\times...\times U_{n}\times Y^{\mathbb{N}}$.
  3. (3)
    $B_{Z}:{\mathcal{O}}(X)\to{\mathcal{O}}(Z),U\mapsto U\cap Z$.
  4. (4)
    $B_{X\sqcup Y}:{\mathcal{O}}(X)\sqcup{\mathcal{O}}(Y)\to{\mathcal{O}}(X\sqcup Y),U\mapsto U$.
  5. (5)
    $B_{X\sqcap Y}:{\mathcal{O}}(X)\times{\mathcal{O}}(Y)\to{\mathcal{O}}(X\sqcap Y),(U,V)\mapsto U\cap V$.
$B_{X\sqcup Y}$ is even a computable base.
Using Theorem 23, Corollary 30 and Proposition 26 we can directly conclude that computable Kolmogorov spaces have very nice closure properties. We note that for this proof we only need the statements on computable prebases that were relatively easy to establish. By ${\mathcal{A}}_{-}(X)$ we denote the space of closed subsets of $X$ represented as complements of open sets in ${\mathcal{O}}(X)$. This space is hence computably isomorphic to ${\mathcal{O}}(X)$. We also consider the spaces ${\mathcal{A}}(X):={\mathcal{A}}_{+}(X)\sqcap{\mathcal{A}}_{-}(X)$ and ${\mathcal{K}}(X):={\mathcal{A}}_{+}(X)\sqcap{\mathcal{K}}_{-}(X)$ (we note that the latter contains the compact, closed and saturated sets).
Corollary 31 (Closure properties of computable Kolmogorov spaces).
If $X$ and $Y$ are computable Kolmogorov spaces, then so are
  1. (1)
    $X\times Y$, $X\sqcup Y$, $X\sqcap Y$, $Y^{\mathbb{N}}$, and every subspace of $X$.
Let $X$ be a represented space. If $Y$ is a computable Kolmogorov space, then so are
  1. (2)
    ${\mathcal{C}}(X,Y)$, ${\mathcal{O}}(X)$, ${\mathcal{A}}_{+}(X)$, ${\mathcal{A}}_{-}(X)$, ${\mathcal{A}}(X)$, ${\mathcal{K}}_{-}(X)$ and ${\mathcal{K}}(X)$.
Analogous statements hold for continuous Kolmogorov spaces.
In fact, some of these results can also be derived from each other in various ways. For instance, the statement on ${\mathcal{C}}(X,Y)$ applied to the special case of $X={\mathbb{N}}$ yields the statement for $Y^{\mathbb{N}}$ and the special case of $Y={\mathbb{S}}$ yields the statements for ${\mathcal{O}}(X)$ (and ${\mathcal{A}}_{-}(X)$). This implies the statement for ${\mathcal{O}}{\mathcal{O}}(X)$, subspaces of which are computably isomorphic to ${\mathcal{A}}_{+}(X)$ and ${\mathcal{K}}_{-}(X)$, respectively. We obtain the following conclusion (see [Sch02a, Sch02b] for the topological version) using the fact that our product and function space constructions satisfy evaluation and currying properties (see [Wei00, Wei87]).
Corollary 32 (Schröder 2002).
The category of computable Kolmogorov spaces is cartesian closed.
For Corollaries 31 and 32 it was sufficient to have computable presubbases for the respective spaces. The fact that we have even computable prebases for some of these spaces allows further conclusions regarding the underlying topologies. Using Corollaries 16 and 30 and Theorem 25 we obtain the following conclusions. By ${\mathcal{O}}(X)\otimes{\mathcal{O}}(Y)$ and $\bigotimes_{i\in{\mathbb{N}}}{\mathcal{O}}(X)$ we denote the respective product topologies, by ${\mathcal{O}}(X)|_{Z}$ the subspace topology of $Z\subseteq X$ and by ${\mathcal{O}}(X)\wedge{\mathcal{O}}(Y)$ the meet topology generated by the base $B_{X\sqcap Y}$.
Corollary 33 (Topologies and set-theoretic constructions).
Let $X$ and $Y$ be represented ${\mathsf{{T}}}_{0}$ spaces and let $Z\subseteq X$ be a subspace. Then:
  1. (1)
    ${\mathcal{O}}(X\times Y)={\mathrm{seq}}({\mathcal{O}}(X)\otimes{\mathcal{O}}(Y))$,
  2. (2)
    ${\mathcal{O}}(X^{\mathbb{N}})={\mathrm{seq}}(\bigotimes_{i\in{\mathbb{N}}}{\mathcal{O}}(X))$,
  3. (3)
    ${\mathcal{O}}(Z)={\mathrm{seq}}({\mathcal{O}}(X)|_{Z})$,
  4. (4)
    ${\mathcal{O}}(X\sqcap Y)={\mathrm{seq}}({\mathcal{O}}(X)\wedge{\mathcal{O}}(Y))$.
We note that for computable Kolmogorov spaces $X$ and $Y$ this result follows from Corollary 30 in combination with Corollary 16. With the help of Theorem 25 we can transfer this result then to arbitrary represented ${\mathsf{{T}}}_{0}$ spaces.
5. Hyperspace and function space topologies
In this section we continue to discuss what can be concluded from our results regarding hyperspace and function space topologies. Once again, these results are well-known by the work of Schröder [Sch02a] and the purpose here is to demonstrate how these results can be obtained with the help of the concept of a computable presubbase or prebase.
As a preparation of our results we need the following lemma, which shows that the subbases $\Diamond$ and $\triangleright$ satisfy the convergent intersection property. This allows us to apply Corollary 12.
Lemma 34 (Countable intersections).
Let $X$ and $Y$ be represented space. Then the presubbases
  1. (1)
    $\Diamond:{\mathcal{O}}(Y)\to{\mathcal{O}}{\mathcal{A}}_{+}(Y)$ and
  2. (2)
    $\triangleright:{\mathcal{K}}_{-}(X)\times{\mathcal{O}}(Y)\to{\mathcal{O}}{\mathcal{C}}(X,Y)$
satisfy the convergent intersection property.
Proof.
Let $(U_{n})_{n\in{\mathbb{N}}_{\infty}}$ and $(K_{n})_{n\in{\mathbb{N}}_{\infty}}$ be converging sequences in ${\mathcal{O}}(Y)$ and ${\mathcal{K}}_{-}(X)$, respectively. We prove that there exists some $k\in{\mathbb{N}}$ such that
  1. (1)
    $\bigcap_{n\in{\mathbb{N}}_{\infty}}\Diamond U_{n}=\Diamond U_{\infty}\cap\bigcap_{n=0}^{k}\Diamond U_{n}$.
  2. (2)
    $\bigcap_{n\in{\mathbb{N}}_{\infty}}(K_{n}\triangleright U_{n})=(K_{\infty}\triangleright U_{\infty})\cap\bigcap_{n=0}^{k}(K_{n}\triangleright U_{n})$.
By Proposition 8 convergence of $(U_{n})_{n\in{\mathbb{N}}_{\infty}}$ implies that there is a $k\in{\mathbb{N}}$ with $U_{\infty}\subseteq\left(\bigcap_{n\geq k}U_{n}\right)^{\circ}$. In particular, $U_{\infty}\subseteq U_{n}$ for all $n\geq k$. The inclusion $\subseteq$” in (1) and (2) is clear in both cases and we need to show why the inverse inclusions hold.
(1) Let $A\in\Diamond U_{\infty}\cap\bigcap_{n=0}^{k}\Diamond U_{n}$. Then, in particular, $A\cap U_{\infty}\not$$\emptyset$ and hence $A\cap U_{n}\not$$\emptyset$ for all $n\geq k$. This proves the claim.
(2) Let $f\in(K_{\infty}\triangleright U_{\infty})\cap\bigcap_{n=0}^{k}(K_{n}\triangleright U_{n})$ be a continuous function $f:X\to Y$. Then, in particular, $f(K_{\infty})\subseteq U_{\infty}$. As $(f(K_{n}))_{n\in{\mathbb{N}}}$ converges in ${\mathcal{K}}_{-}(Y)$, we obtain (see the independently proved statement in Theorem 35 (2)) that there exists a $k^{\prime}\in{\mathbb{N}}$ with $f(K_{n})\subseteq U_{\infty}$ for $n\geq k^{\prime}$ and hence $f(K_{n})\subseteq U_{n}$ for all $n\geq k^{\prime\prime}:=\max(k,k^{\prime})$. This proves the claim. ∎
We use the terminology for topologies introduced in the table in Figure 2 and we add some terminology:
  1. (1)
    The upper Fell topology on the hyperspace ${\mathcal{A}}_{-}(X)$ is the topology inherited (by complementation) from the compact-open topology on ${\mathcal{O}}(X)$ and the Fell topology on ${\mathcal{A}}(X)$ is the meet topology of the lower and upper Fell topologies.
  2. (2)
    The Vietoris topology on ${\mathcal{K}}(X)$ is the meet topology of the lower Fell topology and the upper Vietoris topology.
From Proposition 26 we can obtain information on the topologies of the respective hyperspaces, a fact which we already indicated in Example 27.
Theorem 35 (Hyperspace and function space topologies).
Let $X$ be a represented space and let $Y$ be an admissibly represented ${\mathsf{{T}}}_{0}$ space.
  1. (1)
    ${\mathcal{O}}(X)$ is endowed with the Scott topology, which is the sequentialization of the compact-open topology.
  2. (2)
    ${\mathcal{K}}_{-}(X)$ and ${\mathcal{K}}(X)$ are endowed with topologies that are the sequentializations of the upper Vietoris and the Vietoris topology, respectively.
  3. (3)
    ${\mathcal{A}}_{+}(X)$ and ${\mathcal{A}}(X)$ are endowed with topologies that are the sequentializations of the lower Fell and the Fell topology, respectively.
  4. (4)
    ${\mathcal{C}}(X,Y)$ is endowed with the sequentialization of the compact-open topology.
In particular, all the mentioned spaces are admissibly represented with respect to the given topologies.
Proof.
(1) This is the statement of Corollary 9. That ${\mathcal{O}}{\mathcal{O}}(X)$ is the sequentialization of the topology generated by the sets ${\mathcal{F}}_{K}$ for (saturated) compact $K\subseteq X$ follows also from Proposition 26 in combination with Corollary 16.
(2) That ${\mathcal{O}}{\mathcal{K}}_{-}(X)$ is the sequentialization of the topology generated by the sets $\Box U$ for open $U\subseteq X$ follows from Proposition 26 in combination with Corollary 16.
(3) That ${\mathcal{O}}{\mathcal{A}}_{+}(X)$ is the sequentialization of the lower Fell topology generated by the sets $\Diamond U$ for open $U\subseteq X$ follows from Theorem 2, Corollary 12 and Lemma 34. With the help of Corollary 33 this implies that ${\mathcal{K}}(X)={\mathcal{A}}_{+}(X)\sqcap{\mathcal{K}}_{-}(X)$ and ${\mathcal{A}}(X)={\mathcal{A}}_{+}(X)\sqcap{\mathcal{A}}_{-}(X)$ are endowed with the sequentializations of the Vietoris and Fell topologies, respectively.
(4) If $Y$ is an admissibly represented ${\mathsf{{T}}}_{0}$ space then Theorem 25 allows us to replace it by a computable Kolmogorov space without affecting the topological structure. In fact, the resulting representation is in the same continuous equivalence class as the original and hence so are the corresponding representations of ${\mathcal{C}}(X,Y)$. By Theorem 2, Corollary 12 and Lemma 34 we know that ${\mathcal{O}}{\mathcal{C}}(X,Y)$ is the sequentialization of the compact-open topology on ${\mathcal{C}}(X,Y)$. ∎
The examples of the spaces ${\mathcal{A}}_{+}(X)$ and ${\mathcal{C}}(X,Y)$ show that even in cases where we have only a computable presubbase (and not a computable prebase) it might happen that we obtain as a topology on our space the sequentialization of the topology generated by the presubbase.
6. Epilogue
We close this section with a brief recap of our results from a somewhat different perspective. The results in this article can be viewed as exploiting a Galois connection between maps
[…]
which was already introduced and discussed in [BR25]. Let us fix a set $X$ and let us denote by $\mathsf{REP}_{0}$ the set of representations $\delta:\subseteq{\mathbb{N}}^{\mathbb{N}}\to X$ with a ${\mathsf{{T}}}_{0}$ final topology and let us denote by $\mathsf{PRE}_{0}$ the set of presubbases $B:\subseteq{\mathbb{N}}^{\mathbb{N}}\to 2^{X}$ that induce a ${\mathsf{{T}}}_{0}$ topology on XX. [^3] We call $B:\subseteq{\mathbb{N}}^{\mathbb{N}}\to 2^{X}$ a presubbase if $(B_{y})_{y\in D}$ with $D:={\mathrm{dom}}(B)$ is a presubbase in the sense of our definition.
  1. (1)
    The map $\Delta:\mathsf{PRE}_{0}\to\mathsf{REP}_{0},B\mapsto\delta^{B}$ assigns to any presubbase $B$ the presubbase representation $\delta^{B}:\subseteq{\mathbb{N}}^{\mathbb{N}}\to X$.
  2. (2)
    The map $\nabla:\mathsf{REP}_{0}\to\mathsf{PRE}_{0},\delta\mapsto B_{\delta}$ assigns to every representation $\delta$ of $X$ the induced representation $\delta_{{\mathcal{O}}(X)}$ as presubbase $B_{\delta}:\subseteq{\mathbb{N}}^{\mathbb{N}}\to 2^{X}$.
The presubase $B_{\delta}$ corresponds to the base ${\mathrm{id}}:{\mathcal{O}}(X)\to{\mathcal{O}}(X)$. The maps $\Delta$ and $\nabla$ induce an antitone Galois connection, if we define a computable reducibility $\leq$ for $\mathsf{PRE}_{0}$ as for $\mathsf{REP}_{0}$ (for $\mathsf{REP}_{0}$ we use the usual computable reducibility, as introduced before Theorem 25). Namely for $B_{1},B_{2}\in\mathsf{PRE}_{0}$ we define analogously $B_{1}\leq B_{2}$ if there is some computable $F:\subseteq{\mathbb{N}}^{\mathbb{N}}\to{\mathbb{N}}^{\mathbb{N}}$ with $B_{1}(p)=B_{2}F(p)$ for all $p\in{\mathrm{dom}}(B_{1})$. The proof of the following result is then straightforward and a simple consequence of properties of the transposition operation.
Theorem 36 (Galois connection for representations and presubbases).
Fix a representable space $X$. Then for every $\delta\in\mathsf{REP}_{0}$ and $B\in\mathsf{PRE}_{0}$ we have
$\delta\leq\delta^{B}\iff B\leq B_{\delta}.$
Proof.
Let $\delta:\subseteq{\mathbb{N}}^{\mathbb{N}}\to X$ be a representation with a ${\mathsf{{T}}}_{0}$ topology ${\mathcal{O}}(X)$ and induced representation $\delta_{{\mathcal{O}}(X)}$ of ${\mathcal{O}}(X)$ and let $B:\subseteq{\mathbb{N}}^{\mathbb{N}}\to 2^{X}$ be a presubbase that induces a ${\mathsf{{T}}}_{0}$ topology on $X$. Then we obtain:
$\delta\leq\delta^{B}$ $\iff$ $B^{\mathsf{{T}}}:\subseteq X\to{\mathcal{O}}({\mathbb{N}}^{\mathbb{N}})\mbox{ is computable with respect to $\delta$}$
$\iff$ $B:\subseteq{\mathbb{N}}^{\mathbb{N}}\to{\mathcal{O}}(X)\mbox{ is computable with respect to $\delta_{{\mathcal{O}}(X)}$}$
$\iff$ $B\leq B_{\delta},$
which proves the claim. ∎
We also obtain that ${\mathrm{range}}(\Delta)$ is exactly the set of representations of $X$ that turn $X$ into a computable Kolmogorov space (up to computable equivalence) and ${\mathrm{range}}(\nabla)$ is exactly the set of computable bases of $\mathsf{qcb}_{0}$ topologies for $X$ (up to computable equivalence).
Any antitone Galois connection yields corresponding closure operators, sometimes called monad and comonad. In the case of our Galois connection we obtain:
  1. (1)
    $\Delta\circ\nabla:\mathsf{REP}_{0}\to\mathsf{REP}_{0},\delta\mapsto\delta^{\bullet}:=\delta^{B_{\delta}}$ and
  2. (2)
    $\nabla\circ\Delta:\mathsf{PRE}_{0}\to\mathsf{PRE}_{0},B\mapsto B_{\bullet}:=B_{\delta^{B}}$,
We summarize some properties of these closure operators (see Theorem 25, Corollary 14 and Proposition 22):
  1. (1)
    $\delta\mapsto\delta^{\bullet}$ maps every represented ${\mathsf{{T}}}_{0}$ space $(X,\delta)$ to a computable Kolmogorov space $X^{\bullet}:=(X,\delta^{\bullet})$ with $\delta\leq\delta^{\bullet}$. The induced underlying final topologies and the representations thereof are preserved by this operation (up to computable equivalence).
  2. (2)
    $B\mapsto B_{\bullet}$ maps any presubbase $B$ of a ${\mathsf{{T}}}_{0}$ topology on $X$ to a computable base $B_{\bullet}$ of the sequentialization ${\mathrm{seq}}(\tau)$ of the topology $\tau$ that is generated by the compact intersections $\bigcap_{\mathcal{K}}B$. In particular, ${\mathrm{range}}(B)\subseteq{\mathrm{range}}(B_{\bullet})$. The operation preserves the induced presubbase representations of $X$ (up to computable equivalence).
Since both application are closure operators, a double application does not yield anything new, i.e., $\delta^{\bullet\bullet}\equiv\delta^{\bullet}$ and ${\mathrm{range}}(B_{\bullet\bullet})={\mathrm{range}}(B_{\bullet})$. We give two examples of the action of these closure operator that indicates that these are natural operations. The first example is taken from Example 27, the second one is due to Schröder (personal communication).
Example 37 (Action of closure operators).
  1. (1)
    If we start with the presubbase $B$ for the point-open topology on ${\mathcal{O}}(X)$ then we obtain a base $B_{\bullet}$ of the Scott topology. Both bases induce computably equivalent presubbase representations of ${\mathcal{O}}(X)$.
  2. (2)
    If we start with a decimal representation $\rho$ of the real numbers $\mathbb{R}$}, then we obtain a representation $\rho^{\bullet}$ that is computably equivalent to the Cauchy representation of the reals. Both representations have the same final topology (the Euclidean topology).
In light of the Galois connection between presubbases and representations we can interpret our results as follows. A lot of results that were originally proved with the help of representations and, for instance, the closure operator $\delta\mapsto\delta^{\bullet}$ by Schröder (for example the topological version of Corollary 32 in [Sch02a, Section 4.3]) can equally well be derived with the help of presubbases. Both perspectives correspond to different sides of the same medal and they highlight different aspects of the same mathematical structure.
The purpose of this article was to demonstrate the usefulness of considering presubbases. The benefit of considering presubbases is that they allow us to talk about topologies using their usual subbases (see, e.g., Proposition 26) and the usual constructions of new subbases from given ones (see, e.g., Proposition 29). This enriches the mathematical vocabulary that we have to discuss computable topology.
Acknowledgments
This work was supported by the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) – project number 554999067, by the National Research Foundation of South Africa (NRF) – grant number 151597, and by the Alexander-von-Humboldt Foundation. We would like to thank Matthias Schröder for discussions on the content of this article that have helped to streamline the proofs and their presentations.