### v2

# Computable Bases

Abstract.
In computable analysis typically topological spaces with countable bases are considered. The Theorem of Kreitz-Weihrauch implies that the subbase representation of a second-countable ${\mathsf{{T}}}_{0}$ space is admissible with respect to the topology that the subbase generates. We consider generalizations of this setting to bases that are representable, but not necessarily countable. We introduce the notions of a computable presubbase and a computable prebase. We prove a generalization of the Theorem of Kreitz-Weihrauch for the presubbase representation that shows that any such representation is admissible with respect to the topology generated by compact intersections of the presubbase elements. For computable prebases we obtain representations that are admissible with respect to the topology that they generate. These concepts provide a natural way to investigate many topological spaces that have been studied in computable analysis. The benefit of this approach is that topologies can be described by their usual subbases and standard constructions for such subbases can be applied. Finally we discuss a Galois connection between presubbases and representations of ${\mathsf{{T}}}_{0}$ spaces that indicates that presubbases and representations offer particular views on the same mathematical structure from different perspectives.
1991 Mathematics Subject Classification:
[Theory of computation]: Logic; [Mathematics of computing]: Continuous mathematics.
Version of .
1. Introduction
In the representation based approach to computable analysis, as initiated by Kreitz and Weihrauch [KW85, Wei87, Wei00], we work with representations $\delta:\subseteq{\mathbb{N}}^{\mathbb{N}}\to X$, which are surjective partial maps. In this situation $(X,\delta)$, or briefly $X$, is called a represented space. We say that a function $F:\subseteq{\mathbb{N}}^{\mathbb{N}}\to{\mathbb{N}}^{\mathbb{N}}$ realizes a multivalued map $f:\subseteq X\rightrightarrows Y$ on represented spaces $(X,\delta_{X})$ and $(Y,\delta_{Y})$, in symbols, $F\vdash f$, if $\delta_{Y}F(p)\in f\delta_{X}(p)$ for all $p\in{\mathrm{dom}}(f\delta_{X})$. Then $f$ is called computable or continuous, if it has a realizer of the corresponding type.
On the other hand, every represented space $(X,\delta_{X})$ comes naturally equipped with a topology ${\mathcal{O}}(X)$, namely the final topology
${\mathcal{O}}(X):=\{U\subseteq X:\delta_{X}^{-1}(U)\mbox{ open in }{\mathrm{dom}}(\delta_{X})\}$
induced by the Baire space topology on $\mathbb{N}$$^{\mathbb{N}}$ via the map $\delta_{X}$ on $X$. If not mentioned otherwise, we will assume that ${\mathcal{O}}(X)$ is the final topology of the represented space $X$. The topological spaces whose topology occurs in this way have a special name, they are called $\mathsf{qcb}$ spaces (quotients of countably based spaces). These spaces share some special properties [Sch02b], for instance, they are always sequential, i.e., a subset is closed if and only if it is sequentially closed, which means that it is closed under limits of converging sequences. We also automatically have a representation $\delta_{{\mathcal{O}}(X)}$ of this topology ${\mathcal{O}}(X)$ that is induced by a function space representation of the space ${\mathcal{C}}(X,{\mathbb{S}})$ of the continuous functions $f:X\to{\mathbb{S}}$ with the Sierpiński space ${\mathbb{S}}=\{0,1\}$. This is because $U\subseteq X$ is open if and only if its characteristic function $\chi_{U}:X\to{\mathbb{S}}$ is continuous. Matthias Schröder has developed a theory of computable topology based on these concepts [Sch02b, Sch02a, Sch21] (see also Pauly [Pau16] for a concise presentation of some aspects of computable topology). The open sets $U\subseteq X$ that are computable in ${\mathcal{O}}(X)$ are called c.e. open and these are exactly the sets which computable characteristic function $\chi_{U}:X\to{\mathbb{S}}$.
An essential question is when continuity in terms of the representations $\delta_{X}$ and $\delta_{Y}$ (as defined above) coincides for a singlevalued function $f:X\to Y$ with the ordinary notion of continuity with respect to the topologies ${\mathcal{O}}(X)$ and ${\mathcal{O}}(Y)$. Every function $f$ with a continuous realizer is continuous in the ordinary topological sense, but not necessarily the other way around [Sch02b]. The two concepts of continuity are equivalent if $\delta_{X}$ and $\delta_{Y}$ are so-called admissible representations with respect to these topologies. The notion of admissibility was originally defined by Kreitz and Weihrauch just for second-countable ${\mathsf{{T}}}_{0}$ spaces [KW85] and then later extended to arbitrary topological spaces by Schröder [Sch02b]. We will adopt the latter definition.
Definition 1 (Admissibility).
A representation $\delta_{X}:\subseteq{\mathbb{N}}^{\mathbb{N}}\to X$ is called admissible with respect to a topology $\tau$ on $X$ if the following hold:
  1. (1)
    $\delta_{X}$ is continuous with respect to the topology $\tau$ and
  2. (2)
    any other representation $\delta:\subseteq{\mathbb{N}}^{\mathbb{N}}\to X$ that is continuous with respect to $\tau$ is also continuous with respect to the representation $\delta_{X}$.
A represented space is called admissible if the representation of the space is admissible with respect to the final topology ${\mathcal{O}}(X)$ induced by it.
Admissibility of $\delta_{X}$ means that any continuous representation $\delta$ can be continuously translated into $\delta_{X}$. Hence the names $p\in{\mathbb{N}}^{\mathbb{N}}$ of $\delta_{X}$ provide the minimal information on the points $x\in X$ that is required for the representation to be continuous in the topological sense. Schröder proved [Sch02b, Theorem 7, Lemma 8] that whether a representation $\delta_{X}$ is admissible with respect to some topology $\tau$ only depends on the sequentialization of that topology [^1] The sequentialization ${\mathrm{seq}}(\tau)$ of a topology $\tau$ is the smallest topology that contains $\tau$ and is sequential..
Theorem 2 (Schröder 2002).
Let $(X,\delta_{X})$ be a represented space with final topology ${\mathcal{O}}(X)$ and let $\tau$ be some topology on $X$. Then the following are equivalent:
  1. (1)
    $\delta_{X}$ is admissible with respect to $\tau$.
  2. (2)
    $\delta_{X}$ is admissible with respect to ${\mathrm{seq}}(\tau)$.
If one these conditions holds, then ${\mathrm{seq}}(\tau)={\mathcal{O}}(X)$.
In particular, this means that if $\delta$ is admissible with respect to some topology, then also with respect to ${\mathcal{O}}(X)$. Schröder also showed [Sch02b, Theorem 13] that a topology for which an admissible representation exists is necessarily ${\mathsf{{T}}}_{0}$. The topological spaces that admit an admissible representation are exactly those whose sequentializations are so-called $\mathsf{qcb}_{0}$ spaces (the ${\mathsf{{T}}}_{0}$ spaces among the $\mathsf{qcb}$ spaces). Finally, Schröder also proved that the representation $\delta_{{\mathcal{O}}(X)}$ is always admissible with respect to the Scott topology on ${\mathcal{O}}(X)$ [Sch02a] and that admissibility of a represented space can be characterized with the help of the neighborhood map ${\mathcal{U}}:X\to{\mathcal{O}}{\mathcal{O}}(X),x\mapsto\{U\in{\mathcal{O}}(X):x\in U\}$ of the space as follows.
Proposition 3 (Admissibility).
A represented space $X$ is admissible if and only if its neighborhood map ${\mathcal{U}}:X\to{\mathcal{O}}{\mathcal{O}}(X)$ is a continuous embedding with respect to the involved representations.
Being an embedding with respect to the representations means that $\mathcal{U}$} is injective and $\mathcal{U}$} as well as its partial inverse have continuous realizers. Likewise, we define computable embeddings with computable realizers. The map $\mathcal{U}$} is always computable, hence the essential condition is the one that concerns its inverse. In light of this result one can consider admissibility also as an effectivization of the ${\mathsf{{T}}}_{0}$ property as the map $\mathcal{U}$} is injective if and only if the topology ${\mathcal{O}}(X)$ is ${\mathsf{{T}}}_{0}$ (${\mathsf{{T}}}_{0}$ spaces are also known as a Kolmogorov spaces). Consequently, we call $X$ a computable Kolmogorov space if $\mathcal{U}$} is a computable embedding. [^2] We prefer this terminology over the notion of computable admissibility that is also used for this concept, as one and the same topological space can be a computable Kolmogorov space in different inequivalent ways, whereas topological admissibility characterizes an equivalence class. The notion of admissibility is often linked to this uniqueness property.
Kreitz and Weihrauch [KW85] have shown that every second-countable ${\mathsf{{T}}}_{0}$ space has an admissible representation. Namely, if $B=(B_{n})_{n\in{\mathbb{N}}}$ is a subbase of some topology $\tau$ on $X$, then the subbase representation $\delta^{B}:\subseteq{\mathbb{N}}^{\mathbb{N}}\to X$ with
$\delta^{B}(p)=x:\iff{\mathrm{range}}(p)-1=\{n\in{\mathbb{N}}:x\in B_{n}\}$
for all $p\in{\mathbb{N}}^{\mathbb{N}}$ and $x\in X$ is admissible with respect to $\tau$. Here we do not use ${\mathrm{range}}(p)$ but ${\mathrm{range}}(p)-1:=\{n\in{\mathbb{N}}:n+1\in{\mathrm{range}}(p)\}$ in order to allow for the empty set to be enumerated. We formulate the aforementioned result as a theorem.
Theorem 4 (Kreitz-Weihrauch 1985).
If $X$ is a ${\mathsf{{T}}}_{0}$ space with a countable subbase $B$, then the subbase representation $\delta^{B}$ is admissible with respect to the topology of $X$.
This theorem does not only guarantee that there are sufficiently many nice representations, but it also provides a tool that can be used to show that a concrete given representation $\delta_{X}$ is admissible. One just needs to prove that the representation is topologically equivalent to some subbase representation $\delta^{B}$ with respect to a known subbase $B$ of the topology of interest, i.e., that ${\mathrm{id}}:X\to X$ is continuous in both directions with respect to $\delta_{X}$ and $\delta^{B}$, respectively.
One purpose of this article is to provide a similar tool for general (not necessarily second-countable) ${\mathsf{{T}}}_{0}$ spaces. A second goal is to develop a theory of computable bases for represented spaces very much along the lines of the second-countable case [BR26]. Finally, we also demonstrate that these concepts have interesting applications.
2. The Presubbase Theorem
We start with the following definition that extends the concept of a subbase representation beyond the countable case.
Definition 5 (Presubbase).
Let $X$ be a set. We call a family $(B_{y})_{y\in Y}$ of subsets of $X$ a presubbase for $X$, if $Y$ is a represented space and its transpose
$B^{\mathsf{{T}}}:X\to{\mathcal{O}}(Y),x\mapsto\{y\in Y:x\in B_{y}\}$
is well-defined and injective.
Injectivity of $B^{\mathsf{{T}}}$ implies that $(B_{y})_{y\in Y}$ is a subbase of some ${\mathsf{{T}}}_{0}$ topology on $X$. We note that every countable subbase $B:{\mathbb{N}}\to{\mathcal{O}}(X)$ of a ${\mathsf{{T}}}_{0}$ topology is a particular instance of a presubbase, as $B^{\mathsf{{T}}}:X\to{\mathcal{O}}({\mathbb{N}})$ is always well-defined. Hence, the following definition generalizes the concept of a subbase representation by Kreitz and Weihrauch as it is specified above.
Definition 6 (Presubbase representation).
Let $(B_{y})_{y\in Y}$ be a presubbase of a set $X$. We define the presubbase representation $\delta^{B}:\subseteq{\mathbb{N}}^{\mathbb{N}}\to X$ by
$\delta^{B}(p)=x:\iff\delta_{{\mathcal{O}}(Y)}(p)=\{y\in Y:x\in B_{y}\}$
for all $p\in{\mathbb{N}}^{\mathbb{N}}$ and $x\in X$.
We note that the map $\delta^{B}$ is well-defined as $B^{\mathsf{{T}}}$ is well-defined and injective. For short, we can also write $\delta^{B}=(B^{\mathsf{{T}}})^{-1}\circ\delta_{{\mathcal{O}}(Y)}$. The reason that we speak about a presubbase representation in this general situation and not about a subbase representation is that $\delta^{B}$ is not necessarily admissible with respect to the topology generated by $(B_{y})_{y\in Y}$. However, it is admissible with respect to a closely related topology generated by compact intersections of the sets $B_{y}$, which is our first main result. By a compact set $K\subseteq X$ we mean a set with the property that every open cover has a finite subcover (no Hausdorff condition involved).
Theorem 7 (Presubbase theorem).
Let $(B_{y})_{y\in Y}$ be a presubbase of a set $X$. Then $(X,\delta^{B})$ is a computable Kolmogorov space and $\delta^{B}$ is admissible with respect to the topology $\tau$ on $X$ that is generated by the base sets $X$ and $\bigcap_{y\in K}B_{y}$ for every compact $K\subseteq Y$.
We note that this result generalizes Theorem 4 by Kreitz-Weihrauch as for countable subbases $B:{\mathbb{N}}\to{\mathcal{O}}(X)$ the compact subsets $K\subseteq{\mathbb{N}}$ are exactly the finite subsets and hence the topology generated by $X$ and $\bigcap_{n\in K}B_{n}$ for compact $K\subseteq{\mathbb{N}}$ is exactly the same topology as the topology generated by the subbase $B$ itself.
[…]
Proposition 8 (Scott convergence).
Let $X$ be a sequential topological space. Let $(U_{n})$ be a sequence in ${\mathcal{O}}(X)$ and $U\in{\mathcal{O}}(X)$. Then the following are equivalent:
  1. (1)
    $U_{n}\to U$ with respect to the Scott topology.
  2. (2)
    $U_{n}\to U$ with respect to the topology, which is generated by the base of sets ${\mathcal{F}}_{K}$ over all compact $K\subseteq X$.
  3. (3)
    $U_{n}\to U$ with respect to the topology, which is generated by the subbase of sets ${\mathcal{F}}_{\{x,x_{n}:n\in{\mathbb{N}}\}}$ over all sequences $(x_{n})$ and $x$ in $X$ such that $x_{n}\to x$.
  4. (4)
    $U\subseteq\bigcup_{k\in{\mathbb{N}}}\left(\bigcap_{n\geq k}U_{n}\right)^{\circ}$.
Here $A^{\circ}$ denotes the interior of the set $A$. As a consequence of this result, ${\mathcal{O}}(X)$ is not just admissibly represented with respect to the Scott topology, but also with respect to the compact-open topology. By Theorem 2 we obtain the following characterization of the topology of the space ${\mathcal{O}}(X)$.
Corollary 9 (Schröder 2002).
Let $X$ be a represented space. Then ${\mathcal{O}}(X)$ is endowed with the Scott topology that is the sequentialization of the compact-open topology and the standard representation of ${\mathcal{O}}(X)$ is admissible with respect to both topologies.
Now we are actually prepared to prove Theorem 7.
Proof of Theorem 7.
Since $(B_{y})_{y\in Y}$ is a presubbase of $X$, the map $B^{\mathsf{{T}}}:X\to{\mathcal{O}}(Y)$ is well-defined and injective. We consider the represented space $(X,\delta^{B})$ and we prove that it is a computable Kolmogorov space. We represent ${\mathcal{O}}(X)$ and ${\mathcal{O}}(Y)$ in the usual way, which means that they are endowed with the respective Scott topologies by Corollary 9. The definition $\delta^{B}=(B^{\mathsf{{T}}})^{-1}\circ\delta_{{\mathcal{O}}(Y)}$ ensures that $B^{\mathsf{{T}}}:X\to{\mathcal{O}}(Y)$ becomes a computable embedding with respect to $\delta^{B}$ as a representation of $X$. This implies that $B:Y\to{\mathcal{O}}(X)$ is computable too. We need to prove that also ${\mathcal{U}}:X\to{\mathcal{O}}{\mathcal{O}}(X)$ is a computable embedding. Given ${\mathcal{U}}_{x}=\{U\in{\mathcal{O}}(X):x\in U\}\in{\mathcal{O}}{\mathcal{O}}(X)$ for some $x\in X$, we can compute
$B^{\mathsf{{T}}}_{x}=\{y\in Y:x\in B_{y}\}=B^{-1}({\mathcal{U}}_{x})\in{\mathcal{O}}(Y)$
since $B$ is computable. And since $B^{\mathsf{{T}}}$ is a computable embedding, we can compute $x\in X$ from this set. Hence, also ${\mathcal{U}}:X\to{\mathcal{O}}{\mathcal{O}}(X)$ is a computable embedding and $(X,\delta^{B})$ is a computable Kolmogorov space.
In particular, $\delta^{B}$ is admissible with respect to its final topology ${\mathcal{O}}(X)$ by Proposition 3. We need to prove that it is also admissible with respect to the topology $\tau$. We already know by Corollary 9 that ${\mathcal{O}}(Y)$ is admissibly represented with respect to the compact-open topology generated by the sets ${\mathcal{F}}_{K}=\{U\in{\mathcal{O}}(Y):K\subseteq U\}$ over all compact $K\subseteq Y$. Now we note that
$\bigcap_{y\in K}B_{y}=\{x\in X:K\subseteq B_{x}^{\mathsf{{T}}}\}=(B^{\mathsf{{T}}})^{-1}({\mathcal{F}}_{K})$
and $(B^{\mathsf{{T}}})^{-1}({\mathcal{O}}(Y))=X$. Hence, $\tau$ is the initial topology of $B^{\mathsf{{T}}}$ with respect to the compact-open topology on ${\mathcal{O}}(Y)$, and hence $\delta^{B}=(B^{\mathsf{{T}}})^{-1}\circ\delta_{{\mathcal{O}}(Y)}$ is admissible with respect to $\tau$ by [Sch02b, Section 4.2] (see also [Sch02a, Proposition 4.1.4]). ∎
For simplicity we denote convergent sequences $(x_{n})_{n\in{\mathbb{N}}}$ that converge to $x_{\infty}$ as $(x_{n})_{n\in{\mathbb{N}}_{\infty}}$ with $\mathbb{N}$$_{\infty}$$\mathbb{N}$$\cup\{\infty\}$. By Proposition 8 we could alternatively also use the subbase sets ${\mathcal{F}}_{K}$ with $K=\{y_{n}:n\in{\mathbb{N}}_{\infty}\}$ over all convergent sequences $(y_{n})_{n\in{\mathbb{N}}_{\infty}}$ in $Y$ in the previous proof. Hence, we obtain the following.
Remark 10.
The statement of Theorem 7 remains true if we replace the intersections $\bigcap_{y\in K}B_{y}$ by intersections $\bigcap_{n\in{\mathbb{N}}_{\infty}}B_{y_{n}}$ over all convergent sequences $(y_{n})_{n\in{\mathbb{N}}_{\infty}}$ in $Y$.
General so called $Y$-indexed bases of type $B:Y\to{\mathcal{O}}(X)$ have also been independently considered in [dBSS16, Definition 6.1]. The remark above relates our results regarding presubbases to the so-called sequentially YY–indexed generating systems considered in [dBSS16, Definition 7.1].
When we have a presubbase $B$ of a represented space $X$, then we have to deal in general with at least four different topologies, all generated by $B$ in different ways. We illustrate these in Figure 1. By $\bigcap_{\mathcal{K}}B$ we denote the subbase described in Theorem 7 and by $\bigcap_{\infty}B$ we denote the subbase generated over all convergent sequences, as described in Remark 10.
topology generated
by $B$ as a subbase
$\subseteq\,\,$ topology generated by $\bigcap_{\infty}B$ as a subbase $\subseteq\,\,$ topology generated by $\bigcap_{\mathcal{K}}B$ as a subbase $\subseteq\,\,$ final topology of the representation $\delta^{B}$
Figure 1. Topologies associated to a presubbase $B$ of a represented space.
The fourth (right-hand side) topology in Figure 1 is always the sequentialization of the second and of the third one. These are consequences of Theorem 2, combined with Remark 10 and Theorem 7, respectively. In the case of countable presubbases $B:{\mathbb{N}}\to{\mathcal{O}}(X)$ all four topologies coincide. We introduce a name for presubbases for which the first and the second topologies coincide.
Definition 11 (Convergent intersection property).
We say that a presubbase $B:Y\to{\mathcal{O}}(X)$ of a represented space $X$ satisfies the convergent intersection property if for every convergent sequence $(y_{n})_{n\in{\mathbb{N}}_{\infty}}$ in $Y$ there exists a finite set $F\subseteq Y$ with $\bigcap_{n\in{\mathbb{N}}_{\infty}}B_{y_{n}}=\bigcap_{y\in F}B_{y}$.
For presubbases $B$ with the convergent intersection property all four topologies in Figure 1 have the fourth one as sequentialization.
Corollary 12 (Convergent intersection property).
Let $(B_{y})_{y\in Y}$ be a presubbase of a set $X$ that satisfies the convergent intersection property. Then $\delta^{B}$ is admissible with respect to the topology $\tau$ on $X$ that is generated by $B$.
3. Computable Presubbases, Prebases and Bases
In this section we want to develop a theory of computable bases $B:Y\to{\mathcal{O}}(X)$ along the lines of the theory of computable bases for second-countable spaces [BR26]. Bases with more complicated index sets have already been studied in other contexts, see, e.g., [dBSS16]. Firstly, the proof of Theorem 7 already indicates how we can define computable presubbases.
Definition 13 (Computable presubbase).
Let $X$ and $Y$ be represented spaces. Then $B:Y\to{\mathcal{O}}(X)$ is called a computable presubbase of $X$ if the transpose
$B^{\mathsf{{T}}}:X\to{\mathcal{O}}(Y),x\mapsto\{y\in Y:x\in B_{y}\}$
is a well-defined computable embedding.
We note that $B^{\mathsf{{T}}}$ is well-defined and computable if and only if $B$ is so. Obviously, a presubbase $B$ of $X$ is a computable presubbase of $X$ if and only if the representation of $X$ is computably equivalent to $\delta^{B}$. This observation together with Theorems 7 and 2 yield the following corollary.
Corollary 14 (Presubbases).
If $X$ is a represented space with a computable presubbase $B:Y\to{\mathcal{O}}(X)$, then $X$ is a computable Kolmogorov space and ${\mathcal{O}}(X)={\mathrm{seq}}(\tau)$ for the topology $\tau$ that is generated by the base sets $X$ and $\bigcap_{y\in K}B_{y}$ for every compact $K\subseteq Y$.
In analogy to the countable case we can now define the concept of a computable base. In the second-countable case we demand that every finite intersection of subbase sets should be computably representable as a countable union of subbase sets. If we replace “finite” by “compact” and “countable” by “overt”, then we obtain the general concept of a computable base. By ${\mathcal{K}}_{-}(Y)$ we denote the set of saturated compact subsets $K\subseteq Y$ represented via continuous maps $\forall_{K}:{\mathcal{O}}(X)\to{\mathbb{S}}$ and by ${\mathcal{A}}_{+}(Y)$ we denote the space of closed subsets $A\subseteq Y$ represented via continuous maps $\exists_{A}:{\mathcal{O}}(X)\to{\mathbb{S}}$. Here $\forall_{K}$ is just the characteristic function of ${\mathcal{F}}_{K}$ and $\exists_{A}$ is the characteristic function of ${\mathcal{T}}_{A}:=\{U\in{\mathcal{O}}(X):A\cap U\not$$\varnothing\}$ (see [Pau16, Sch21]). We recall that a set $A\subseteq X$ is called overt, if $\exists_{A}$ is computable. Using the above terminology we obtain the following straightforward definition.
Definition 15 (Computable prebase).
Let $X$ be a represented space with a computable presubbase $B:Y\to{\mathcal{O}}(X)$. Then $B$ is called a computable prebase of $X$, if there is a computable $R:{\mathcal{K}}_{-}(Y)\rightrightarrows{\mathcal{A}}_{+}(Y)$ such that
$\bigcap_{y\in K}B_{y}=\bigcup_{y\in A}B_{y}$
for every $K\in{\mathcal{K}}_{-}(Y)$ and $A\in R(K)$. We work with the assumption that $\bigcap_{y\in\varnothing}B_{y}=X$. We call $B$ a computable base of $X$ if $B$ is actually a base of $X$.
Our definition of a computable base is a computable analogue of the definition of a YY–base provided in [dBSS16, Definition 6.1]. However, we note that the more important concept for us is that of a computable prebase, which is related to the sequential bases in [dBSS16, Definition 7.1].
From a purely topological perspective, computable prebases have in particular the property that compact intersections are open in the topology generated by the prebase. Hence, by Corollary 14 computable prebases characterize the topology of their spaces up to sequentialization.
Corollary 16 (Computable prebases).
If $X$ is a represented space that has a computable prebase that generates a topology $\tau$, then ${\mathcal{O}}(X)={\mathrm{seq}}(\tau)$.
In the following we will need a couple of computable operations for the spaces ${\mathcal{O}}(X)$, ${\mathcal{A}}_{+}(X)$ and ${\mathcal{K}}_{-}(X)$, respectively. Most of these operations have been considered before (see [Pau16]). We recall that a subset $K\subseteq X$ of a topological space $X$ is called saturated if $K={\mathrm{sat}}(K):=\bigcap{\mathcal{F}}_{K}$. Here ${\mathrm{sat}}(K)$ is called the saturation of $K$.
Proposition 17 (Computable operations on hyperspaces).
Let $X,Y$ be represented spaces. Then the following hold:
  1. (1)
    ${\mathcal{U}}:X\to{\mathcal{O}}{\mathcal{O}}(X)$ is computable (neighborhood map)
  2. (2)
    ${\mathrm{inj}}:X\to{\mathcal{A}}_{+}(X),x\mapsto\overline{\{x\}}$ is computable (closed injection)
  3. (3)
    ${\mathrm{inj}}:X\to{\mathcal{K}}_{-}(X),x\mapsto{\mathrm{sat}}\{x\}$ is computable (compact injection)
  4. (4)
    ${\mathcal{K}}_{-}:{\mathcal{C}}(X,Y)\to{\mathcal{C}}({\mathcal{K}}_{-}(X),{\mathcal{K}}_{-}(Y)),f\mapsto(K\mapsto{\mathrm{sat}}f(K))$ (compact images)
  5. (5)
    ${\mathcal{A}}_{+}:{\mathcal{C}}(X,Y)\to{\mathcal{C}}({\mathcal{A}}_{+}(X),{\mathcal{A}}_{+}(Y)),f\mapsto(A\mapsto\overline{f(A)})$ (closed images)
  6. (6)
    $\sec:X\times{\mathcal{O}}(X\times Y)\to{\mathcal{O}}(Y),(x,U)\mapsto\{y\in Y:(x,y)\in U\}$ is computable (section)
  7. (7)
    $\times:{\mathcal{O}}(X)\times{\mathcal{O}}(Y)\to{\mathcal{O}}(X\times Y),(V,U)\mapsto V\times U$ is computable (product)
  8. (8)
    $\times:{\mathcal{A}}_{+}(X)\times{\mathcal{A}}_{+}(Y)\to{\mathcal{A}}_{+}(X\times Y),(A,B)\mapsto A\times B$ is computable (product)
  9. (9)
    $\bigcup:{\mathcal{A}}_{+}{\mathcal{O}}(X)\to{\mathcal{O}}(X),{\mathcal{A}}\mapsto\bigcup{\mathcal{A}}$ is computable (overt union of open)
  10. (10)
    $\bigcap:{\mathcal{K}}_{-}{\mathcal{O}}(X)\to{\mathcal{O}}(X),{\mathcal{K}}\mapsto\bigcap{\mathcal{K}}$ is computable (compact intersection of open)
  11. (11)
    $\bigcup:{\mathcal{A}}_{+}{\mathcal{A}}_{+}(X)\to{\mathcal{A}}_{+}(X),{\mathcal{A}}\mapsto\overline{\bigcup{\mathcal{A}}}$ is computable (overt union of overt)
  12. (12)
    $\bigcup:{\mathcal{K}}_{-}{\mathcal{K}}_{-}(X)\to{\mathcal{K}}_{-}(X),{\mathcal{K}}\mapsto{\mathrm{sat}}\bigcup{\mathcal{K}}$ is computable (compact union of compact)
  13. (13)
    ${\mathcal{F}}:{\mathcal{O}}(X)\to{\mathcal{K}}_{-}{\mathcal{O}}(X),U\mapsto{\mathcal{F}}_{U}$ is a computable embedding (compact filter)
  14. (14)
    ${\mathcal{F}}:{\mathcal{K}}_{-}(X)\to{\mathcal{O}}{\mathcal{O}}(X),K\mapsto{\mathcal{F}}_{K}$ is a computable embedding (open filter)
  15. (15)
    ${\mathcal{T}}:{\mathcal{A}}_{+}(X)\to{\mathcal{O}}{\mathcal{O}}(X),A\mapsto{\mathcal{T}}_{A}$ is a computable embedding (trace)
  16. (16)
    $\Box:{\mathcal{O}}(X)\to{\mathcal{O}}{\mathcal{K}}_{-}(X),U\mapsto\{K:K\subseteq U\}$ is a computable embedding (box)
  17. (17)
    $\Diamond:{\mathcal{O}}(X)\to{\mathcal{O}}{\mathcal{A}}_{+}(X),U\mapsto\{A:A\cap U\not$$\varnothing\}$ is a computable embedding (diamond)
  18. (18)
    ${\mathcal{C}}{\mathcal{O}}:{\mathcal{C}}(X,Y)\to{\mathcal{O}}({\mathcal{K}}_{-}(X)\times{\mathcal{O}}(Y)),f\mapsto\{(K,U)\in{\mathcal{K}}_{-}(X)\times{\mathcal{O}}(Y):f(K)\subseteq U\}$
    is computable (a computable embedding if $Y$ is computable Kolmogorov) (compact-open)
For arbitrary sets ${\mathcal{A}},{\mathcal{K}}\subseteq{\mathcal{O}}(X)$ we have $\bigcup{\mathcal{A}}=\bigcup\overline{{\mathcal{A}}}$ and $\bigcap{\mathcal{K}}=\bigcap{\mathrm{sat}}({\mathcal{K}})$.
Proof.
The results (1)–(10) can all be found in [Pau16], see Propositions 4.2, 5.5, 7.4, Section 9 and Corollaries 10.2 and 10.4 therein. We briefly discuss (11)–(18), which are all easy to see.
(13) We have ${\mathcal{F}}_{U}\subseteq{\mathcal{U}}\iff U\in{\mathcal{U}}$ and hence $\forall_{{\mathcal{F}}_{U}}({\mathcal{U}})=\chi_{\mathcal{U}}(U)$. Hence $\mathcal{F}$} is computable. We also have $x\in U\iff U\in{\mathcal{U}}_{x}$ and hence $\chi_{U}(x)=\forall_{{\mathcal{F}}_{U}}({\mathcal{U}}_{x})$, which proves that $\mathcal{F}$} is a computable embedding.
(14) We have $U\in{\mathcal{F}}_{K}\iff K\subseteq U$ and hence $\chi_{{\mathcal{F}}_{K}}(U)=\forall_{K}(U)$.
(15) We have $U\cap{\mathcal{T}}_{A}\not=\varnothing\iff A\cap U\not$$\varnothing$ and hence $\chi_{{\mathcal{T}}_{A}}(U)=\exists_{A}(U)$.
(16) We have $K\in\Box(U)\iff K\subseteq U\iff U\in{\mathcal{F}}_{K}$. Hence computability of $\mathcal{F}$} in (14) implies computability of $\Box$. We also have $x\in U\iff{\mathrm{sat}}\{x\}\subseteq U\iff{\mathrm{sat}}\{x\}\in\Box(U)$. Hence $\Box$ is a computable embedding as ${\mathrm{inj}}:X\to{\mathcal{K}}_{-}(X)$ is computable by (3).
(17) We have $A\in\Diamond U\iff A\cap U\not=\varnothing\iff U\in{\mathcal{T}}_{A}$. Hence computability of $\Diamond$ follows from (15). We also have $x\in U\iff\overline{\{x\}}\subseteq U\iff\overline{\{x\}}\in\Diamond U$. Hence $\Diamond$ is a computable embedding as ${\mathrm{inj}}:X\to{\mathcal{A}}_{+}(X),x\mapsto\overline{\{x\}}$ is computable by (2).
(11) We have $\overline{\bigcup{\mathcal{A}}}\cap U\not=\varnothing\iff\bigcup{\mathcal{A}}\cap U\not=\varnothing\iff(\exists A\in{\mathcal{A}})\;A\in\Diamond U\iff\exists_{\mathcal{A}}(\Diamond U)=1$. Hence (11) follows from (17).
(12) We have ${\mathrm{sat}}\bigcup{\mathcal{K}}\subseteq U\iff\bigcup{\mathcal{K}}\subseteq U\iff(\forall K\in{\mathcal{K}})\;K\in\Box U\iff\forall_{\mathcal{K}}(\Box U)=1$. Hence, (12) follows from (16).
(18) Since $f(K)\subseteq U\iff{\mathrm{sat}}f(K)\subseteq U$, computability of ${\mathcal{C}}{\mathcal{O}}$ follows from (1) and (4). If $Y$ is a computable Kolmogorov space, then ${\mathcal{U}}:Y\to{\mathcal{O}}{\mathcal{O}}(Y)$ is a computable embedding. Since $U\in{\mathcal{U}}_{f(x)}\iff f(x)\in U\iff{\mathrm{sat}}f\{x\}\subseteq U$, it follows that ${\mathcal{C}}{\mathcal{O}}$ is a computable embedding in this case too.
For arbitrary ${\mathcal{A}},{\mathcal{K}}\subseteq{\mathcal{O}}(X)$ we have
  • $x\in\bigcup\overline{{\mathcal{A}}}\iff\overline{{\mathcal{A}}}\cap{\mathcal{U}}_{x}\not=\varnothing\iff{\mathcal{A}}\cap{\mathcal{U}}_{x}\not=\varnothing\iff x\in\bigcup{\mathcal{A}}$ and
  • $x\in\bigcap{\mathrm{sat}}{\mathcal{K}}\iff{\mathrm{sat}}{\mathcal{K}}\subseteq{\mathcal{U}}_{x}\iff{\mathcal{K}}\subseteq{\mathcal{U}}_{x}\iff x\in\bigcap{\mathcal{K}}$,
which proves the additional remark. ∎
Often, we can establish that a computable presubbase is even a computable prebase because it is already closed under compact intersections.
Lemma 18 (Closure under compact intersections).
Let $X$ be a represented space with a computable presubbase $B:Y\to{\mathcal{O}}(X)$. If there is a computable $R:{\mathcal{K}}_{-}(Y)\rightrightarrows Y$} such that $\bigcap_{y\in K}B_{y}=B_{z}$ for every $z\in R(K)$, then $B$ is a computable prebase.
Proof.
This follows as ${\mathrm{inj}}:Y\to{\mathcal{A}}_{+}(Y),z\mapsto\overline{\{z\}}$ is computable and $\bigcup_{y\in\overline{\{z\}}}B_{y}=B_{z}$ by Proposition 17. ∎
Every computable presubbase can be converted into a computable prebase just by taking the closure under compact intersections.
Proposition 19 (Computable prebases).
Let $X$ be a represented space with a computable presubbase $B:Y\to{\mathcal{O}}(X)$. Then
$\cap_{\mathcal{K}}B:{\mathcal{K}}_{-}(Y)\to{\mathcal{O}}(X),K\mapsto\bigcap_{y\in K}B_{y},$
is a computable prebase with the understanding that the empty intersection is $X$.
Proof.
We use Proposition 17. Firstly, we note that $\cap_{\mathcal{K}}B(\varnothing)=X$. Next, we need to prove that $\cap_{\mathcal{K}}B$ is a computable presubbase, i.e., that
$(\cap_{\mathcal{K}}B)^{\mathsf{{T}}}:X\to{\mathcal{O}}{\mathcal{K}}_{-}(Y),x\mapsto\{K\in{\mathcal{K}}_{-}(Y):x\in\cap_{y\in K}B_{y}\}$
is well-defined and a computable embedding. Since $B^{\mathsf{{T}}}$ is a computable embedding, it follows that $B$ is computable and hence so is $\cap_{\mathcal{K}}B$. Hence $(\cap_{\mathcal{K}}B)^{\mathsf{{T}}}$ also is well-defined and computable. Now, given $(\cap_{\mathcal{K}}B)^{\mathsf{{T}}}_{x}$, we can compute $B^{\mathsf{{T}}}_{x}=\{y\in Y:x\in B_{y}\}\in{\mathcal{O}}(Y)$, as ${\mathrm{inj}}:Y\to{\mathcal{K}}_{-}(Y),y\mapsto{\mathrm{sat}}\{y\}$ is computable and the saturation is irrelevant as $\cap_{y\in K}B_{y}=\cap_{y\in{\mathrm{sat}}(K)}B_{y}$. That is $B^{\mathsf{{T}}}_{x}={\mathrm{inj}}^{-1}((\bigcap_{\mathcal{K}}B)^{\mathsf{{T}}}_{x})$. As $B^{\mathsf{{T}}}$ is a computable embedding, this implies that $(\cap_{\mathcal{K}}B)^{\mathsf{{T}}}$ is a computable embedding too.
Finally, we need to prove that $\cap_{\mathcal{K}}B$ is a computable prebase, i.e., that all compact intersections of it can be obtained computably as overt unions. To this end, we prove that $\cap_{\mathcal{K}}\cap_{\mathcal{K}}B:{\mathcal{K}}_{-}{\mathcal{K}}_{-}(Y)\to{\mathcal{O}}(X)$ is computable. Since $\bigcup:{\mathcal{K}}_{-}{\mathcal{K}}_{-}(Y)\to{\mathcal{K}}_{-}(Y),{\mathcal{K}}\mapsto{\mathrm{sat}}\bigcup{\mathcal{K}}$ is computable and
$\bigcap_{K\in{\mathcal{K}}}\bigcap_{y\in K}B_{y}=\bigcap_{y\in\bigcup{\mathcal{K}}}B_{y}=\bigcap_{y\in{\mathrm{sat}}(\bigcup{\mathcal{K}})}B_{y},$
the claim follows with Lemma 18. ∎
In fact, every computable Kolmogorov space has a computable base, namely simply the identity. This corresponds to the fact in classical topology that every topology is a base of itself.
Proposition 20 (Identity as a base).
Let $X$ be a represented space. Then the identity ${\mathrm{id}}:{\mathcal{O}}(X)\to{\mathcal{O}}(X)$ is a computable base/prebase/presubbase of $X$ if and only if $X$ is a computable Kolmogorov space.
Proof.
It follows from ${\mathrm{id}}^{\mathsf{{T}}}={\mathcal{U}}:X\to{\mathcal{O}}{\mathcal{O}}(X)$ that ${\mathrm{id}}:{\mathcal{O}}(X)\to{\mathcal{O}}(X)$ is a computable presubbase if and only if $X$ is a computable Kolmogorov space. If $\mathrm{id}$} is a computable presubbase, then it is automatically a computable base by Lemma 18 and Proposition 17. ∎
Altogether, we obtain the following characterization of computable Kolmogorov spaces in terms of their bases.
Theorem 21 (Computable Kolmogorov spaces and bases).
Let $X$ be a represented space. Then the following are pairwise equivalent:
  1. (1)
    $X$ is a computable Kolmogorov space.
  2. (2)
    $X$ has a computable presubbase.
  3. (3)
    $X$ has a computable prebase.
  4. (4)
    $X$ has a computable base.
  5. (5)
    ${\mathrm{id}}:{\mathcal{O}}(X)\to{\mathcal{O}}(X)$ is a computable base of $X$.
Proof.
[…]
By a represented ${\mathsf{{T}}}_{0}$ space we mean a represented space $X$ whose final topology is ${\mathsf{{T}}}_{0}$. We note that another consequence of the above results is that every represented ${\mathsf{{T}}}_{0}$ space $X$ can be converted into a computable Kolmogorov space. This is because ${\mathrm{id}}:{\mathcal{O}}(X)\to{\mathcal{O}}(X)$ is always a presubbase. In fact, this conversion preserves the topology and the equivalence class of its representation. For two representations $\delta_{1}$ of a set $X_{1}$ and $\delta_{2}$ of a set $X_{2}$ we write $\delta_{1}\leq\delta_{2}$ if $X_{1}\subseteq X_{2}$ and the representation $\delta_{1}$ can be computably reduced to $\delta_{2}$, i.e., if the identity ${\mathrm{id}}:X_{1}\to X_{2}$ is computable.
Theorem 22 (Represented spaces as computable Kolmogorov spaces).
Let $X$ be a represented space with a ${\mathsf{{T}}}_{0}$ topology ${\mathcal{O}}(X)$. Then we can endow $X$ with another representation that turns $X$ into a computable Kolmogorov space without changing the topology ${\mathcal{O}}(X)$ (and without changing the computable equivalence class of the representation of ${\mathcal{O}}(X)$).
Proof.
Let $(X,\delta)$ be a represented space with ${\mathsf{{T}}}_{0}$ topology ${\mathcal{O}}(X)$ and canonical representation $\delta^{\circ}$ of ${\mathcal{O}}(X)$ and $\delta^{\circ\circ}$ of ${\mathcal{O}}{\mathcal{O}}(X)$. Then ${\mathrm{id}}:{\mathcal{O}}(X)\to{\mathcal{O}}(X)$ is a presubbase of this space, as ${\mathrm{id}}^{\mathsf{{T}}}={\mathcal{U}}:X\to{\mathcal{O}}{\mathcal{O}}(X)$ is well-defined and injective. Hence, the presubbase representation $\delta^{\bullet}={\mathcal{U}}^{-1}\circ\delta^{\circ\circ}$ turns $X$ into a computable Kolmogorov space $X^{\bullet}$ by the Presubbase Theorem 7. Let $\delta^{\bullet\circ}$ denote the representation of the open sets ${\mathcal{O}}(X^{\bullet})$ induced by $\delta^{\bullet}$. We need to prove that ${\mathcal{O}}(X)={\mathcal{O}}(X^{\bullet})$ and that $\delta^{\circ}\equiv\delta^{\bullet\circ}$. Since $\mathcal{U}$} is computable, it is clear that $\delta\leq\delta^{\bullet}$ and hence $\delta^{\bullet\circ}\leq\delta^{\circ}$ and ${\mathcal{O}}(X^{\bullet})\subseteq{\mathcal{O}}(X)$. We need to prove the inverse reduction. If $x=\delta^{\bullet}(p)$ and $U=\delta^{\circ}(q)$ are given by $p,q\in{\mathbb{N}}^{\mathbb{N}}$, then we have $x\in U\iff U\in{\mathcal{U}}_{x}\iff\delta^{\circ}(q)\in\delta^{\circ\circ}(p)$ which can be confirmed with the help of $p,q\in{\mathbb{N}}^{\mathbb{N}}$. This shows $\delta^{\circ}\leq\delta^{\bullet\circ}$ and ${\mathcal{O}}(X)\subseteq{\mathcal{O}}(X^{\bullet})$. ∎
In fact, $\delta\mapsto\delta^{\bullet}$ is a closure operator in the lattice of representations of ${\mathsf{{T}}}_{0}$ spaces that was introduced and studied by Schröder [Sch02a]. We will get back to discussing this closure operator in the Epilogue.
4. Special Bases and Spaces
In this section we want to discuss some special types of bases and characterizations of special types of spaces by bases. The remainder of this article is independent of the results of this section. Firstly, we introduce two notions of bases that were considered in the countable case in [BR26]. A Lacombe base reflects the fact that open sets can be written as unions of base sets. A Nogina base reflects the fact that for open sets and points therein we can find base sets in between. We call a map $f:X\to Y$ computably surjective if it is computable, surjective and has a multivalued computable inverse $F:Y\rightrightarrows X$.
Definition 23 (Lacombe and Nogina bases).
Let $X,Y$ be represented spaces with a map $B:Y\to{\mathcal{O}}(X)$. Then
  1. (1)
    $B$ is called a computable Lacombe base of $X$ if
    $\bigcup B:{\mathcal{A}}_{+}(Y)\to{\mathcal{O}}(X),A\mapsto\bigcup_{y\in A}B_{y}$
    is computably surjective. We use $\bigcup_{y\in\varnothing}B_{y}=\varnothing$.
  2. (2)
    $B$ is called a computable Nogina base of $X$ if
    $N:\subseteq X\times{\mathcal{O}}(X)\rightrightarrows Y,(x,U)\mapsto\{y\in Y:x\in B_{y}\subseteq U\}$
    with ${\mathrm{dom}}(N):=\{(x,U):x\in U\}$ is well-defined and computable.
A definition similar to that of a Lacombe base has also been considered in a constructive setting before [BL12, Definition 6.3]. It is easy to see that the mere existence of computable Lacombe or Nogina bases is not a particularly interesting property as every represented space has such bases.
Proposition 24 (Identity as Lacombe and Nogina base).
Let $X$ be a represented space. Then the identity ${\mathrm{id}}:{\mathcal{O}}(X)\to{\mathcal{O}}(X)$ is a computable Lacombe base and a computable Nogina base of $X$.
Proof.
It is obvious that ${\mathrm{id}}:{\mathcal{O}}(X)\to{\mathcal{O}}(X)$ is a computable Nogina base and it is easy to see that it also is a computable Lacombe base. This follows from Proposition 17 since ${\mathrm{inj}}:{\mathcal{O}}(X)\to{\mathcal{A}}_{+}{\mathcal{O}}(X),U\mapsto\overline{\{U\}}$ is computable and $\bigcup\overline{\{U\}}=U$. ∎
It is more interesting to ask how different base properties for a given $B:Y\to{\mathcal{O}}(X)$ are logically related to each other. We will prove a number of implications, some of which only hold for specific types of spaces $Y$ or $X$. It is a consequence of Proposition 17 that every computable Lacombe base is a computable base, provided the space is a computable Kolmogorov space.
Proposition 25 (Computable Lacombe bases).
Let $X$ be a computable Kolmogorov space. Then every computable Lacombe base of $X$ is also a computable base.
Proof.
We use Proposition 17. If $\bigcup B:{\mathcal{A}}_{+}(Y)\to{\mathcal{O}}(X)$ is computable, then so is the map $B:Y\to{\mathcal{O}}(X)$. This is because ${\mathrm{inj}}:Y\to{\mathcal{A}}_{+}(Y),y\mapsto\overline{\{y\}}$ is computable and $\bigcup_{z\in\overline{\{y\}}}B_{z}=B_{y}$. We also have that $\bigcap:{\mathcal{K}}_{-}{\mathcal{O}}(X)\to{\mathcal{O}}(X)$ is computable and $\bigcup B:{\mathcal{A}}_{+}(Y)\to{\mathcal{O}}(X)$ has a computable right inverse $S:{\mathcal{O}}(X)\rightrightarrows{\mathcal{A}}_{+}(Y)$. Hence, given a set $K\in{\mathcal{K}}_{-}(Y)$ we can compute ${\mathcal{K}}:={\mathrm{sat}}B(K)={\mathrm{sat}}\{B_{y}:y\in K\}\in{\mathcal{K}}_{-}{\mathcal{O}}(X)$ and hence $\bigcap{\mathcal{K}}=\bigcap_{y\in K}B_{y}\in{\mathcal{O}}(X)$. Using $S$ we obtain $A\in{\mathcal{A}}_{+}(Y)$ with $\bigcap_{y\in K}B_{y}=\bigcup_{y\in A}B_{y}$. Altogether this describes a computation of a multivalued $R:{\mathcal{K}}_{-}(Y)\rightrightarrows{\mathcal{A}}_{+}(Y)$ as it is required for a computable prebase. Because $\bigcup B:{\mathcal{A}}_{+}(Y)\to{\mathcal{O}}(X)$ is surjective, $B$ is actually a base of $X$. It remains to prove that $B:Y\to{\mathcal{O}}(X)$ is a computable presubbase, i.e., that the transpose $B^{\mathsf{{T}}}:X\to{\mathcal{O}}(Y)$ is injective and has a computable right inverse. If the set $V=\{y\in Y:x\in B_{y}\}\in{\mathcal{O}}(Y)$ is given, then we can also compute the set ${\mathcal{U}}_{x}=\{U\in{\mathcal{O}}(X):x\in U\}\in{\mathcal{O}}{\mathcal{O}}(X)$. This is because given $U\in{\mathcal{O}}(X)$ we can compute $A\in{\mathcal{A}}_{+}(Y)$ with $U=\bigcup_{y\in A}B_{y}$ and hence $x\in U\iff(\exists y\in A)\;x\in B_{y}\iff A\cap V\not$$\varnothing$. But given ${\mathcal{U}}_{x}\in{\mathcal{O}}{\mathcal{O}}(X)$ we can compute $x\in X$, as $X$ is a computable Kolmogorov space. This completes the proof that $B:Y\to{\mathcal{O}}(X)$ is a computable base of $X$. ∎
This result generalizes [BR26, Proposition 4.4] for the second-countable case. We call a represented space computably second-countable, if it has a computable base $B:{\mathbb{N}}\to X$. The following characterization is well-known [BR26, Theorem 6.1].
Proposition 26 (Computable second-countability).
Let $X$ be a computable Kolmogorov space. Then $X$ is computably second-countable if and only if $X$ admits a computably equivalent computably open representation.
For such spaces every computable base $B:{\mathbb{N}}\to X$ is even a computable Lacombe base [BR26, Lemma 3.4]. We can expand the equivalence to a larger class of index spaces $Y$.
As a preparation, we need to study the representation $\delta_{{\mathcal{O}}(Y)}$ of open subsets of $Y$. If $q\in{\mathbb{N}}^{\mathbb{N}}$, then we write for short $n\in q:\iff n\in{\mathrm{range}}(q)-1\iff(\exists i)\;q(i)=n+1$, i.e., if $n$ is “enumerated” into $q$ (where 0 serves as a dummy symbol). We use an analogous notation $i\in w$ for words $w\in{\mathbb{N}}^{*}$. Let $\mathsf{w}_{i}$ be the ii–th word according to some standard numbering of ${\mathbb{N}}^{+}={\mathbb{N}}^{*}\setminus\{\varepsilon\}$, the set of words without the empty word $\varepsilon$.
Proposition 27 (Synthetic versus final representation).
Let $(Y,\delta_{Y})$ be a represented space. If we define a representation $\delta_{{\mathcal{O}}(Y)}$ of ${\mathcal{O}}(Y)$ by
$\delta_{{\mathcal{O}}(Y)}(q)=V:\iff\delta_{Y}^{-1}(V)={\mathrm{dom}}(\delta_{Y})\cap\bigcup_{j\in q}\mathsf{w}_{j}{\mathbb{N}}^{\mathbb{N}}$
for all $q\in{\mathbb{N}}^{\mathbb{N}}$ and $V\in{\mathcal{O}}(Y)$, then $\chi:{\mathcal{O}}(Y)\to{\mathcal{C}}(Y,{\mathbb{S}}),U\mapsto\chi_{U}$ is a computable isomorphism and hence $\delta_{{\mathcal{O}}(Y)}$ is computably equivalent to the synthetic representation of ${\mathcal{O}}(Y)$, defined via continuous characteristic functions $\chi_{U}:X\to{\mathbb{S}}$.
Proof.
Given $U\in{\mathcal{O}}(Y)$ by some $\delta_{{\mathcal{O}}(Y)}$–name $q$ and $y\in Y$ by some $\delta_{Y}$–name $p$, we can easily evaluate $\chi_{U}(y)$: we just search for some $j\in q$ with $\mathsf{w}_{j}\sqsubseteq p$ and if we find one, then we output $1$, otherwise 0. By currying, this proves that $\chi$ is computable. Now given $\chi_{U}\in{\mathcal{C}}(Y,{\mathbb{S}})$, we have a word function $f:{\mathbb{N}}^{*}\to{\mathbb{N}}^{*}$ that approximates a realizer of $\chi_{U}$. We just need to search for words $w$ such that $f(w)$ contains a $1$ and enumerate corresponding $j$ with $\mathsf{w}_{j}=w$ into $q$. If it happens that $w$ is the empty word, then we list instead all words $w$ of length $1$ in $q$. This proves that the inverse of $\chi$ is computable. ∎
We note that the representation $\delta_{{\mathcal{O}}(Y)}$ is not necessarily total, because the enumeration of certain words $\mathsf{w}_{j}$ in a name $q$ enforces the enumeration of other words $\mathsf{w}_{k}$ in the enumeration $q$ due to extensionality. However, every word $w\in{\mathbb{N}}^{*}$ can be extended to a name of $Y$ and hence $w{\mathbb{N}}^{\mathbb{N}}\cap{\mathrm{dom}}(\delta_{{\mathcal{O}}(Y)})\not$$\varnothing$. We now investigate this representation on cylinders.
Proposition 28 (Cylinder and filter).
Let $(Y,\delta_{Y})$ be a represented space and let $\delta_{{\mathcal{O}}(Y)}$ be defined as in Proposition 27. Then we obtain
$\delta_{{\mathcal{O}}(Y)}(w{\mathbb{N}}^{\mathbb{N}})={\mathcal{F}}_{P_{w}}\text{ with }P_{w}:=\bigcup_{j\in w}\delta_{Y}(\mathsf{w}_{j}{\mathbb{N}}^{\mathbb{N}})$
for each $w\in{\mathbb{N}}^{*}$, where ${\mathcal{F}}_{P_{w}}=\{U\in{\mathcal{O}}(Y):P_{w}\subseteq U\}$.
Proof.
For $q\in{\mathrm{dom}}(\delta_{{\mathcal{O}}(Y)})$ we obtain $V:=\delta_{{\mathcal{O}}(Y)}(q)=\bigcup_{j\in q}\delta_{Y}(\mathsf{w}_{j}{\mathbb{N}}^{\mathbb{N}})$. Hence
$\delta_{{\mathcal{O}}(Y)}(w{\mathbb{N}}^{\mathbb{N}})$ $=$ $\{\delta_{{\mathcal{O}}(Y)}(q):q\in w{\mathbb{N}}^{\mathbb{N}}\cap{\mathrm{dom}}(\delta_{{\mathcal{O}}(Y)})\}$
$=$ $\left\{\bigcup_{j\in q}\delta_{Y}(\mathsf{w}_{j}{\mathbb{N}}^{\mathbb{N}}):q\in w{\mathbb{N}}^{\mathbb{N}}\cap{\mathrm{dom}}(\delta_{{\mathcal{O}}(Y)})\right\}$
$=$ $\left\{\bigcup_{j\in w}\delta_{Y}(\mathsf{w}_{j}{\mathbb{N}}^{\mathbb{N}})\cup\bigcup_{j\in r}\delta_{Y}(\mathsf{w}_{j}{\mathbb{N}}^{\mathbb{N}}):wr\in w{\mathbb{N}}^{\mathbb{N}}\cap{\mathrm{dom}}(\delta_{{\mathcal{O}}(Y)})\right\}$
$=$ $\{V\in{\mathcal{O}}(Y):P_{w}\subseteq V\}$
$=$ ${\mathcal{F}}_{P_{w}}.$
Here $\supseteq$” in the second to last step holds, because if $r\in{\mathrm{dom}}(\delta_{{\mathcal{O}}(Y)})$, $w\in{\mathbb{N}}^{*}$ and $V:=\delta_{{\mathcal{O}}(Y)}(r)\supseteq P_{w}$, then $wr\in{\mathrm{dom}}(\delta_{{\mathcal{O}}(Y)})$ and $\delta_{{\mathcal{O}}(Y)}(wr)=V$. ∎
In order to make the dual behavior of a representation and its induced representation of open subsets formally precise, we define the following notion.
Definition 29 (Computably locally compact representations).
We call a representation $\delta_{Y}:\subseteq{\mathbb{N}}^{\mathbb{N}}\to Y$ of $Y$ computably locally compact if
$\lambda:{\mathbb{N}}^{+}\to{\mathcal{K}}_{-}(Y),w\mapsto\delta_{Y}(w{\mathbb{N}}^{\mathbb{N}})$
is well-defined and computable.
We recall that a representation $\delta_{Y}$ is called computably open if ${\mathcal{O}}({\mathbb{N}}^{\mathbb{N}})\to{\mathcal{O}}(Y),U\mapsto\delta_{Y}(U)$ is well-defined and computable, or equivalently, if ${\mathbb{N}}^{+}\to{\mathcal{O}}(Y),w\mapsto\delta_{Y}(w{\mathbb{N}}^{\mathbb{N}})$ is well-defined and computable. Now we can formulate the following duality, where we use the dual computability properties of the filter maps $U\mapsto{\mathcal{F}}_{U}$ and $K\mapsto{\mathcal{F}}_{K}$ from Proposition 17 together with the fact that finite unions can be computed for open and compact sets.
Corollary 30 (Duality).
Let $Y$ be a represented space. Then the following hold:
  1. (1)
    If $Y$ has a computably equivalent computably open representation, then ${\mathcal{O}}(Y)$ has a computably equivalent computably locally compact representation.
  2. (2)
    If $Y$ has a computably equivalent computably locally compact representation, then ${\mathcal{O}}(Y)$ has a computably equivalent computably open representation.
We note that this duality is related to some purely topological characterizations in [Sch04, Sch16]. Now we want to discuss implications of these observations for bases.
Proposition 31 (Normal form).
Let $(X,\delta_{X})$ and $(Y,\delta_{Y})$ be represented spaces with a presubbase $B:Y\to{\mathcal{O}}(X)$ and $\delta_{{\mathcal{O}}(X)},\delta_{{\mathcal{O}}(Y)}$ defined as in Proposition 27, where $\delta_{X}=\delta^{B}$ is the presubbase representation of $B$. Then we obtain
$U=\bigcup_{i\in p}\bigcap_{y\in K_{\mathsf{w}_{i}}}B_{y}.$
for each $p\in{\mathrm{dom}}(\delta_{{\mathcal{O}}(X)})$ with $U:=\delta_{{\mathcal{O}}(X)}(p)$ and sets $K_{w}:=\bigcup_{j\in w}\delta_{Y}(\mathsf{w}_{j}{\mathbb{N}}^{\mathbb{N}})$ for $w\in{\mathbb{N}}^{*}$. In particular, if $\delta_{Y}$ is computably locally compact and $B$ is a computable base, then $B$ is a Lacombe base.
Proof.
Let $\delta_{{\mathcal{O}}(X)}(p)=U$. With $\delta_{X}=(B^{\mathsf{{T}}})^{-1}\circ\delta_{{\mathcal{O}}(Y)}$ we obtain with Proposition 28
$U=\delta_{X}\left(\bigcup_{i\in p}\mathsf{w}_{i}{\mathbb{N}}^{\mathbb{N}}\right)=\bigcup_{i\in p}(B^{\mathsf{{T}}})^{-1}\circ\delta_{{\mathcal{O}}(Y)}(\mathsf{w}_{i}{\mathbb{N}}^{\mathbb{N}})=\bigcup_{i\in p}(B^{\mathsf{{T}}})^{-1}({\mathcal{F}}_{K_{\mathsf{w}_{i}}})=\bigcup_{i\in p}\bigcap_{y\in K_{\mathsf{w}_{i}}}B_{y}.$
If $\delta_{Y}$ is computably locally compact, then we can compute the sets $K_{\mathsf{w}_{i}}\in{\mathcal{K}}_{-}(Y)$ and if $B$ is a computable base, then we can compute $A_{i}\in{\mathcal{K}}_{-}(Y)$ uniformly in $i$ such that $\bigcup_{y\in A_{i}}B_{y}=\bigcap_{y\in K_{\mathsf{w}_{i}}}B_{y}$. Finally, from the sets $A_{i}$ we can compute the union $A:=\overline{\bigcup_{i\in p}A_{i}}$ in ${\mathcal{A}}_{+}(Y)$ by Proposition 17 and we obtain
$\bigcup_{y\in A}B_{y}=\bigcup_{y\in\bigcup_{i\in p}A_{i}}B_{y}=\bigcup_{i\in p}\bigcup_{y\in A_{i}}B_{y}=\bigcup_{i\in p}\bigcap_{y\in K_{\mathsf{w}_{i}}}B_{y}=U.$
This proves that $B$ is a computable Lacombe base. ∎
This result allows us to formulate a computable base theorem for a class of bases with more general index space $Y$. We say that $Y$ is computably locally compactly represented if $Y$ has a computably equivalent representation that is computably locally compact. Some obvious examples of such spaces are $\mathbb{N}$}, ${\mathbb{N}}\times 2^{\mathbb{N}}$ and ${\mathbb{R}}^{n}$. However, we note that computably locally compactly represented spaces are not necessarily locally compact themselves. Together with Propositions 25 and 31 we obtain the following.
Theorem 32 (Base theorem).
Let $Y$ be computably locally compactly represented space and $X$ a computable Kolmogorov space with a map $B:Y\to{\mathcal{O}}(X)$. Then the following are equivalent:
  1. (1)
    $B:Y\to{\mathcal{O}}(X)$ is a computable base of $X$.
  2. (2)
    $B:Y\to{\mathcal{O}}(X)$ is a computable Lacombe base of $X$.
However, this result looks perhaps more useful than it is. While we have extended the base theorem from the case of $Y={\mathbb{N}}$ to a larger class of bases $B:Y\to{\mathcal{O}}(X)$ with more general index spaces $Y$, we have not extended it to a larger class of spaces $X$. This is because of the following result.
Theorem 33 (Computable second-countability).
Let $X$ be a computable Kolmogorov space. Then the following are equivalent:
  1. (1)
    $X$ is computably second-countable.
  2. (2)
    $X$ admits a computable base $B:{\mathbb{N}}\to{\mathcal{O}}(X)$.
  3. (3)
    $X$ admits a computable base $B:Y\to{\mathcal{O}}(X)$ with a computably locally compactly represented space $Y$.
  4. (4)
    ${\mathcal{O}}(X)$ is computably locally compactly represented.
Proof.
The condition (2) is our definition of (1). Clearly, (2) implies (3), as $\mathbb{N}$} is computably locally compactly represented. Also (3) implies (1): If $Y$ is computably locally compactly represented, and $B:Y\to{\mathcal{O}}(X)$ is a computable base of $X$, then $B^{\mathsf{{T}}}:X\to{\mathcal{O}}(Y)$ is a computable embedding. By Corollary 30 the space ${\mathcal{O}}(Y)$ has a computably equivalent computably open representation. This implies that ${\mathcal{O}}(Y)$ is computably second-countable by Proposition 26 (${\mathcal{O}}(Y)$ is always a computable Kolmogorov space). Since $X$ can be computably embedded into ${\mathcal{O}}(Y)$, it is also computably second-countable. Finally, (1) implies (4) by Corollary 30 and Proposition 26, and (4) implies (3) by Proposition 20 . ∎
This theorem characterizes the scope of spaces that have computable bases with computably locally compactly represented index space $Y$. On the first sight, it seems to make Theorem 32 less useful. On the other hand, we will see in Example 42 that even for computably second-countable spaces $X$ Theorem 32 does not hold for arbitrary index spaces $Y$.
Nogina bases seem to be less suitable for a general treatment of computable topology, as they refer to points, which might not be easily accessible. However, if the index space $Y$ is sufficiently nice, then every Lacombe base is a Nogina base. If the space $X$ itself is nice, then every Nogina base yields a Lacombe base in a canonical way.
In order to express this canonical representation, we denote for every represented space $(X,\delta)$ by $(X_{\bot},\delta_{\bot})$ a version of the space to which we attach a bottom element $\bot\not\in X$. That is $X_{\bot}:=X\cup\{\bot\}$ and $\delta_{\bot}$ is defined by
$\delta_{\bot}(\widehat{0}):=\bot\text{ and }\delta_{\bot}(0^{n}1p):=\delta(p)$
for all $p\in{\mathrm{dom}}(\delta)$ and $n\in{\mathbb{N}}$. In other words, the element $\bot$ is attached to $X$ such that ${\mathrm{id}}:X\to X_{\bot}$ is a computable embedding and $X$ is a c.e. open subspace of $X_{\bot}$. This yields the least possible one-point compactification of $X$, and its topology is sometimes called the open extension topology. For instance, $\{1\}_{\bot}$ is computably isomorphic to $\mathbb{S}$}.
For every map $B:Y\to{\mathcal{O}}(X)$ we define $B_{\bot}:Y_{\bot}\to{\mathcal{O}}(X)$ as extension of $B$ with $B(\bot)=\varnothing$. It is clear that $B_{\bot}$ is computable if $B$ is computable, as we can map the input to a prefix of a name of $\varnothing$ until it is clear that the input is different from $\bot$.
computable basecomputable Nogina basecomputable Lacombe base $Y$ computably
quasi-Polish
$Y$ computably locally
compactly represented
Figure 2. Bases $B:Y\to{\mathcal{O}}(X)$ for computable Kolmogorov spaces $X$.
We say that a represented space $X$ is computably complete, if its representation is computably equivalent to a total representation $\delta:{\mathbb{N}}^{\mathbb{N}}\to X$. And a represented space $X$ is called a computable quasi-Polish space, if it is complete and is computable second-countable (i.e., admits a computable basis $B:{\mathbb{N}}\to X$).
Proposition 34 (Nogina and Lacombe bases).
Let $X$ and $Y$ be represented spaces.
  1. (1)
    If $B:Y\to{\mathcal{O}}(X)$ is a computable Lacombe base of $X$ and $Y$ is a computable quasi-Polish space, then $B$ is a computable Nogina base of $X$.
  2. (2)
    If $B:Y\to{\mathcal{O}}(X)$ is a computable Nogina base of $X$ and $X$ is a complete space, then $B_{\bot}:Y_{\bot}\to{\mathcal{O}}(X)$ is a computable Lacombe base of $X$.
Proof.
(1) Let $B:Y\to{\mathcal{O}}(X)$ be a computable Lacombe base of $X$. Then given some $U\in{\mathcal{O}}(X)$ we can compute some $A\in{\mathcal{A}}_{+}(Y)$ such that $U=\bigcup_{y\in A}B_{y}$. Since $Y$ is a computable quasi-Polish space, we can compute some sequence $(y_{n})_{n}$ in $Y$ with $A=\overline{\{y_{n}:n\in{\mathbb{N}}\}}$ by [BPS20, Corollary 21], which implies $U=\bigcup_{n\in{\mathbb{N}}}B_{y_{n}}$ (using the statement at the bottom of Proposition 17). Now given some $x\in X$ with $x\in U$, we can compute some $n\in{\mathbb{N}}$ with $x\in B_{y_{n}}\subseteq U$. Hence, $B$ is a computable Nogina base of $X$.
(2) Let $B:Y\to{\mathcal{O}}(X)$ be a computable Nogina base and let $X$ be complete. We can extend the Nogina operation $N$ to a total computable operation
$N:X\times{\mathcal{O}}(X)\rightrightarrows Y_{\bot},(x,U)\mapsto\left\{\begin{array}[]{ll}\{y\in Y:x\in B_{y}\subseteq U\}&\text{if $x\in U$}\\ \bot&\text{otherwise}\end{array}\right.$.
Hence, there is a computable function $F:\subseteq{\mathbb{N}}^{\mathbb{N}}\times{\mathbb{N}}^{\mathbb{N}}\to Y_{\bot}$ such that for every name $p$ of a point $x\in X$ and every name $q$ of an open $U\in{\mathcal{O}}(X)$ the value $F(p,q)$ satisfies $x\in B_{\bot}(F(p,q))\subseteq U$ if $x\in U$ and $B_{\bot}(F(p,q))=\varnothing$ otherwise. Since $X$ is complete (and hence has a total representation), we obtain by currying a computable function $G:\subseteq{\mathbb{N}}^{\mathbb{N}}\to{\mathcal{C}}({\mathbb{N}}^{\mathbb{N}},Y_{\bot})$. Hence, given a name $q$ of an open set $U\in{\mathcal{O}}(X)$, we can compute $A:=\overline{G(q)({\mathbb{N}}^{\mathbb{N}})}\in{\mathcal{A}}_{+}(Y_{\bot})$ by Proposition 17 and we obtain
$\bigcup_{y\in A}B_{\bot}(y)=\bigcup_{y\in F({\mathbb{N}}^{\mathbb{N}}\times\{q\})}B_{\bot}(y)=U.$
Hence, $B_{\bot}:Y_{\bot}\to{\mathcal{O}}(X)$ is a computable Lacombe base of $X$. ∎
We could weaken the condition on completeness of $X$ even to represented spaces $X$ that admit a representation with an overt domain. In Figure 2 some of the implications between types of bases are displayed.
Next we want to discuss how overt spaces can be characterized in terms of bases. We recall that a represented space $X$ is called overt if $X$ is a computable point in ${\mathcal{A}}_{+}(X)$, or in other words, if ${\mathcal{O}}(X)^{\times}:=\{U\in{\mathcal{O}}(X):U\not$$\varnothing\}$ is a c.e. open subspace of ${\mathcal{O}}(X)$. By ${\mathrm{id}}^{\times}:{\mathcal{O}}(X)^{\times}\to{\mathcal{O}}(X)$ we denote the identity restricted to ${\mathcal{O}}(X)^{\times}$. We also need to following natural generalization of the finite intersection property.
Definition 35 (Compact intersection property).
Let $X$ be a represented space.
  1. (1)
    We say that $X$ has the compact intersection property, if $\bigcap K\not$$\varnothing$ for all $K\in{\mathcal{K}}_{-}({\mathcal{O}}(X)^{\times})$ implies $\bigcap{\mathcal{O}}(X)^{\times}\not$$\varnothing$.
  2. (2)
    We say that $X$ has the computable compact intersection property, if $\bigcap K\not$$\varnothing$ for all computable $K\in{\mathcal{K}}_{-}({\mathcal{O}}(X)^{\times})$ implies that there is a computable $x\in\bigcap{\mathcal{O}}(X)^{\times}$.
We note that the compact intersection property is satisfied by many spaces, among them all sober spaces. This is because the premise of the compact intersection property implies that all pairs of non-empty open subsets of $X$ intersect, which means that $X$ itself is irreducible. If it is sober, then it follows that $X=\overline{\{x\}}$, which shows the conclusion is satisfied. On the other hand, if $X$ is not-sober, then $X\sqcup\{0\}$ is also not sober and yet satisfies the compact intersection property, because the premise is not satisfied.
Now we can prove a characterization of overt Kolmogorov spaces $X$ via certain bases that avoid the empty set. We say that a map $B:Y\to{\mathcal{O}}(X)$ avoids the empty set, if ${\mathrm{range}}(B)\subseteq{\mathcal{O}}(X)^{\times}$.
Theorem 36 (Overt spaces and bases).
Let $X$ be a computable Kolmogorov space. Then the following are pairwise equivalent:
  1. (1)
    $X$ is overt.
  2. (2)
    $X$ has a computable Lacombe base that avoids the empty set.
  3. (3)
    ${\mathrm{id}}^{\times}:{\mathcal{O}}(X)^{\times}\to{\mathcal{O}}(X)$ is a computable Lacombe base of $X$.
If $X$, additionally, satisfies the computable compact intersection property, then the following condition is equivalent to each of the aforementioned ones:
  1. (4)
    ${\mathrm{id}}^{\times}:{\mathcal{O}}(X)^{\times}\to{\mathcal{O}}(X)$ is a computable base of $X$.
Proof.
[…]
$U\not=\varnothing\iff A\not=\varnothing\iff A\cap Y\not$$\varnothing$
and the latter property is c.e. in $A$ since $Y$ is c.e. open in itself. Hence $X$ is overt.
[…]
$A:=\left\{\begin{array}[]{ll}\overline{\{U\}}&\mbox{if $U\not=\varnothing$}\\ \varnothing&\mbox{otherwise}\end{array}\right.\in{\mathcal{A}}_{+}({\mathcal{O}}(X)^{\times})$
as overtness of $X$ implies that given ${\mathcal{U}}\in{\mathcal{O}}({\mathcal{O}}(X)^{\times})$ the following condition is c.e.:
$A\cap{\mathcal{U}}\not=\varnothing\iff U\not=\varnothing\text{ and }U\in{\mathcal{U}}.$
We obtain $\bigcup_{V\in A}{\mathrm{id}}^{\times}(V)=U$ and hence $\mathrm{id}$$^{\times}$ is a computable Lacombe base.
[…]
$\forall_{K\vee U}({\mathcal{U}})\iff(\forall W\in K\vee U)\,W\in{\mathcal{U}}\iff(\forall V\in K)\,V\cup U\in{\mathcal{U}}\iff\forall_{K}(f_{U}^{-1}({\mathcal{U}})).$
Finally, ${\mathcal{O}}(X)\to{\mathcal{K}}_{-}({\mathcal{O}}(X)^{\times}),U\mapsto K\vee U$ is computable by Proposition 17, as the function $U\mapsto{\mathcal{C}}({\mathcal{O}}(X)^{\times},{\mathcal{O}}(X)^{\times}),U\mapsto f_{U}$ is computable. ∎
It is obvious that ${\mathrm{id}}^{\times}:{\mathcal{O}}(X)^{\times}\to{\mathcal{O}}(X)$ is a computable Nogina base for every represented space. Hence, the existence of such a base cannot be used to characterize overt spaces. It is also easy to see that every space $X$ with two non-empty disjoint c.e. open subsets $U_{1},U_{2}\in X$ satisfies the computable compact intersection property (since $K={\mathrm{sat}}\{U_{1},U_{2}\}\in{\mathcal{K}}_{-}({\mathcal{O}}(X)^{\times})$ is computable then). This observation yields the following negative result.
Corollary 37.
If $X$ is a non-overt computable Kolmogorov space with two disjoint non-empty c.e. open subsets, then ${\mathrm{id}}^{\times}:{\mathcal{O}}(X)^{\times}\to{\mathcal{O}}(X)$ is a computable Nogina base, which is neither a computable Lacombe base nor a computable base.
Every non-c.e. subspace $X\subseteq{\mathbb{N}}$ is an example of a non-overt computable Kolmogorov space with two non-empty c.e. open subsets. Also computable subbases that avoid the empty set cannot be used to characterize overtness. This holds even in the second-countable case with $Y={\mathbb{N}}$.
Example 38.
Let $a,b\in A\subseteq{\mathbb{N}}$ with $a\not=b$ and $B_{2n}:=\{a,n\}$, $B_{2n+1}:=\{b,n\}$ for all $n\in{\mathbb{N}}$. Then $B:{\mathbb{N}}\to{\mathcal{O}}({\mathbb{N}}),n\mapsto B_{n}$ is a computable base of $\mathbb{N}$} and hence we obtain a computable subbase $B^{\prime}:{\mathbb{N}}\to{\mathcal{O}}(A),n\mapsto B_{n}\cap A$ of $A$ (see also Proposition 46). Clearly, $B^{\prime}$ avoids the empty set, but $A$ is overt if and only if it is c.e.
In the second-countable case, computable bases $B:{\mathbb{N}}\to{\mathcal{O}}(X)$ that avoid the empty set can be used to characterize overt spaces [BR26, Proposition 6.5]. We prove that this does not hold for arbitrary bases. In fact, we will show that the compact intersection property in Theorem 36 was in some sense not just sufficient but also necessary. We proceed in several steps.
Lemma 39.
Let $X$ be a computable Kolmogorov space that satisfies $\bigcap{\mathcal{O}}(X)^{\times}=\varnothing$. Then ${\mathrm{id}}^{\times}:{\mathcal{O}}(X)^{\times}\to{\mathcal{O}}(X)$ is a computable presubbase of $X$.
Proof.
Since $X$ is a computable Kolmogorov space, it follows that the neighborhood map ${\mathcal{U}}:X\to{\mathcal{O}}{\mathcal{O}}(X)$ is a computable embedding. We need to prove that
${\mathcal{U}}^{\times}:X\to{\mathcal{O}}({\mathcal{O}}(X)^{\times}),x\mapsto{\mathcal{U}}^{\times}_{x}:=\{U\in{\mathcal{O}}(X)^{\times}:x\in U\}$
is a computable embedding too. We first note that ${\mathcal{O}}(X)^{\times}$ is an open and hence sequential subspace of ${\mathcal{O}}(X)$. Thus, ${\mathcal{O}}({\mathcal{O}}(X)^{\times})=\{{\mathcal{V}}\cap{\mathcal{O}}(X)^{\times}:{\mathcal{V}}\in{\mathcal{O}}{\mathcal{O}}(X)\}$. Since the identity $\iota:{\mathcal{O}}(X)^{\times}\to{\mathcal{O}}(X)$} is a computable embedding, we obtain that ${\mathcal{U}}^{\times}=\iota^{-1}\circ{\mathcal{U}}$ is computable too. We need to show that the partial inverse of $\mathcal{U}$$^{\times}$ is computable. We consider the representation $\delta_{{\mathcal{O}}({\mathcal{O}}(X)^{\times})}$ according to Proposition 27. Let $q\in{\mathbb{N}}^{\mathbb{N}}$ be such that $\delta_{{\mathcal{O}}({\mathcal{O}}(X)^{\times})}(q)={\mathcal{U}}^{\times}_{x}$ for some $x\in X$ and let $j\in q$. Then for every $U\in{\mathcal{O}}(X)^{\times}$
$U\in\delta_{{\mathcal{O}}(X)^{\times}}(\mathsf{w}_{j}{\mathbb{N}}^{\mathbb{N}})=\delta_{{\mathcal{O}}(X)}(\mathsf{w}_{j}{\mathbb{N}}^{\mathbb{N}})\cap{\mathcal{O}}(X)^{\times}\Longrightarrow x\in U.$
Since $\bigcap{\mathcal{O}}(X)^{\times}=\varnothing$, there is some $U\in{\mathcal{O}}(X)^{\times}$ with $x\not\in U$ and hence $U\not\in\delta_{{\mathcal{O}}(X)^{\times}}(\mathsf{w}_{j}{\mathbb{N}}^{\mathbb{N}})$. But this implies $\varnothing\not\in\delta_{{\mathcal{O}}(X)}(\mathsf{w}_{j}{\mathbb{N}}^{\mathbb{N}})$, since otherwise $\delta_{{\mathcal{O}}(X)}(\mathsf{w}_{j}{\mathbb{N}}^{\mathbb{N}})={\mathcal{O}}(X)$ would follow. But if $\varnothing\not\in\delta_{{\mathcal{O}}(X)}(\mathsf{w}_{j}{\mathbb{N}}^{\mathbb{N}})$ for all $j\in q$, then $q\in{\mathrm{dom}}(\delta_{{\mathcal{O}}{\mathcal{O}}(X)})$ and $\delta_{{\mathcal{O}}{\mathcal{O}}(X)}(q)={\mathcal{U}}_{x}$. But this means that given a name $q$ for ${\mathcal{U}}^{\times}_{x}$, we actually already have a name for ${\mathcal{U}}_{x}$ and hence we can compute $x\in X$ and $\mathcal{U}$$^{\times}$ is a computable embedding. ∎
From this result it follows that we even obtain a computable base if the space does not satisfy the compact intersection property.
Proposition 40.
If $X$ is a computable Kolmogorov space that does not satisfy the compact intersection property, then ${\mathrm{id}}^{\times}:{\mathcal{O}}(X)^{\times}\to{\mathcal{O}}(X)$ is a computable base of $X$.
Proof.
That $X$ does not satisfy the compact intersection property implies that $\bigcap K\not$$\varnothing$ for all $K\in{\mathcal{K}}_{-}({\mathcal{O}}(X)^{\times})$ and yet $\bigcap{\mathcal{O}}(X)^{\times}=\varnothing$. The latter condition implies that $\mathrm{id}$$^{\times}$ is a computable presubbase by Lemma 39. We prove that $\mathrm{id}$$^{\times}$ is even a computable base. We claim that the map $\bigcap:{\mathcal{K}}_{-}({\mathcal{O}}(X)^{\times})\to{\mathcal{O}}(X)^{\times},K\mapsto\bigcap K$ is well-defined and computable. Then it follows by Lemma 18 that $\mathrm{id}$$^{\times}$ is a computable base (since it is obviously also a base). We now prove the claim. Firstly, by the non-emptiness assumption we have $\bigcap K\not$$\varnothing$ for all $K\in{\mathcal{K}}_{-}({\mathcal{O}}(X)^{\times})$. Hence, for the well-definedness we still need to show that $\bigcap K$ is open, which we prove together with the computability of the map. We obtain for $K\in{\mathcal{K}}_{-}({\mathcal{O}}(X)^{\times})$
$x\in\bigcap K\iff(\forall U\in K)\,x\in U\iff(\forall U\in K)\,U\in{\mathcal{U}}^{\times}_{x}\iff\forall_{K}({\mathcal{U}}^{\times}_{x})=1,$
i.e., $\chi_{\bigcap K}=\forall_{K}\circ{\mathcal{U}}^{\times}$. This proves that $\bigcap:{\mathcal{K}}_{-}({\mathcal{O}}(X)^{\times})\to{\mathcal{O}}(X)$ is computable and well-defined. As the identity $\iota:{\mathcal{O}}(X)^{\times}\to{\mathcal{O}}(X)$ is a computable embedding, we can also compute $\bigcap$ with ${\mathcal{O}}(X)^{\times}$ as output space, as claimed. ∎
Finally, together with Proposition 36 we obtain the desired sufficient criterion for $\mathrm{id}$$^{\times}$ being a computable base, but not a computable Lacombe base.
Corollary 41.
If $X$ is a computable Kolmogorov space that is neither overt, nor satisfies the compact intersection property, then ${\mathrm{id}}^{\times}:{\mathcal{O}}(X)^{\times}\to{\mathcal{O}}(X)$ is a computable base, but not a computable Lacombe base.
We still need to prove that there are actually spaces that satisfy the requirements given in this corollary. The natural numbers with the cofinite topology are a natural example of a space that does not satisfy the compact intersection property. However, this space is overt. But a non-overt version of this space yields a counterexample. And the space $X$ in this counterexample is even computably second-countable.
Example 42 (Non-overt cofinite space of natural numbers).
We consider the space $X={\mathbb{N}}$ with the cofinite topology, induced by the base $B:{\mathbb{N}}\to{\mathcal{O}}(X)$ with
  1. (1)
    $B_{2n}:={\mathbb{N}}\setminus{\mathrm{range}}(\mathsf{w}_{n})$,
  2. (2)
    $B_{2n+1}:=\varnothing$ if $n\not\in A$,
  3. (3)
    $B_{2n+1}:={\mathbb{N}}$ if $n\in A$
for all $n\in{\mathbb{N}}$, where $A\subseteq{\mathbb{N}}$ is some set that is not c.e. and $\mathsf{w}:{\mathbb{N}}\to{\mathbb{N}}^{*}$ is a standard numbering of words. We endow $X$ with the subbase representation $\delta^{B}$ induced by this subbase. Then $X$ is a computably second-countable computable Kolmogorov space that is not overt and does not satisfy the compact intersection property. Hence, ${\mathrm{id}}^{\times}:{\mathcal{O}}(X)^{\times}\to{\mathcal{O}}(X)$ is a computable base of $X$ that is not a computable Lacombe base.
Proof.
It is clear that $X$ is a computably second-countable computable Kolmogorov space. Then $B:{\mathbb{N}}\to{\mathcal{O}}(X)$ is computable for this space and $B_{2n+1}\not=\varnothing\iff n\in A$. Hence $X$ is not overt. We have ${\mathcal{O}}(X)^{\times}=\{U\subseteq{\mathbb{N}}:{\mathbb{N}}\setminus U\text{ finite}\}$. Hence, $X$ does not satisfy the finite intersection property in the sense that $U_{1}\cap U_{2}\cap...\cap U_{n}\not$$\varnothing$ for any choice of finitely many $U_{1},...,U_{n}\in{\mathcal{O}}(X)^{\times}$, but $\bigcap{\mathcal{O}}(X)^{\times}=\varnothing$. Hence, by Corollary 41 it suffices to show that ${\mathcal{K}}_{-}({\mathcal{O}}(X)^{\times})$ contains only finite sets. Since ${\mathcal{O}}(X)^{\times}$ is an open subset of ${\mathcal{O}}(X)$, it is a sequential subspace and hence ${\mathcal{O}}({\mathcal{O}}(X)^{\times})=\{{\mathcal{U}}\cap{\mathcal{O}}(X)^{\times}:{\mathcal{U}}\in{\mathcal{O}}{\mathcal{O}}(X)\}$. Since every subset of $X$ is compact, in particular, the sets $B_{2n}$ are compact and hence
${\mathcal{U}}_{n}:=\{U\in{\mathcal{O}}(X)^{\times}:B_{2n}\subseteq U\}={\mathcal{F}}_{B_{2n}}\cap{\mathcal{O}}(X)^{\times}\in{\mathcal{O}}({\mathcal{O}}(X)^{\times})$
for all $n\in{\mathbb{N}}$. The sets ${\mathcal{U}}_{n}$ form an open cover of ${\mathcal{O}}(X)^{\times}$ and hence every compact set $K\in{\mathcal{K}}_{-}({\mathcal{O}}(X)^{\times})$ has a finite subcover of the form $K\subseteq{\mathcal{U}}_{1}\cup...\cup{\mathcal{U}}_{m}$ for some $m\in{\mathbb{N}}$. This implies that $K$ is finite since all the ${\mathcal{U}}_{n}$ are finite. This completes the proof. ∎
5. Closure properties
In this section we provide a number of examples of computable presubbases and prebases and we demonstrate that these concepts can be used to derive interesting results. All these results are essentially well-known (mostly due to the work of Schröder [Sch02a]). The purpose here is not to claim originality, but to show how these results can be easily derived using the concepts of bases. We start with providing a number of examples that show that computable presubbases of hyper and function spaces occur very naturally.
Proposition 43 (Computable presubbases).
Let $X$ be a represented space and $Y$ a computable Kolmogorov space. The following are computable presubbases:
  1. (1)
    ${\mathcal{U}}:X\to{\mathcal{O}}{\mathcal{O}}(X),x\mapsto\{U\in{\mathcal{O}}(X):x\in U\}$.
  2. (2)
    ${\mathcal{F}}:{\mathcal{K}}_{-}(X)\to{\mathcal{O}}{\mathcal{O}}(X),K\mapsto\{U\in{\mathcal{O}}(X):K\subseteq U\}$.
  3. (3)
    $\Box:{\mathcal{O}}(X)\to{\mathcal{O}}{\mathcal{K}}_{-}(X),U\mapsto\{K\in{\mathcal{K}}_{-}(X):K\subseteq U\}$.
  4. (4)
    $\Diamond:{\mathcal{O}}(X)\to{\mathcal{O}}{\mathcal{A}}_{+}(X),U\mapsto\{A\in{\mathcal{A}}_{+}(X):A\cap U\not$$\varnothing\}$.
  5. (5)
    $\triangleright:{\mathcal{K}}_{-}(X)\times{\mathcal{O}}(Y)\to{\mathcal{O}}{\mathcal{C}}(X,Y),(K,U)\mapsto\{f\in{\mathcal{C}}(X,Y):f(K)\subseteq U\}$.
The maps $\mathcal{F}$} and $\Box$ are even computable prebases.
Proof.
We have ${\mathcal{U}}^{\mathsf{{T}}}={\mathrm{id}}_{{\mathcal{O}}(X)}$, $\Box^{\mathsf{{T}}}={\mathcal{F}}$, $\Diamond^{\mathsf{{T}}}={\mathcal{T}}$ and $\triangleright^{\mathsf{{T}}}={\mathcal{C}}{\mathcal{O}}$ where the maps $\mathcal{F}$}, $\mathcal{T}$}, ${\mathcal{C}}{\mathcal{O}}$ and $\Box$ are computable embeddings according to Proposition 17. Hence, $\mathcal{U}$}, $\mathcal{F}$}, $\Box$, $\Diamond$ and $\triangleright$ are all computable presubbases.
We still need to prove that $\mathcal{F}$} and $\Box$ are even computable prebases. Given ${\mathcal{K}}\in{\mathcal{K}}_{-}{\mathcal{K}}_{-}(X)$ we can compute ${\mathrm{sat}}\bigcup_{K\in{\mathcal{K}}}K\in{\mathcal{K}}_{-}(X)$ according to Proposition 17 and because $\bigcap_{K\in{\mathcal{K}}}{\mathcal{F}}_{K}={\mathcal{F}}_{{\mathrm{sat}}\bigcup_{K\in{\mathcal{K}}}K}$ it follows by Lemma 18 and Proposition 17 that $\mathcal{F}$} is a computable prebase. Likewise, given ${\mathcal{K}}\in{\mathcal{K}}_{-}{\mathcal{O}}(X)$ we can compute $\bigcap_{U\in{\mathcal{K}}}U\in{\mathcal{O}}(X)$ and because $\bigcap_{U\in{\mathcal{K}}}\Box(U)=\Box(\bigcap_{U\in{\mathcal{K}}}U)$ it follows by Lemma 18 that $\Box$ is a computable prebase. ∎
In particular, we can conclude that all the involved spaces ${\mathcal{O}}(X)$, ${\mathcal{K}}_{-}(X)$, ${\mathcal{A}}_{+}(X)$ and ${\mathcal{C}}(X,Y)$ are computable Kolmogorov spaces under the given conditions (see Corollary 48). The maps specified in Proposition 43 are all known as subbases of some well-known topologies on the respective spaces. We summarize the terminology in the table in Figure 3.
space subbase name of topology
${\mathcal{C}}(X,Y)$ $\triangleright$ compact-open topology
${\mathcal{A}}_{+}(X)$ $\Diamond$ lower Fell topology
${\mathcal{K}}_{-}(X)$ $\Box$ upper Vietoris topology
${\mathcal{O}}(X)$ $\mathcal{F}$} compact-open topology
${\mathcal{O}}(X)$ $\mathcal{U}$} point-open topology
Figure 3. Function and hyperspace topologies and their subbases.
Proposition 43 can also be used as an example to illustrate the different scopes of presubbases, prebases, and bases, respectively.
Example 44.
Let $X$ be some represented space.
  1. (1)
    The neighborhood map $\mathcal{U}$} is a computable presubbase of ${\mathcal{O}}(X)$ that generates the point-open topology on ${\mathcal{O}}(X)$.
  2. (2)
    The canonical computable prebase associated to $\mathcal{U}$} (see Proposition 19) is $\bigcap_{\mathcal{K}}{\mathcal{U}}={\mathcal{F}}$, i.e., the filter map $\mathcal{F}$} that generates the compact-open topology on ${\mathcal{O}}(X)$, whose sequentialization is hence the Scott topology on ${\mathcal{O}}(X)$.
  3. (3)
    The map $\mathcal{F}$} is a computable base of ${\mathcal{O}}(X)$ with respect to the Scott topology if and only if the space $X$ is consonant.
  4. (4)
    The map ${\mathrm{id}}_{{\mathcal{O}}{\mathcal{O}}(X)}$ is always a computable (Lacombe) base of ${\mathcal{O}}(X)$ with respect to the Scott topology.
These examples show that in the general case the concept of a computable presubbase and of a computable prebase might actually be more fruitful and interesting than the concept of a computable base. We have lots of natural presubbases and prebases that reveal useful information about our spaces.
With the following result we provide a few natural constructions of prebases, where we exploit that the index spaces are overt. Overt spaces are exactly those spaces that allow computable projections on open sets.
Proposition 45 (Overt space).
A represented spaces $X$ is overt if and only if the projection
${\mathrm{pr}}_{Y}:{\mathcal{O}}(Y\times X)\to{\mathcal{O}}(Y),U\mapsto\{y\in Y:(\exists x\in X)\;(y,x)\in U\}$
is computable for every represented space $Y$
Proof.
We use the fact that sections $U_{y}:=\{x\in X:(y,x)\in U\}$ and products are computable by Proposition 17.
[…]
Now we can prove the following closure properties for bases $B:R\to{\mathcal{O}}(X)$ with overt index spaces $R$. We note that ${\mathcal{O}}(X)$ is always overt, hence every computable Kolmogorov space $X$ has a computable base ${\mathrm{id}}_{{\mathcal{O}}(X)}:{\mathcal{O}}(X)\to{\mathcal{O}}(X)$ with an overt index space. By $X\sqcup Y$ we denote the coproduct of two represented spaces $X$ and $Y$, which is the set $X\sqcup Y=(\{0\}\times X)\cup(\{1\}\times Y)$ with its canonical representation and by $X\sqcap Y$ we denote the meet $X\sqcap Y=X\cap Y$, which is represented with names that are pairs, where one component represents the point as a point in $X$ and the second component the same point as a point in $Y$. We can identify the space $X\sqcap Y$ with the subspace $\{(x,y)\in X\times Y:x=y\}$ of $X\times Y$. Hence, the results on $X\sqcap Y$ could also be derived from the results on products and subspaces.
Proposition 46 (Constructions of overt prebases).
Let $X$ and $Y$ be represented spaces and let $Z\subseteq X$. Let $S$ and $R$ be overt represented spaces. If $B_{X}:R\to{\mathcal{O}}(X)$ and $B_{Y}:S\to{\mathcal{O}}(Y)$ are computable presubbases (prebases), then so are:
  1. (1)
    $B_{X\times Y}:R\times S\to{\mathcal{O}}(X\times Y),(r,s)\mapsto B_{X}(r)\times B_{Y}(s)$.
  2. (2)
    $B_{Y^{\mathbb{N}}}:S^{*}\to{\mathcal{O}}(Y^{\mathbb{N}}),(s_{1},...,s_{n})\mapsto B_{Y}(s_{1})\times...\times B_{Y}(s_{n})\times Y^{\mathbb{N}}$.
  3. (3)
    $B_{Z}:R\to{\mathcal{O}}(Z),r\mapsto B_{X}(r)\cap Z$.
  4. (4)
    $B_{X\sqcup Y}:R\sqcup S\to{\mathcal{O}}(X\sqcup Y),t\mapsto\left\{\begin{array}[]{ll}B_{X}(t)&\mbox{if $t\in R$}\\ B_{Y}(t)&\mbox{if $t\in S$}\end{array}\right.$.
  5. (5)
    $B_{X\sqcap Y}:R\times S\to{\mathcal{O}}(X\sqcap Y),(r,s)\mapsto B_{X}(r)\cap B_{Y}(s)$.
If $B_{X}$ and $B_{Y}$ are even computable bases, then so is $B_{X\sqcup Y}$. Overtness of $R$ is not needed for (3).
Proof.
Let $B_{X}$ and $B_{Y}$ be computable presubbases. We obtain the transposes
  1. (1)
    $B_{X\times Y}^{\mathsf{{T}}}:X\times Y\to{\mathcal{O}}(R\times S),(x,y)\mapsto B_{X}^{\mathsf{{T}}}(x)\times B_{Y}^{\mathsf{{T}}}(y)$.
  2. (2)
    $B_{Y^{\mathbb{N}}}^{\mathsf{{T}}}:Y^{\mathbb{N}}\to{\mathcal{O}}(S^{*}),(y_{n})_{n\in{\mathbb{N}}}\mapsto\bigsqcup_{n\in{\mathbb{N}}}\left(B_{Y}^{\mathsf{{T}}}(y_{0})\times...\times B_{Y}^{\mathsf{{T}}}(y_{n-1})\right)$.
  3. (3)
    $B_{Z}^{\mathsf{{T}}}:Z\to{\mathcal{O}}(R),z\mapsto B_{X}^{\mathsf{{T}}}(z)$.
  4. (4)
    $B_{X\sqcup Y}^{\mathsf{{T}}}:X\sqcup Y\to{\mathcal{O}}(R\sqcup S),z\mapsto\left\{\begin{array}[]{ll}B_{X}^{\mathsf{{T}}}(z)&\mbox{if $z\in X$}\\ B_{Y}^{\mathsf{{T}}}(z)&\mbox{if $z\in Y$}\end{array}\right.$.
  5. (5)
    $B_{X\sqcap Y}^{\mathsf{{T}}}:X\sqcap Y\to{\mathcal{O}}(R\times S),z\mapsto B_{X}^{\mathsf{{T}}}(z)\times B_{Y}^{\mathsf{{T}}}(z)$.
It is easy to see that all these maps are computable and we are going to show that they are computable embeddings. Hence the maps listed in the proposition are all computable presubbases.
(1) and (2) We prove this in detail for the countable case of $B_{Y^{\mathbb{N}}}$ in (2) and we leave the finite case to the reader. Given $U:=\bigsqcup_{n\in{\mathbb{N}}}\left(B_{Y}^{\mathsf{{T}}}(y_{0})\times...\times B_{Y}^{\mathsf{{T}}}(y_{n-1})\right)\in{\mathcal{O}}(S^{*})$ for some $(y_{n})_{n\in{\mathbb{N}}}\in Y^{\mathbb{N}}$ and $n\in{\mathbb{N}}$, we can compute the sections
$V:=\sec_{n+1}(U)=B_{Y}^{\mathsf{{T}}}(y_{0})\times...\times B_{Y}^{\mathsf{{T}}}(y_{n})\in{\mathcal{O}}(S^{n+1})$
and since $S$ is overt, and hence $S^{n}$ too, we can then compute the projection $W:=B_{Y}^{\mathsf{{T}}}(y_{n})\in{\mathcal{O}}(S)$ by Proposition 45 (which also holds uniformly in $n$). From $W$ we can reconstruct $y_{n}$, as $B_{Y}^{\mathsf{{T}}}$ is a computable embedding.
(3) It immediately follows that $B_{Z}^{\mathsf{{T}}}$ is a computable embedding because $B_{X}^{\mathsf{{T}}}$ is one. Overtness of $R$ is not required here.
(4) Given $U\in{\mathcal{O}}(R\sqcup S)$, where $R\sqcup S=(\{0\}\times R)\cup(\{1\}\times S)$, we can compute the sections $U_{0}\in{\mathcal{O}}(R)$ and $U_{1}\in{\mathcal{O}}(S)$. If $U$ is in the image of $B_{X\sqcup Y}^{\mathsf{{T}}}$ exactly one of $U_{0}$ and $U_{1}$ is non-empty and since $R$ and $S$ are overt, we can find out which one it is. If, for instance, $U_{0}\not$$\varnothing$, then $U_{0}=B_{X}^{\mathsf{{T}}}(x)$ and since $B_{X}^{\mathsf{{T}}}$ is a computable embedding, we can reconstruct $x\in X$ from $U_{0}$. Likewise, we can reconstruct $y\in Y$ in the case of $U_{1}\not$$\varnothing$.
(5) Given $U=B_{X}^{\mathsf{{T}}}(z)\times B_{Y}^{\mathsf{{T}}}(z)\in{\mathcal{O}}(R\times S)$ we can compute the projections $B_{X}^{\mathsf{{T}}}(z)$ and $B_{Y}^{\mathsf{{T}}}(z)$ by Proposition 46 and hence we can reconstruct $z\in X$ as well as $z\in Y$, as $B_{X}^{\mathsf{{T}}}$ and $B_{Y}^{\mathsf{{T}}}$ are computable embeddings. This yields $z\in X\sqcap Y$.
Let now $B_{X}$ and $B_{Y}$ even be computable prebases. We claim that in this case all the maps listed in the proposition are computable prebases too. To this end, we need to prove that compact intersections of the given map can be computably obtained as overt unions.
(1) and (2) We prove the claim for $B_{Y^{\mathbb{N}}}$. Given $K\in{\mathcal{K}}_{-}(S^{*})$ we can compute the projection $K_{0}:={\mathrm{pr}}_{0}(K)\in{\mathcal{K}}_{-}({\mathbb{N}})$ onto the natural number component and we can compute some upper bound $m\in{\mathbb{N}}$ on $\max(K_{0})$. Hence, we can compute also the projections $K_{i}:={\mathrm{sat}}({\mathrm{pr}}_{i}(K))\in{\mathcal{K}}_{-}(S)$ on the components i=1,,mi=1,...,m of $K$ by Proposition 17 (which holds uniformly in $i$). For each i=1,,mi=1,...,m we can compute some set $A_{i}\in{\mathcal{A}}_{+}(S)$ with $\bigcap_{s\in K_{i}}B_{Y}(s)=\bigcup_{s\in A_{i}}B_{Y}(s)$. By Proposition 17 (which holds analogously for finite products) we can hence compute $A:=A_{1}\times...\times A_{m}\in{\mathcal{A}}_{+}(S^{*})$ and we obtain (with implicitly bound $n\in{\mathbb{N}}$):
$\bigcap_{(s_{1},...,s_{n})\in K}B_{Y^{\mathbb{N}}}(s_{1},...,s_{n})$ $=$ $\bigcap_{(s_{1},...,s_{n})\in K}(B_{Y}(s_{1})\times...\times B_{Y}(s_{n})\times Y^{\mathbb{N}})$
$=$ $\left(\bigcap_{s_{1}\in K_{1}}B_{Y}(s_{1})\right)\times...\times\left(\bigcap_{s_{m}\in K_{m}}B_{Y}(s_{m})\right)\times Y^{\mathbb{N}}$
$=$ $\left(\bigcup_{s_{1}\in A_{1}}B_{Y}(s_{1})\right)\times...\times\left(\bigcup_{s_{m}\in A_{m}}B_{Y}(s_{m})\right)\times Y^{\mathbb{N}}$
$=$ $\bigcup_{(s_{1},...,s_{m})\in A}B_{Y^{\mathbb{N}}}(s_{1},...,s_{m}).$
We note that the second and fourth equalities hold as the conditions on the $s_{i}$ are independent and for the second equality we also use that the saturated sets $K_{i}={\mathrm{sat}}({\mathrm{pr}}_{i}(K))$ can be replaced by the unsaturated sets ${\mathrm{pr}}_{i}(K)$ in the intersections by Proposition 17.
(3) This is obvious as the intersection with $Z$ can be associated to a compact intersection as well as to an overt union.
(4) It is easy to see that given $K\in{\mathcal{K}}_{-}(R\sqcup S)$ we can compute the sections $K_{0}\in{\mathcal{K}}_{-}(R)$ and $K_{1}\in{\mathcal{K}}_{-}(S)$. This is because $K_{0}\subseteq U\iff K\subseteq U\sqcup Y$ and $K_{1}\subseteq V\iff K\subseteq X\sqcup V$ for $U\in{\mathcal{O}}(R)$ and $V\in{\mathcal{O}}(S)$. With these sections we obtain
$\bigcap_{t\in K}B_{X\sqcup Y}(t)$ $=$ $\left(\bigcap_{r\in K_{0}}B_{X}(t)\right)\sqcup\left(\bigcap_{s\in K_{1}}B_{Y}(s)\right)$
$=$ $\left(\bigcup_{r\in A_{0}}B_{X}(t)\right)\sqcup\left(\bigcup_{s\in A_{1}}B_{Y}(s)\right)=\bigcup_{t\in A_{0}\sqcup A_{1}}B_{X\sqcup Y}(t)$
where $A_{0}\in{\mathcal{A}}_{+}(R)$ and $A_{1}\in{\mathcal{A}}_{+}(S)$ are sets that we can compute from $K_{0}$ and $K_{1}$ by assumption. Hence we can also compute $A:=A_{0}\sqcup A_{1}\in{\mathcal{A}}_{+}(R\sqcup S)$ because $A\cap U\not$$\varnothing$ for $U\in{\mathcal{O}}(R\sqcup S)$ if and only if $A_{0}\cap U_{0}\not$$\varnothing$ or $A_{1}\cap U_{1}\not$$\varnothing$.
(5) If $K\in{\mathcal{K}}_{-}(R\times S)$, then we can compute the projections $K_{1}:={\mathrm{sat}}{\mathrm{pr}}_{1}(K)\in{\mathcal{K}}_{-}(R)$ and $K_{2}:={\mathrm{sat}}{\mathrm{pr}}_{2}(K)\in{\mathcal{K}}_{-}(S)$. We obtain
$\bigcap_{(r,s)\in K}B_{X\sqcap Y}(r,s)$ $=$ $\bigcap_{r\in K_{1}}B_{X}(r)\cap\bigcap_{s\in K_{2}}B_{Y}(s)$
$=$ $\left(\bigcup_{r\in A_{1}}B_{X}(r)\right)\cap\left(\bigcup_{s\in A_{2}}B_{Y}(s)\right)=\bigcup_{(r,s)\in A_{1}\times A_{2}}B_{X}(r)\cap B_{Y}(s)$
and by assumption we can compute corresponding sets $A_{1}\in{\mathcal{A}}_{+}(R)$ and $A_{2}\in{\mathcal{A}}_{+}(S)$.
Let now $B_{X}$ and $B_{Y}$ even be computable bases. That $B_{X\sqcup Y}$ is a base follows from ${\mathcal{O}}(X\sqcup Y)={\mathcal{O}}(X)\sqcup{\mathcal{O}}(Y)$. ∎
Using Proposition 46 and the computable base ${\mathrm{id}}_{{\mathcal{O}}(X)}:{\mathcal{O}}(X)\to{\mathcal{O}}(X)$ we even obtain specific computable prebases for spaces associated to computable Kolmogorov spaces.
Corollary 47 (Computable prebases).
Let $X$ and $Y$ be computable Kolmogorov spaces and let $Z\subseteq X$. Then the following are computable prebases:
  1. (1)
    $B_{X\times Y}:{\mathcal{O}}(X)\times{\mathcal{O}}(Y)\to{\mathcal{O}}(X\times Y),(U,V)\mapsto U\times V$.
  2. (2)
    $B_{Y^{\mathbb{N}}}:{\mathcal{O}}(X)^{*}\to{\mathcal{O}}(Y^{\mathbb{N}}),(U_{1},...,U_{n})\mapsto U_{1}\times...\times U_{n}\times Y^{\mathbb{N}}$.
  3. (3)
    $B_{Z}:{\mathcal{O}}(X)\to{\mathcal{O}}(Z),U\mapsto U\cap Z$.
  4. (4)
    $B_{X\sqcup Y}:{\mathcal{O}}(X)\sqcup{\mathcal{O}}(Y)\to{\mathcal{O}}(X\sqcup Y),U\mapsto U$.
  5. (5)
    $B_{X\sqcap Y}:{\mathcal{O}}(X)\times{\mathcal{O}}(Y)\to{\mathcal{O}}(X\sqcap Y),(U,V)\mapsto U\cap V$.
$B_{X\sqcup Y}$ is even a computable base.
Using Theorem 21, Corollary 47 and Proposition 43 we can directly conclude that computable Kolmogorov spaces have very nice closure properties. We note that for this proof we only need the statements on computable prebases that were relatively easy to establish. By ${\mathcal{A}}_{-}(X)$ we denote the space of closed subsets of $X$ represented as complements of open sets in ${\mathcal{O}}(X)$. This space is hence computably isomorphic to ${\mathcal{O}}(X)$. We also consider the spaces ${\mathcal{A}}(X):={\mathcal{A}}_{+}(X)\sqcap{\mathcal{A}}_{-}(X)$ and ${\mathcal{K}}(X):={\mathcal{A}}_{+}(X)\sqcap{\mathcal{K}}_{-}(X)$ (we note that the latter contains the compact, closed and saturated sets).
Corollary 48 (Closure properties of computable Kolmogorov spaces).
If $X$ and $Y$ are computable Kolmogorov spaces, then so are
  1. (1)
    $X\times Y$, $X\sqcup Y$, $X\sqcap Y$, $Y^{\mathbb{N}}$, and every subspace of $X$.
Let $X$ be a represented space. If $Y$ is a computable Kolmogorov space, then so are
  1. (2)
    ${\mathcal{C}}(X,Y)$, ${\mathcal{O}}(X)$, ${\mathcal{A}}_{+}(X)$, ${\mathcal{A}}_{-}(X)$, ${\mathcal{A}}(X)$, ${\mathcal{K}}_{-}(X)$ and ${\mathcal{K}}(X)$.
Analogous statements hold for continuous Kolmogorov spaces.
In fact, some of these results can also be derived from each other in various ways. For instance, the statement on ${\mathcal{C}}(X,Y)$ applied to the special case of $X={\mathbb{N}}$ yields the statement for $Y^{\mathbb{N}}$ and the special case of $Y={\mathbb{S}}$ yields the statements for ${\mathcal{O}}(X)$ (and ${\mathcal{A}}_{-}(X)$). This implies the statement for ${\mathcal{O}}{\mathcal{O}}(X)$, subspaces of which are computably isomorphic to ${\mathcal{A}}_{+}(X)$ and ${\mathcal{K}}_{-}(X)$, respectively. We obtain the following conclusion (see [Sch02a, Sch02b] for the topological version) using the fact that our product and function space constructions satisfy evaluation and currying properties (see [Wei00, Wei87]).
Corollary 49 (Schröder 2002).
The category of computable Kolmogorov spaces with computable maps as morphisms is cartesian closed.
For Corollaries 48 and 49 it was sufficient to have computable presubbases for the respective spaces. The fact that we have even computable prebases for some of these spaces allows further conclusions regarding the underlying topologies. Using Corollaries 16 and 47 we obtain the following conclusions. By ${\mathcal{O}}(X)\otimes{\mathcal{O}}(Y)$ and $\bigotimes_{i\in{\mathbb{N}}}{\mathcal{O}}(X)$ we denote the respective product topologies, by ${\mathcal{O}}(X)|_{Z}$ the subspace topology of $Z\subseteq X$ and by ${\mathcal{O}}(X)\wedge{\mathcal{O}}(Y)$ the meet topology generated by the base $B_{X\sqcap Y}$.
Corollary 50 (Topologies and set-theoretic constructions).
Let $X$ and $Y$ be admissibly represented ${\mathsf{{T}}}_{0}$ spaces and let $Z\subseteq X$ be a subspace. Then:
  1. (1)
    ${\mathcal{O}}(X\times Y)={\mathrm{seq}}({\mathcal{O}}(X)\otimes{\mathcal{O}}(Y))$,
  2. (2)
    ${\mathcal{O}}(X^{\mathbb{N}})={\mathrm{seq}}(\bigotimes_{i\in{\mathbb{N}}}{\mathcal{O}}(X))$,
  3. (3)
    ${\mathcal{O}}(Z)={\mathrm{seq}}({\mathcal{O}}(X)|_{Z})$,
  4. (4)
    ${\mathcal{O}}(X\sqcap Y)={\mathrm{seq}}({\mathcal{O}}(X)\wedge{\mathcal{O}}(Y))$.
This result follows from Corollary 47 in combination with Corollary 16. With the help of Theorem 22 we can transfer (1) and (2) of this result then to arbitrary represented ${\mathsf{{T}}}_{0}$ spaces, since $\delta\mapsto\delta^{\bullet}$ commutes with products [Sch02b].
6. Hyperspace and function space topologies
In this section we continue to discuss what can be concluded from our results regarding hyperspace and function space topologies. Once again, these results are well-known by the work of Schröder [Sch02a] and the purpose here is to demonstrate how these results can be obtained with the help of the concept of a computable presubbase or prebase.
As a preparation of our results we need the following lemma, which shows that the subbases $\Diamond$ and $\triangleright$ satisfy the convergent intersection property. This allows us to apply Corollary 12.
Lemma 51 (Countable intersections).
Let $X$ and $Y$ be represented space. Then the presubbases
  1. (1)
    $\Diamond:{\mathcal{O}}(Y)\to{\mathcal{O}}{\mathcal{A}}_{+}(Y)$ and
  2. (2)
    $\triangleright:{\mathcal{K}}_{-}(X)\times{\mathcal{O}}(Y)\to{\mathcal{O}}{\mathcal{C}}(X,Y)$
satisfy the convergent intersection property.
Proof.
Let $(U_{n})_{n\in{\mathbb{N}}_{\infty}}$ and $(K_{n})_{n\in{\mathbb{N}}_{\infty}}$ be converging sequences in ${\mathcal{O}}(Y)$ and ${\mathcal{K}}_{-}(X)$, respectively. We prove that there exists some $k\in{\mathbb{N}}$ such that
  1. (1)
    $\bigcap_{n\in{\mathbb{N}}_{\infty}}\Diamond U_{n}=\Diamond U_{\infty}\cap\bigcap_{n=0}^{k}\Diamond U_{n}$.
  2. (2)
    $\bigcap_{n\in{\mathbb{N}}_{\infty}}(K_{n}\triangleright U_{n})=(K_{\infty}\triangleright U_{\infty})\cap\bigcap_{n=0}^{k}(K_{n}\triangleright U_{n})$.
By Proposition 8 convergence of $(U_{n})_{n\in{\mathbb{N}}_{\infty}}$ implies that there is a $k\in{\mathbb{N}}$ with $U_{\infty}\subseteq\left(\bigcap_{n\geq k}U_{n}\right)^{\circ}$. In particular, $U_{\infty}\subseteq U_{n}$ for all $n\geq k$. The inclusion $\subseteq$” in (1) and (2) is clear in both cases and we need to show why the inverse inclusions hold.
(1) Let $A\in\Diamond U_{\infty}\cap\bigcap_{n=0}^{k}\Diamond U_{n}$. Then, in particular, $A\cap U_{\infty}\not$$\varnothing$ and hence $A\cap U_{n}\not$$\varnothing$ for all $n\geq k$. This proves the claim.
(2) Let $f\in(K_{\infty}\triangleright U_{\infty})\cap\bigcap_{n=0}^{k}(K_{n}\triangleright U_{n})$ be a continuous function $f:X\to Y$. Then, in particular, $f(K_{\infty})\subseteq U_{\infty}$. As $(f(K_{n}))_{n\in{\mathbb{N}}}$ converges in ${\mathcal{K}}_{-}(Y)$, we obtain (see the independently proved statement in Theorem 52 (2)) that there exists a $k^{\prime}\in{\mathbb{N}}$ with $f(K_{n})\subseteq U_{\infty}$ for $n\geq k^{\prime}$ and hence $f(K_{n})\subseteq U_{n}$ for all $n\geq k^{\prime\prime}:=\max(k,k^{\prime})$. This proves the claim. ∎
We use the terminology for topologies introduced in the table in Figure 3 and we add some terminology:
  1. (1)
    The upper Fell topology on the hyperspace ${\mathcal{A}}_{-}(X)$ is the topology inherited (by complementation) from the compact-open topology on ${\mathcal{O}}(X)$ and the Fell topology on ${\mathcal{A}}(X)$ is the meet topology of the lower and upper Fell topologies.
  2. (2)
    The Vietoris topology on ${\mathcal{K}}(X)$ is the meet topology of the lower Fell topology and the upper Vietoris topology.
From Proposition 43 we can obtain information on the topologies of the respective hyperspaces, a fact which we already indicated in Example 44.
Theorem 52 (Hyperspace and function space topologies).
Let $X$ be a represented space and let $Y$ be an admissibly represented ${\mathsf{{T}}}_{0}$ space.
  1. (1)
    ${\mathcal{O}}(X)$ is endowed with the Scott topology, which is the sequentialization of the compact-open topology.
  2. (2)
    ${\mathcal{K}}_{-}(X)$ and ${\mathcal{K}}(X)$ are endowed with topologies that are the sequentializations of the upper Vietoris and the Vietoris topology, respectively.
  3. (3)
    ${\mathcal{A}}_{+}(X)$ and ${\mathcal{A}}(X)$ are endowed with topologies that are the sequentializations of the lower Fell and the Fell topology, respectively.
  4. (4)
    ${\mathcal{C}}(X,Y)$ is endowed with the sequentialization of the compact-open topology.
In particular, all the mentioned spaces are admissibly represented with respect to the given topologies.
Proof.
(1) This is the statement of Corollary 9. That ${\mathcal{O}}{\mathcal{O}}(X)$ is the sequentialization of the topology generated by the sets ${\mathcal{F}}_{K}$ for (saturated) compact $K\subseteq X$ follows also from Proposition 43 in combination with Corollary 16.
(2) That ${\mathcal{O}}{\mathcal{K}}_{-}(X)$ is the sequentialization of the topology generated by the sets $\Box U$ for open $U\subseteq X$ follows from Proposition 43 in combination with Corollary 16.
(3) That ${\mathcal{O}}{\mathcal{A}}_{+}(X)$ is the sequentialization of the lower Fell topology generated by the sets $\Diamond U$ for open $U\subseteq X$ follows from Theorem 2, Corollary 12 and Lemma 51. With the help of Corollary 50 this implies that ${\mathcal{K}}(X)={\mathcal{A}}_{+}(X)\sqcap{\mathcal{K}}_{-}(X)$ and ${\mathcal{A}}(X)={\mathcal{A}}_{+}(X)\sqcap{\mathcal{A}}_{-}(X)$ are endowed with the sequentializations of the Vietoris and Fell topologies, respectively.
(4) If $Y$ is an admissibly represented ${\mathsf{{T}}}_{0}$ space then Theorem 22 allows us to replace it by a computable Kolmogorov space without affecting the topological structure. In fact, the resulting representation is in the same continuous equivalence class as the original and hence so are the corresponding representations of ${\mathcal{C}}(X,Y)$. By Theorem 2, Corollary 12 and Lemma 51 we know that ${\mathcal{O}}{\mathcal{C}}(X,Y)$ is the sequentialization of the compact-open topology on ${\mathcal{C}}(X,Y)$. ∎
The examples of the spaces ${\mathcal{A}}_{+}(X)$ and ${\mathcal{C}}(X,Y)$ show that even in cases where we have only a computable presubbase (and not a computable prebase) it might happen that we obtain as a topology on our space the sequentialization of the topology generated by the presubbase.
7. Epilogue
We close this section with a brief recap of our results from a somewhat different perspective. The results in this article can be viewed as exploiting a Galois connection between maps
[…]
Let us fix a set $X$ and let us denote by $\mathsf{REP}_{0}$ the set of representations $\delta:\subseteq{\mathbb{N}}^{\mathbb{N}}\to X$ with a ${\mathsf{{T}}}_{0}$ final topology and let us denote by $\mathsf{PRE}_{0}$ the set of presubbases $B:\subseteq{\mathbb{N}}^{\mathbb{N}}\to 2^{X}$ that induce a ${\mathsf{{T}}}_{0}$ topology on XX. [^3] We call $B:\subseteq{\mathbb{N}}^{\mathbb{N}}\to 2^{X}$ a presubbase if $(B_{y})_{y\in D}$ with $D:={\mathrm{dom}}(B)$ is a presubbase in the sense of our definition.
  1. (1)
    The map $\Delta:\mathsf{PRE}_{0}\to\mathsf{REP}_{0},B\mapsto\delta^{B}$ assigns to any presubbase $B$ the presubbase representation $\delta^{B}:\subseteq{\mathbb{N}}^{\mathbb{N}}\to X$.
  2. (2)
    The map $\nabla:\mathsf{REP}_{0}\to\mathsf{PRE}_{0},\delta\mapsto B_{\delta}$ assigns to every representation $\delta$ of $X$ the induced representation $\delta_{{\mathcal{O}}(X)}$ as presubbase $B_{\delta}:\subseteq{\mathbb{N}}^{\mathbb{N}}\to 2^{X}$.
The presubase $B_{\delta}$ corresponds to the base ${\mathrm{id}}:{\mathcal{O}}(X)\to{\mathcal{O}}(X)$. The maps $\Delta$ and $\nabla$ induce an antitone Galois connection, if we define a computable reducibility $\leq$ for $\mathsf{PRE}_{0}$ as for $\mathsf{REP}_{0}$ (for $\mathsf{REP}_{0}$ we use the usual computable reducibility, as introduced before Theorem 22). Namely for $B_{1},B_{2}\in\mathsf{PRE}_{0}$ we define analogously $B_{1}\leq B_{2}$ if there is some computable $F:\subseteq{\mathbb{N}}^{\mathbb{N}}\to{\mathbb{N}}^{\mathbb{N}}$ with $B_{1}(p)=B_{2}F(p)$ for all $p\in{\mathrm{dom}}(B_{1})$. The proof of the following result is then straightforward and a simple consequence of properties of the transposition operation.
Theorem 53 (Galois connection for representations and presubbases).
Fix a representable space $X$. Then for every $\delta\in\mathsf{REP}_{0}$ and $B\in\mathsf{PRE}_{0}$ we have
$\delta\leq\delta^{B}\iff B\leq B_{\delta}.$
Proof.
Let $\delta:\subseteq{\mathbb{N}}^{\mathbb{N}}\to X$ be a representation with a ${\mathsf{{T}}}_{0}$ topology ${\mathcal{O}}(X)$ and induced representation $\delta_{{\mathcal{O}}(X)}$ of ${\mathcal{O}}(X)$ and let $B:\subseteq{\mathbb{N}}^{\mathbb{N}}\to 2^{X}$ be a presubbase that induces a ${\mathsf{{T}}}_{0}$ topology on $X$. Then we obtain:
$\delta\leq\delta^{B}$ $\iff$ $B^{\mathsf{{T}}}:\subseteq X\to{\mathcal{O}}({\mathbb{N}}^{\mathbb{N}})\mbox{ is computable with respect to $\delta$}$
$\iff$ $B:\subseteq{\mathbb{N}}^{\mathbb{N}}\to{\mathcal{O}}(X)\mbox{ is computable with respect to $\delta_{{\mathcal{O}}(X)}$}$
$\iff$ $B\leq B_{\delta},$
which proves the claim. ∎
We also obtain that ${\mathrm{range}}(\Delta)$ is exactly the set of representations of $X$ that turn $X$ into a computable Kolmogorov space (up to computable equivalence) and ${\mathrm{range}}(\nabla)$ is exactly the set of representations of $\mathsf{qcb}_{0}$ topologies for $X$ (up to computable equivalence).
Any antitone Galois connection yields corresponding closure operators, sometimes called monad and comonad. In the case of our Galois connection we obtain:
  1. (1)
    $\Delta\circ\nabla:\mathsf{REP}_{0}\to\mathsf{REP}_{0},\delta\mapsto\delta^{\bullet}:=\delta^{B_{\delta}}$ and
  2. (2)
    $\nabla\circ\Delta:\mathsf{PRE}_{0}\to\mathsf{PRE}_{0},B\mapsto B_{\bullet}:=B_{\delta^{B}}$,
We summarize some properties of these closure operators (see Theorem 22, Corollary 14 and Proposition 20):
  1. (1)
    $\delta\mapsto\delta^{\bullet}$ maps every represented ${\mathsf{{T}}}_{0}$ space $(X,\delta)$ to a computable Kolmogorov space $X^{\bullet}:=(X,\delta^{\bullet})$ with $\delta\leq\delta^{\bullet}$. The induced underlying final topologies and the representations thereof are preserved by this operation (up to computable equivalence).
  2. (2)
    $B\mapsto B_{\bullet}$ maps any presubbase $B$ of a ${\mathsf{{T}}}_{0}$ topology on $X$ to a computable base $B_{\bullet}$ of the sequentialization ${\mathrm{seq}}(\tau)$ of the topology $\tau$ that is generated by the compact intersections $\bigcap_{\mathcal{K}}B$. In particular, ${\mathrm{range}}(B)\subseteq{\mathrm{range}}(B_{\bullet})$. The operation preserves the induced presubbase representations of $X$ (up to computable equivalence).
Since both application are closure operators, a double application does not yield anything new, i.e., $\delta^{\bullet\bullet}\equiv\delta^{\bullet}$ and ${\mathrm{range}}(B_{\bullet\bullet})={\mathrm{range}}(B_{\bullet})$. We give two examples of the action of these closure operator that indicates that these are natural operations. The first example is taken from Example 44, the second one is due to Schröder (personal communication).
Example 54 (Action of closure operators).
  1. (1)
    If we start with the presubbase $B$ for the point-open topology on ${\mathcal{O}}(X)$ then we obtain a base $B_{\bullet}$ of the Scott topology. Both bases induce computably equivalent presubbase representations of ${\mathcal{O}}(X)$.
  2. (2)
    If we start with a decimal representation $\rho$ of the real numbers $\mathbb{R}$}, then we obtain a representation $\rho^{\bullet}$ that is computably equivalent to the Cauchy representation of the reals. Both representations have the same final topology (the Euclidean topology).
In light of the Galois connection between presubbases and representations we can interpret our results as follows. A lot of results that were originally proved with the help of representations and, for instance, the closure operator $\delta\mapsto\delta^{\bullet}$ by Schröder (for example the topological version of Corollary 49 in [Sch02a, Section 4.3]) can equally well be derived with the help of presubbases. Both perspectives correspond to different sides of the same medal and they highlight different aspects of the same mathematical structure.
The purpose of this article was to demonstrate the usefulness of considering presubbases. The benefit of considering presubbases is that they allow us to talk about topologies using their usual subbases (see, e.g., Proposition 43) and the usual constructions of new subbases from given ones (see, e.g., Proposition 46). This enriches the mathematical vocabulary that we have to discuss computable topology.
Acknowledgments
This work was supported by the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) – project number 554999067, by the National Research Foundation of South Africa (NRF) – grant number 151597, by the Alexander-von-Humboldt Foundation and by the Research Institute for Mathematical Sciences, an international joint research center based at Kyoto University. We would like to thank Matthias Schröder for discussions on the content of this article that have helped to streamline some proofs and their presentations.